The Caml Light system release 0.74

Documentation and user’s manual

Xavier Leroy December 2, 1997

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1997 Institut National de Recherche en Informatique et Automatique 2

Contents

I Getting started 7

1 Installation instructions 9 1.1 The Unix version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 1.2 The Macintosh version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 1.3 The MS-Windows version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

II The Caml Light language reference manual 11

2 The core Caml Light language 13 2.1 Lexical conventions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.2 Global names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 2.3 Values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.4 Type expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 2.5 Constants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 2.6 Patterns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 2.7 Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 2.8 Global definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 2.9 Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 2.10 Module implementations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 2.11 Module interfaces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

3 Language extensions 35 3.1 Streams, parsers, and printers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 3.2 Guards . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 3.3 Range patterns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 3.4 Recursive definitions of values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 3.5 Local definitions using where . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 3.6 Mutable variant types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 3.7 String access . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 3.8 Alternate syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 3.9 Infix symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 3.10 Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

III The Caml Light commands 43

4 Batch compilation (camlc) 45 4.1 Overview of the compiler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 4.2 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 4.3 Modules and the file system . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 4.4 Common errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

5 The toplevel system (camllight) 53 5.1 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 5.2 Toplevel control functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 5.3 The toplevel and the module system . . . . . . . . . . . . . . . . . . . . . . . . . . 58 5.4 Common errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 5.5 Building custom toplevel systems: camlmktop . . . . . . . . . . . . . . . . . . . . . 60 5.6 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

6 The runtime system (camlrun) 63 6.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 6.2 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 6.3 Common errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

7 The librarian (camllibr) 67 7.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 7.2 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 7.3 Turning code into a library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

8 Lexer and parser generators (camllex, camlyacc) 71 8.1 Overview of camllex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 8.2 Syntax of lexer definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 8.3 Overview of camlyacc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 8.4 Syntax of grammar definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 8.5 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 8.6 A complete example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76

9 The debugger (camldebug) 79 9.1 Compiling for debugging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 9.2 Invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 9.3 Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 9.4 Executing a program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 9.5 Breakpoints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 9.6 The call stack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 9.7 Examining variable values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 9.8 Controlling the debugger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 9.9 Miscellaneous commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 10 Profiling (camlpro) 91

10.1 Compiling for profiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 10.2 Profiling an execution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 10.3 Printing profiling information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 10.4 Known bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

11 Using Caml Light under Emacs 95 11.1 Updating your .emacs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 11.2 The caml editing mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 11.3 Running the toplevel as an inferior process . . . . . . . . . . . . . . . . . . . . . . . 96 11.4 Running the debugger as an inferior process . . . . . . . . . . . . . . . . . . . . . . 97

12 Interfacing C with Caml Light 99 12.1 Overview and compilation information . . . . . . . . . . . . . . . . . . . . . . . . . 99 12.2 The value type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 12.3 Representation of Caml Light data types . . . . . . . . . . . . . . . . . . . . . . . . 102 12.4 Operations on values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 12.5 Living in harmony with the garbage collector . . . . . . . . . . . . . . . . . . . . . 105 12.6 A complete example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107

IV The Caml Light library 111

13 The core library 113 13.1 bool: boolean operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 13.2 builtin: base types and constructors . . . . . . . . . . . . . . . . . . . . . . . . . 114 13.3 char: character operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 13.4 eq: generic comparisons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 13.5 exc: exceptions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 13.6 fchar: character operations, without sanity checks . . . . . . . . . . . . . . . . . . 117 13.7 float: operations on floating-point numbers . . . . . . . . . . . . . . . . . . . . . 117 13.8 fstring: string operations, without sanity checks . . . . . . . . . . . . . . . . . . 119 13.9 fvect: operations on vectors, without sanity checks . . . . . . . . . . . . . . . . . 119 13.10 int: operations on integers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 13.11 io: buffered input and output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 13.12 list: operations on lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 13.13 pair: operations on pairs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 13.14 ref: operations on references . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 13.15 stream: operations on streams . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 13.16 string: string operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 13.17 vect: operations on vectors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134

14 The standard library 137 14.1 arg: parsing of command line arguments . . . . . . . . . . . . . . . . . . . . . . . . 137 14.2 baltree: basic balanced binary trees . . . . . . . . . . . . . . . . . . . . . . . . . . 138 14.3 filename: operations on file names . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 14.4 format: pretty printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 14.5 gc: memory management control and statistics . . . . . . . . . . . . . . . . . . . . 147 14.6 genlex: a generic lexical analyzer . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 14.7 hashtbl: hash tables and hash functions . . . . . . . . . . . . . . . . . . . . . . . . 150 14.8 lexing: the run-time library for lexers generated by camllex . . . . . . . . . . . . 151 14.9 map: association tables over ordered types . . . . . . . . . . . . . . . . . . . . . . . 153 14.10 parsing: the run-time library for parsers generated by camlyacc . . . . . . . . . . 153 14.11 printexc: a catch-all exception handler . . . . . . . . . . . . . . . . . . . . . . . . 154 14.12 printf: formatting printing functions . . . . . . . . . . . . . . . . . . . . . . . . . 154 14.13 queue: queues . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 14.14 random: pseudo-random number generator . . . . . . . . . . . . . . . . . . . . . . . 157 14.15 set: sets over ordered types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 14.16 sort: sorting and merging lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 14.17 stack: stacks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 14.18 sys: system interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159

15 The graphics library 163 15.1 graphics: machine-independent graphics primitives . . . . . . . . . . . . . . . . . 164

16 The unix library: Unix system calls 171 16.1 unix: interface to the Unix system . . . . . . . . . . . . . . . . . . . . . . . . . . . 171

17 The num library: arbitrary-precision rational arithmetic 191 17.1 num: operations on numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 17.2 arith_status: flags that control rational arithmetic . . . . . . . . . . . . . . . . . 194

18 The str library: regular expressions and string processing 197 18.1 str: regular expressions and high-level string processing . . . . . . . . . . . . . . . 197

V Appendix 201

19 Further reading 203 19.1 Programming in ML . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 19.2 Descriptions of ML dialects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 19.3 Implementing functional programming languages . . . . . . . . . . . . . . . . . . . 205 19.4 Applications of ML . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206

Index to the library 207

Index of keywords 216

Foreword

This manual documents the release 0.74 of the Caml Light system. It is organized as follows.

• Part I, “Getting started”, explains how to install Caml Light on your machine.

• Part II, “The Caml Light language reference manual”, is the reference description of the Caml Light language.

• Part III, “The Caml Light commands”, documents the Caml Light compiler, toplevel system, and programming utilities.

• Part IV, “The Caml Light library”, describes the modules provided in the standard library.

• Part V, “Appendix”, contains a short bibliography, an index of all identifiers defined in the standard library, and an index of Caml Light keywords.

Conventions

The Caml Light system comes in several versions: for Unix machines, for Macintoshes, and for PCs. The parts of this manual that are specific to one version are presented as shown below:

Unix: This is material specific to the Unix version.

Mac: This is material specific to the Macintosh version.

PC: This is material specific to the PC version.

License

The Caml Light system is copyright \begin{matrix}c \\ ◯ \end{matrix} 1989, 1990, 1991, 1992, 1993, 1994, 1995, 1996, 1997 Institut National de Recherche en Informatique et en Automatique (INRIA). INRIA holds all ownership rights to the Caml Light system. See the file COPYRIGHT in the distribution for the copyright notice.

The Caml Light system can be freely redistributed. More precisely, INRIA grants any user of the Caml Light system the right to reproduce it, provided that the copies are distributed under the conditions given in the COPYRIGHT file. The present documentation is distributed under the same conditions.

6 Foreword

Availability by FTP

The complete Caml Light distribution resides on the machine ftp.inria.fr. The distribution files can be transferred by anonymous FTP:

Host:  ftp.inria.fr (Internet address 192.93.2.54)

Login name: anonymous
Password:  your e-mail address

Directory:  lang/caml-light

Files:  see the index in file README

Part I

Chapter 1

Getting started Installation instructions

This chapter explains how to install Caml Light on your machine.

1.1 The Unix version

Requirements. Any machine that runs under one of the various flavors of the Unix operating system, and that has a flat, non-segmented, 32-bit or 64-bit address space. 4M of RAM, 2M of free disk space. The graphics library requires X11 release 4 or later.

Installation. The Unix version is distributed in source format, as a compressed tar file named cl74unix.tar.gz. To extract, move to the directory where you want the source files to reside, transfer cl74unix.tar.gz to that directory, and execute

        zcat cl74unix.tar.gz | tar xBf -

This extracts the source files in the current directory. The file INSTALL contains complete instructions on how to configure, compile and install Caml Light. Read it and follow the instructions.

Troubleshooting. See the file INSTALL.

1.2 The Macintosh version

Requirements. Any Macintosh with at least 1M of RAM (2M is recommended), running System 6 or 7. About 850K of free space on the disk. The parts of the Caml Light system that support batch compilation currently require the Macintosh Programmer’s Workshop (MPW) version 3.2. MPW is Apple’s development environment, and it is distributed by APDA, Apple’s Programmers and Developers Association. See the file READ ME in the distribution for APDA’s address.

Installation. Create the folder where the Caml Light files will reside. Double-click on the file cl74macbin.sea from the distribution. This displays a file dialog box. Open the folder where the Caml Light files will reside, and click on the Extract button. This will re-create all files from the distribution in the Caml Light folder.

To test the installation, double-click on the application Caml Light. The “Caml Light output” window should display something like

        >       Caml Light version 0.74
        #

In the “Caml Light input” window, enter 1+2;; and press the Return key. The “Caml Light output” window should display:

        >       Caml Light version 0.74
        #1+2;;
        - : int = 3
        #

Select “Quit” from the “File” menu to return to the Finder.

If you have MPW, you can install the batch compilation tools as follows. The tools and scripts from the tools folder must reside in a place where MPW will find them as commands. There are two ways to achieve this result: either copy the files in the tools folder to the Tools or the Scripts folder in your MPW folder; or keep the files in the tools folder and add the following line to your UserStartup file (assuming Caml Light resides in folder Caml Light on the disk named

My HD):
        Set Commands "{Commands},My HD:Caml Light:tools:"

In either case, you now have to edit the camlc script, and replace the string

        Macintosh HD:Caml Light:lib:

(in the first line) with the actual pathname of the lib folder. For example, if you put Caml Light in folder Caml Light on the disk named My HD, the first line of camlc should read:

        Set stdlib "My HD:Caml Light:lib:"

Troubleshooting. Here is one commonly encountered problem.

Cannot find file stream.zi

(Displayed in the “Caml Light output” window, with an alert box telling you that Caml Light has terminated abnormally.) This is an installation error. The folder named lib in the distribution must always be in the same folder as the Caml Light application. It’s OK to move the application to another folder; but remember to move the lib directory to the same folder. (To return to the Finder, first select “Quit” from the “File” menu.)

1.3 The MS-Windows version

Requirements. A PC equipped with a 80386, 80486 or Pentium processor, running MS Windows 3.x, Windows 95 or Windows NT. About 3M of free space on the disk. At least 8M of RAM is recommended.

Installation. The MS-Windows version is distributed as a self-extracting, self-installing archive named cl74win.exe. Simply run it and follow the steps of the installation program.

Part II

The Caml Light language reference manual The core Caml Light language

Foreword

This document is intended as a reference manual for the Caml Light language. It lists all language constructs, and gives their precise syntax and informal semantics. It is by no means a tutorial introduction to the language: there is not a single example. A good working knowledge of the language, as provided by the companion tutorial Functional programming using Caml Light, is assumed.

No attempt has been made at mathematical rigor: words are employed with their intuitive meaning, without further definition. As a consequence, the typing rules have been left out, by lack of the mathematical framework required to express them, while they are definitely part of a full formal definition of the language. The reader interested in truly formal descriptions of languages from the ML family is referred to The definition of Standard ML and Commentary on Standard ML, by Milner, Tofte and Harper, MIT Press.

Warning

Several implementations of the Caml Light language are available, and they evolve at each release. Consequently, this document carefully distinguishes the language and its implementations. Implementations can provide extra language constructs; moreover, all points left unspecified in this reference manual can be interpreted differently by the implementations. The purpose of this reference manual is to specify those features that all implementations must provide.

Notations

The syntax of the language is given in BNF-like notation. Terminal symbols are set in typewriter font (like this). Non-terminal symbols are set in italic font (like that). Square brackets […] denote optional components. Curly brackets {…} denotes zero, one or several repetitions of the enclosed components. Curly bracket with a trailing plus sign {…}^{+} denote one or several repetitions of the enclosed components. Parentheses (…) denote grouping.

2.1 Lexical conventions

Blanks

The following characters are considered as blanks: space, newline, horizontal tabulation, carriage return, line feed and form feed. Blanks are ignored, but they separate adjacent identifiers, literals and keywords that would otherwise be confused as one single identifier, literal or keyword.

Comments

Comments are introduced by the two characters (*, with no intervening blanks, and terminated by the characters *), with no intervening blanks. Comments are treated as blank characters. Comments do not occur inside string or character literals. Nested comments are correctly handled.

Identifiers

ident ::= letter {letter | 0 … 9 | _}
letter ::= A … Z | a … z

Identifiers are sequences of letters, digits and _(the underscore character), starting with a letter. Letters contain at least the 52 lowercase and uppercase letters from the ASCII set. Implementations can recognize as letters other characters from the extended ASCII set. Identifiers cannot contain two adjacent underscore characters (__). Implementation may limit the number of characters of an identifier, but this limit must be above 256 characters. All characters in an identifier are meaningful.

Integer literals

integer-literal ::= [-] {0 … 9}^{+}
| [-] (0x | 0X) {0 … 9 | A … F | a … f}^{+}
| [-] (0o | 0O) {0 … 7}^{+}
| [-] (0b | 0B) {0 … 1}^{+}

An integer literal is a sequence of one or more digits, optionally preceded by a minus sign. By default, integer literals are in decimal (radix 10). The following prefixes select a different radix:

Prefix Radix
0x, 0X hexadecimal (radix 16)
0o, 0O octal (radix 8)
0b, 0B binary (radix 2)

(The initial 0 is the digit zero; the O for octal is the letter O.)

Floating-point literals

float-literal ::= [-] {0 … 9}^{+} [. {0 … 9}] [(e | E) [+ | -] {0 … 9}^{+}]

Floating-point decimals consist in an integer part, a decimal part and an exponent part. The integer part is a sequence of one or more digits, optionally preceded by a minus sign. The decimal part is a decimal point followed by zero, one or more digits. The exponent part is the character e or E followed by an optional + or - sign, followed by one or more digits. The decimal part or the exponent part can be omitted, but not both to avoid ambiguity with integer literals.

Character literals

char-literal ::= ` regular-char `

                 | `\ (\ | ` | n | t | b | r) `

| `\ (0 … 9) (0 … 9) (0 … 9) `

Character literals are delimited by ` (backquote) characters. The two backquotes enclose either one character different from ` and \, or one of the escape sequences below:

Sequence Character denoted
\\  backslash (\)
\`  backquote (`)
\n  newline (LF)
\r  return (CR)
\t  horizontal tabulation (TAB)
\b  backspace (BS)
\ddd  the character with ASCII code ddd in decimal

String literals

string-literal ::= " {string-character} "
string-character ::= regular-char
| \ (\ | " | n | t | b | r)
| \ (0 … 9) (0 … 9) (0 … 9)

String literals are delimited by " (double quote) characters. The two double quotes enclose a sequence of either characters different from " and \, or escape sequences from the table below:

Sequence Character denoted
\\  backslash (\)
\"  double quote (")
\n  newline (LF)
\r  return (CR)
\t  horizontal tabulation (TAB)
\b  backspace (BS)
\ddd  the character with ASCII code ddd in decimal

Implementations must support string literals up to 2^{16} − 1 characters in length (65535 characters).

Keywords

The identifiers below are reserved as keywords, and cannot be employed otherwise:

        and    as     begin       do      done      downto
        else   end    exception   for     fun       function
        if     in     let         match   mutable   not
        of     or     prefix      rec     then      to
        try    type   value       where   while     with

The following character sequences are also keywords:

        #    !    !=   &    (    )    *    *.   +    +.
        ,    -    -.   ->   .    .(   /    /.   :    ::
        :=   ;    ;;   <    <.   <-   <=   <=.  <>   <>.
        =    =.   ==   >    >.   >=   >=.  @    [    [|
        ]    ^___{    |    |]   }    '

Ambiguities

Lexical ambiguities are resolved according to the “longest match” rule: when a character sequence can be decomposed into two tokens in several different ways, the decomposition retained is the one with the longest first token.

2.2 Global names

Global names are used to denote value variables, value constructors (constant or non-constant), type constructors, and record labels. Internally, a global name consists of two parts: the name of the defining module (the module name), and the name of the global inside that module (the local name). The two parts of the name must be valid identifiers. Externally, global names have the following syntax:

global-name ::= ident
| ident __ident

The form ident __ident is called a qualified name. The first identifier is the module name, the second identifier is the local name. The form ident is called an unqualified name. The identifier is the local name; the module name is omitted. The compiler infers this module name following the completion rules given below, therefore transforming the unqualified name into a full global name.

To complete an unqualified identifier, the compiler checks a list of modules, the opened modules, to see if they define a global with the same local name as the unqualified identifier. When one is found, the identifier is completed into the full name of that global. That is, the compiler takes as module name the name of an opened module that defines a global with the same local name as the unqualified identifier. If several modules satisfy this condition, the one that comes first in the list of opened modules is selected.

The list of opened modules always includes the module currently being compiled (checked first). (In the case of a toplevel-based implementation, this is the module where all toplevel definitions are entered.) It also includes a number of standard library modules that provide the initial environment (checked last). In addition, the #open and #close directives can be used to add or remove modules from that list. The modules added with #open are checked after the module currently being compiled, but before the initial standard library modules.

variable ::= global-name
| prefix operator-name
operator-name ::= + | - | * | / | mod | +. | -. | *. | /.

                                  | @ | ^| ! | := | = | <> | == | != | !
                                  | < | <= | > | <= | <. | <=. | >. | <=.

cconstr ::= global-name

                                  | []
                                  | ()

ncconstr ::= global-name

                                  | prefix::

typeconstr ::= global-name
label ::= global-name

Depending on the context, global names can stand for global variables (variable), constant value constructors (cconstr), non-constant value constructors (ncconst), type constructors (typeconstr), or record labels (label). For variables and value constructors, special names built with prefix and an operator name are recognized. The tokens [] and () are also recognized as built-in constant constructors (the empty list and the unit value).

The syntax of the language restricts labels and type constructors to appear in certain positions, where no other kind of global names are accepted. Hence labels and type constructors have their own name spaces. Value constructors and value variables live in the same name space: a global name in value position is interpreted as a value constructor if it appears in the scope of a type declaration defining that constructor; otherwise, the global name is taken to be a value variable. For value constructors, the type declaration determines whether a constructor is constant or not.

2.3 Values

This section describes the kinds of values that are manipulated by Caml Light programs.

2.3.1 Base values

Integer numbers

Integer values are integer numbers from −2^{30} to 2^{30} − 1, that is −1073741824 to 1073741823. Implementations may support a wider range of integer values.

Floating-point numbers

Floating-point values are numbers in floating-point representation. Everything about floating-point values is implementation-dependent, including the range of representable numbers, the number of significant digits, and the way floating-point results are rounded.

Characters

Character values are represented as 8-bit integers between 0 and 255. Character codes between 0 and 127 are interpreted following the ASCII standard. The interpretation of character codes between 128 and 255 is implementation-dependent.

Character strings

String values are finite sequences of characters. Implementations must support strings up to 2^{16} −1 characters in length (65535 characters). Implementations may support longer strings.

2.3.2 Tuples

Tuples of values are written (v_{1}, …, v_{n}), standing for the n-tuple of values v_{1} to v_{n}. Tuples of up to 2^{14} − 1 elements (16383 elements) must be supported, though implementations may support tuples with more elements.

2.3.3 Records

Record values are labeled tuples of values. The record value written {label_{1} = v_{1}; …; label_{n} = v_{n}} associates the value v_{i} to the record label label_{i}, for i = 1 … n. Records with up to 2^{14} − 1 fields (16383 fields) must be supported, though implementations may support records with more fields.

2.3.4 Arrays

Arrays are finite, variable-sized sequences of values of the same type. Arrays of length up to 2^{14} −1 (16383 elements) must be supported, though implementations may support larger arrays.

2.3.5 Variant values

Variant values are either a constant constructor, or a pair of a non-constant constructor and a value. The former case is written cconstr; the latter case is written ncconstr(v), where v is said to be the argument of the non-constant constructor ncconstr.

The following constants are treated like built-in constant constructors:

Constant Constructor
false  the boolean false
true  the boolean true
()  the “unit” value
[]  the empty list

2.3.6 Functions

Functional values are mappings from values to values.

2.4 Type expressions

typexpr ::= ' ident
| ( typexpr )
| typexpr -> typexpr
| typexpr {* typexpr}^{+}
| typeconstr
| typexpr typeconstr
| ( typexpr {, typexpr} ) typeconstr

The table below shows the relative precedences and associativity of operators and non-closed type constructions. The constructions with higher precedences come first.

Operator  Associativity
Type constructor application –

                      *  –

->  right

Type expressions denote types in definitions of data types as well as in type constraints over patterns and expressions.

Type variables

The type expression ' ident stands for the type variable named ident. In data type definitions, type variables are names for the data type parameters. In type constraints, they represent unspecified types that can be instantiated by any type to satisfy the type constraint.

Parenthesized types

The type expression ( typexpr ) denotes the same type as typexpr.

Function types

The type expression typexpr_{1} -> typexpr_{2} denotes the type of functions mapping arguments of type typexpr_{1} to results of type typexpr_{2}.

Tuple types

The type expression typexpr_{1} * … * typexpr_{n} denotes the type of tuples whose elements belong to types typexpr_{1}, … typexpr_{n} respectively.

Constructed types

Type constructors with no parameter, as in typeconstr, are type expressions.

The type expression typexpr typeconstr, where typeconstr is a type constructor with one parameter, denotes the application of the unary type constructor typeconstr to the type typexpr.

The type expression (typexpr_{1}, …, typexpr_{n}) typeconstr, where typeconstr is a type constructor with n parameters, denotes the application of the n-ary type constructor typeconstr to the types typexpr_{1} through typexpr_{n}.

2.5 Constants

constant ::= integer-literal
| float-literal
| char-literal
| string-literal
| cconstr

The syntactic class of constants comprises literals from the four base types (integers, floatingpoint numbers, characters, character strings), and constant constructors.

2.6 Patterns

pattern ::= ident

| _

| pattern as ident
| ( pattern )
| ( pattern : typexpr )
| pattern | pattern
| constant
| ncconstr pattern
| pattern , pattern {, pattern}
| { label = pattern {; label = pattern} }

| []

| [ pattern {; pattern} ]
| pattern :: pattern

The table below shows the relative precedences and associativity of operators and non-closed pattern constructions. The constructions with higher precedences come first.

Operator  Associativity
Constructor application –
::  right

,  –

|  left

as  –

Patterns are templates that allow selecting data structures of a given shape, and binding identifiers to components of the data structure. This selection operation is called pattern matching; its outcome is either “this value does not match this pattern”, or “this value matches this pattern, resulting in the following bindings of identifiers to values”.

Variable patterns

A pattern that consists in an identifier matches any value, binding the identifier to the value. The pattern _also matches any value, but does not bind any identifier.

Alias patterns

The pattern pattern_{1} as ident matches the same values as pattern_{1}. If the matching against pattern_{1} is successful, the identifier ident is bound to the matched value, in addition to the bindings performed by the matching against pattern_{1}.

Parenthesized patterns

The pattern ( pattern_{1} ) matches the same values as pattern_{1}. A type constraint can appear in a parenthesized patterns, as in ( pattern_{1} : typexpr ). This constraint forces the type of pattern_{1} to be compatible with type.

“Or” patterns

The pattern pattern_{1} | pattern_{2} represents the logical “or” of the two patterns pattern_{1} and pattern_{2}. A value matches pattern_{1} | pattern_{2} either if it matches pattern_{1} or if it matches pattern_{2}. The two sub-patterns pattern_{1} and pattern_{2} must contain no identifiers. Hence no bindings are returned by matching against an “or” pattern.

Constant patterns

A pattern consisting in a constant matches the values that are equal to this constant.

Variant patterns

The pattern ncconstr pattern_{1} matches all variants whose constructor is equal to ncconstr, and whose argument matches pattern_{1}.

The pattern pattern_{1} :: pattern_{2} matches non-empty lists whose heads match pattern_{1}, and whose tails match pattern_{2}. This pattern behaves like prefix::( pattern_{1} , pattern_{2} ).

The pattern [ pattern_{1} ; …; pattern_{n} ] matches lists of length n whose elements match pattern_{1} … pattern_{n}, respectively. This pattern behaves like pattern_{1} :: …:: pattern_{n} ::[].

Tuple patterns

The pattern pattern_{1} , …, pattern_{n} matches n-tuples whose components match the patterns pattern_{1} through pattern_{n}. That is, the pattern matches the tuple values (v_{1}, …, v_{n}) such that pattern_{i} matches v_{i} for i = 1, …, n.

Record patterns

The pattern { label_{1} = pattern_{1} ; …; label_{n} = pattern_{n} } matches records that define at least the labels label_{1} through label_{n}, and such that the value associated to label_{i} match the pattern pattern_{i}, for i = 1, …, n. The record value can define more labels than label_{1} … label_{n}; the values associated to these extra labels are not taken into account for matching.

2.7 Expressions

expr ::= ident
| variable
| constant
| ( expr )

| begin expr end

| ( expr : typexpr )
| expr , expr {, expr}
| ncconstr expr
| expr :: expr
| [ expr {; expr} ]
| [| expr {; expr} |]
| { label = expr {; label = expr} }
| expr expr
| prefix-op expr
| expr infix-op expr
| expr . label
| expr . label <- expr
| expr .( expr )
| expr .( expr )<- expr
| expr & expr
| expr or expr
| if expr then expr [else expr]

| while expr do expr done

| for ident = expr (to | downto) expr do expr done
| expr ; expr
| match expr with simple-matching
| fun multiple-matching
| function simple-matching
| try expr with simple-matching
| let [rec] let-binding {and let-binding} in expr
simple-matching ::= pattern -> expr {| pattern -> expr}
multiple-matching ::= pattern-list -> expr {| pattern-list -> expr}
pattern-list ::= pattern {pattern}
let-binding ::= pattern = expr
| variable pattern-list = expr
prefix-op ::= - | -. | !
infix-op ::= + | - | * | / | mod | +. | -. | *. | /. | ** | @ | ^| ! | :=

| = | <> | == | != | < | <= | > | >= | <. | <=. | >. | >=.

The table below shows the relative precedences and associativity of operators and non-closed constructions. The constructions with higher precedence come first.

Construction or operator  Associativity

!  –
.   .(  –

function application  left
constructor application  –
-   -. (prefix)  –
**  right
mod  left

*   *.   /   /.  left
+   +.   -   -.  left

::  right
@ ^right
comparisons (=  ==  < etc.) left

not  –

&  left
or  left

,  –

<-   :=  right

if  –

;  right

let match fun function try –

2.7.1 Simple expressions

Constants

Expressions consisting in a constant evaluate to this constant.

Variables

Expressions consisting in a variable evaluate to the value bound to this variable in the current evaluation environment. The variable can be either a qualified identifier or a simple identifier. Qualified identifiers always denote global variables. Simple identifiers denote either a local variable, if the identifier is locally bound, or a global variable, whose full name is obtained by qualifying the simple identifier, as described in section 2.2.

Parenthesized expressions

The expressions ( expr ) and begin expr end have the same value as expr. Both constructs are semantically equivalent, but it is good style to use begin … end inside control structures:

if . . . then begin . . . ; . . . end else begin . . . ; . . . end

and ( …) for the other grouping situations.

Parenthesized expressions can contain a type constraint, as in ( expr : type ). This constraint forces the type of expr to be compatible with type.

Function abstraction

The most general form of function abstraction is:

fun pattern_{1}^{1} … pattern_{1}^{m} -> expr_{1}
| …
| pattern_{n}^{1} … pattern_{n}^{m} -> expr_{n}

This expression evaluates to a functional value with m curried arguments. When this function is applied to m values v_{1} … v_{m}, the values are matched against each pattern row pattern_{i}^{1} … pattern_{i}^{m} for i from 1 to n. If one of these matchings succeeds, that is if the value v_{j} matches the pattern pattern_{i}^{j} for all j = 1, …, m, then the expression expr_{i} associated to the selected pattern row is evaluated, and its value becomes the value of the function application. The evaluation of expr_{i} takes place in an environment enriched by the bindings performed during the matching.

If several pattern rows match the arguments, the one that occurs first in the function definition is selected. If none of the pattern rows matches the argument, the exception Match_failure is raised.

If the function above is applied to less than m arguments, a functional value is returned, that represents the partial application of the function to the arguments provided. This partial application is a function that, when applied to the remaining arguments, matches all arguments against the pattern rows as described above. Matching does not start until all m arguments have been provided to the function; hence, partial applications of the function to less than m arguments never raise

Match_failure.

All pattern rows in the function body must contain the same number of patterns. A variable must not be bound more than once in one pattern row.

Functions with only one argument can be defined with the function keyword instead of fun:

function pattern_{1} -> expr_{1}
| …
| pattern_{n} -> expr_{n}

The function thus defined behaves exactly as described above. The only difference between the two forms of function definition is how a parsing ambiguity is resolved. The two forms cconstr pattern (two patterns in a row) and ncconstr pattern (one pattern) cannot be distinguished syntactically. Function definitions introduced by fun resolve the ambiguity to the former form; function definitions introduced by function resolve it to the latter form (the former form makes no sense in this case).

Function application

Function application is denoted by juxtaposition of expressions. The expression expr_{1} expr_{2} … expr_{n} evaluates the expressions expr_{1} to expr_{n}. The expression expr_{1} must evaluate to a functional value, which is then applied to the values of expr_{2}, …, expr_{n}. The order in which the expressions expr_{1}, …, expr_{n} are evaluated is not specified.

Local definitions

The let and let rec constructs bind variables locally. The construct

let pattern_{1} = expr_{1} and … and pattern_{n} = expr_{n} in expr

evaluates expr_{1} … expr_{n} in some unspecified order, then matches their values against the patterns pattern_{1} … pattern_{n}. If the matchings succeed, expr is evaluated in the environment enriched by the bindings performed during matching, and the value of expr is returned as the value of the whole let expression. If one of the matchings fails, the exception Match_failure is raised.

An alternate syntax is provided to bind variables to functional values: instead of writing

ident =fun pattern_{1} … pattern_{m} -> expr

in a let expression, one may instead write

ident pattern_{1} … pattern_{m} = expr

Both forms bind ident to the curried function with m arguments and only one case,

pattern_{1} … pattern_{m} -> expr.

Recursive definitions of variables are introduced by let rec:

letrec pattern_{1} = expr_{1} and … and pattern_{n} = expr_{n} in expr

The only difference with the let construct described above is that the bindings of variables to values performed by the pattern-matching are considered already performed when the expressions expr_{1} to expr_{n} are evaluated. That is, the expressions expr_{1} to expr_{n} can reference identifiers that are bound by one of the patterns pattern_{1}, …, pattern_{n}, and expect them to have the same value as in expr, the body of the let rec construct.

The recursive definition is guaranteed to behave as described above if the expressions expr_{1} to expr_{n} are function definitions (fun … or function …), and the patterns pattern_{1} … pattern_{n} consist in a single variable, as in:

letrec ident_{1} =fun … and … and ident_{n} =fun … in expr

This defines ident_{1} … ident_{n} as mutually recursive functions local to expr. The behavior of other forms of let rec definitions is implementation-dependent.

2.7.2 Control constructs

Sequence

The expression expr_{1} ; expr_{2} evaluates expr_{1} first, then expr_{2}, and returns the value of expr_{2}.

Conditional

The expression if expr_{1} then expr_{2} else expr_{3} evaluates to the value of expr_{2} if expr_{1} evaluates to the boolean true, and to the value of expr_{3} if expr_{1} evaluates to the boolean false.

The else expr_{3} part can be omitted, in which case it defaults to else().

Case expression

The expression

                             match expr

with pattern_{1} -> expr_{1}
| …
| pattern_{n} -> expr_{n}

matches the value of expr against the patterns pattern_{1} to pattern_{n}. If the matching against pattern_{i} succeeds, the associated expression expr_{i} is evaluated, and its value becomes the value of the whole match expression. The evaluation of expr_{i} takes place in an environment enriched by the bindings performed during matching. If several patterns match the value of expr, the one that occurs first in the match expression is selected. If none of the patterns match the value of expr, the exception Match_failure is raised.

Boolean operators

The expression expr_{1} & expr_{2} evaluates to true if both expr_{1} and expr_{2} evaluate to true; otherwise, it evaluates to false. The first component, expr_{1}, is evaluated first. The second component, expr_{2}, is not evaluated if the first component evaluates to false. Hence, the expression expr_{1} & expr_{2} behaves exactly as

                           if expr_{1} then expr_{2} elsefalse.

The expression expr_{1} or expr_{2} evaluates to true if one of expr_{1} and expr_{2} evaluates to true; otherwise, it evaluates to false. The first component, expr_{1}, is evaluated first. The second component, expr_{2}, is not evaluated if the first component evaluates to true. Hence, the expression expr_{1} or expr_{2} behaves exactly as

                            if expr_{1} thentrueelse expr_{2}.

Loops

The expression while expr_{1} do expr_{2} done repeatedly evaluates expr_{2} while expr_{1} evaluates to true. The loop condition expr_{1} is evaluated and tested at the beginning of each iteration. The whole while … done expression evaluates to the unit value ().

The expression for ident = expr_{1} to expr_{2} do expr_{3} done first evaluates the expressions expr_{1} and expr_{2} (the boundaries) into integer values n and p. Then, the loop body expr_{3} is repeatedly evaluated in an environment where the local variable named ident is successively bound to the values n, n + 1, …, p − 1, p. The loop body is never evaluated if n > p.

The expression for ident = expr_{1} downto expr_{2} do expr_{3} done first evaluates the expressions expr_{1} and expr_{2} (the boundaries) into integer values n and p. Then, the loop body expr_{3} is repeatedly evaluated in an environment where the local variable named ident is successively bound to the values n, n − 1, …, p + 1, p. The loop body is never evaluated if n < p.

In both cases, the whole for expression evaluates to the unit value ().

Exception handling

The expression

try expr
with pattern_{1} -> expr_{1}
| …
| pattern_{n} -> expr_{n}

evaluates the expression expr and returns its value if the evaluation of expr does not raise any exception. If the evaluation of expr raises an exception, the exception value is matched against the patterns pattern_{1} to pattern_{n}. If the matching against pattern_{i} succeeds, the associated expression expr_{i} is evaluated, and its value becomes the value of the whole try expression. The evaluation of expr_{i} takes place in an environment enriched by the bindings performed during matching. If several patterns match the value of expr, the one that occurs first in the try expression is selected. If none of the patterns matches the value of expr, the exception value is raised again, thereby transparently “passing through” the try construct.

2.7.3 Operations on data structures

Products

The expression expr_{1} , …, expr_{n} evaluates to the n-tuple of the values of expressions expr_{1} to expr_{n}. The evaluation order for the subexpressions is not specified.

Variants

The expression ncconstr expr evaluates to the variant value whose constructor is ncconstr, and whose argument is the value of expr.

For lists, some syntactic sugar is provided. The expression expr_{1} :: expr_{2} stands for the constructor prefix :: applied to the argument ( expr_{1} , expr_{2} ), and therefore evaluates to the list whose head is the value of expr_{1} and whose tail is the value of expr_{2}. The expression [ expr_{1} ; …; expr_{n} ] is equivalent to expr_{1} :: …:: expr_{n} :: [], and therefore evaluates to the list whose elements are the values of expr_{1} to expr_{n}.

Records

The expression { label_{1} = expr_{1} ; …; label_{n} = expr_{n} } evaluates to the record value { label_{1} = v_{1} ; …; label_{n} = v_{n} }, where v_{i} is the value of expr_{i} for i = 1, …, n. The labels label_{1} to label_{n} must all belong to the same record types; all labels belonging to this record type must appear exactly once in the record expression, though they can appear in any order. The order in which expr_{1} to expr_{n} are evaluated is not specified.

The expression expr_{1} . label evaluates expr_{1} to a record value, and returns the value associated to label in this record value.

The expression expr_{1} . label <- expr_{2} evaluates expr_{1} to a record value, which is then modified in-place by replacing the value associated to label in this record by the value of expr_{2}. This operation is permitted only if label has been declared mutable in the definition of the record type. The whole expression expr_{1} . label <- expr_{2} evaluates to the unit value ().

Arrays

The expression [| expr_{1} ; …; expr_{n} |] evaluates to a n-element array, whose elements are initialized with the values of expr_{1} to expr_{n} respectively. The order in which these expressions are evaluated is unspecified.

The expression expr_{1} .( expr_{2} ) is equivalent to the application vect_item expr_{1} expr_{2}. In the initial environment, the identifier vect_item resolves to a built-in function that returns the value of element number expr_{2} in the array denoted by expr_{1}. The first element has number 0; the last element has number n − 1, where n is the size of the array. The exception Invalid_argument is raised if the access is out of bounds.

The expression expr_{1} .( expr_{2} ) <- expr_{3} is equivalent to vect_assign expr_{1} expr_{2} expr_{3}. In the initial environment, the identifier vect_assign resolves to a built-in function that modifies in-place the array denoted by expr_{1}, replacing element number expr_{2} by the value of expr_{3}. The exception Invalid_argument is raised if the access is out of bounds. The built-in function returns (). Hence, the whole expression expr_{1} .( expr_{2} )<- expr_{3} evaluates to the unit value ().

This behavior of the two constructs expr_{1} .( expr_{2} ) and expr_{1} .( expr_{2} )<- expr_{3} may change if the meaning of the identifiers vect_item and vect_assign is changed, either by redefinition or by modification of the list of opened modules. See the discussion below on operators.

2.7.4 Operators

The operators written infix-op in the grammar above can appear in infix position (between two expressions). The operators written prefix-op in the grammar above can appear in prefix position (in front of an expression).

The expression prefix-op expr is interpreted as the application ident expr, where ident is the identifier associated to the operator prefix-op in the table below. Similarly, the expression expr_{1} infix-op expr_{2} is interpreted as the application ident expr_{1} expr_{2}, where ident is the identifier associated to the operator infix-op in the table below. The identifiers written ident above are then evaluated following the rules in section 2.7.1. In the initial environment, they evaluate to built-in functions whose behavior is described in the table. The behavior of the constructions prefix-op expr and expr_{1} infix-op expr_{2} may change if the meaning of the identifiers associated to prefix-op or infix-op is changed, either by redefinition of the identifiers, or by modification of the list of opened modules, through the #open and #close directives.

Operator Associated Behavior in the default environment

identifier

+ prefix + Integer addition.

- (infix) prefix - Integer subtraction.

- (prefix) minus Integer negation.

* prefix * Integer multiplication. / prefix / Integer division. Raise Division_by_zero if second argu-

ment is zero. The result is unspecified if either argument is negative.

mod prefix mod Integer modulus. Raise Division_by_zero if second argument is zero. The result is unspecified if either argument is negative.

+. prefix +. Floating-point addition. -. (infix) prefix -. Floating-point subtraction. -. (prefix) minus_float Floating-point negation. *. prefix *. Floating-point multiplication. /. prefix /. Floating-point division. Raise Division_by_zero if second

argument is zero.

** prefix ** Floating-point exponentiation. @ prefix @ List concatenation. ^prefix ^String concatenation. ! prefix ! Dereferencing (return the current contents of a reference). := prefix := Reference assignment (update the reference given as first

argument with the value of the second argument).

= prefix = Structural equality test. <> prefix <> Structural inequality test. == prefix == Physical equality test. != prefix != Physical inequality test. < prefix < Test “less than” on integers. <= prefix <= Test “less than or equal ” on integers. > prefix > Test “greater than” on integers. >= prefix >= Test “greater than or equal” on integers. <. prefix <. Test “less than” on floating-point numbers. <=. prefix <=. Test “less than or equal ” on floating-point numbers. >. prefix >. Test “greater than” on floating-point numbers. >=. prefix >=. Test “greater than or equal” on floating-point numbers.

The behavior of the +, -, *, /, mod, +., -., *. or /. operators is unspecified if the result falls outside of the range of representable integers or floating-point numbers, respectively. See chapter 13 for a more precise description of the behavior of the operators above.

2.8 Global definitions

This section describes the constructs that bind global identifiers (value variables, value constructors, type constructors, record labels).

2.8.1 Type definitions

type-definition ::= type typedef {and typedef }
typedef ::= type-params ident = constr-decl {| constr-decl}
| type-params ident ={ label-decl {; label-decl} }
| type-params ident == typexpr
| type-params ident
type-params ::= nothing
| ' ident
| (' ident {,' ident} )
constr-decl ::= ident
| ident of typexpr
label-decl ::= ident : typexpr
| mutable ident : typexpr

Type definitions bind type constructors to data types: either variant types, record types, type abbreviations, or abstract data types.

Type definitions are introduced by the type keyword, and consist in one or several simple definitions, possibly mutually recursive, separated by the and keyword. Each simple definition defines one type constructor.

A simple definition consists in an identifier, possibly preceded by one or several type parameters, and followed by a data type description. The identifier is the local name of the type constructor being defined. (The module name for this type constructor is the name of the module being compiled.) The optional type parameters are either one type variable ' ident, for type constructors with one parameter, or a list of type variables (' ident_{1}, …, ' ident_{n}), for type constructors with several parameters. These type parameters can appear in the type expressions of the right-hand side of the definition.

Variant types

The type definition typeparams ident = constr-decl_{1} | … | constr-decl_{n} defines a variant type. The constructor declarations constr-decl_{1}, …, constr-decl_{n} describe the constructors associated to this variant type. The constructor declaration ident of typexpr declares the local name ident (in the module being compiled) as a non-constant constructor, whose argument has type typexpr. The constructor declaration ident declares the local name ident (in the module being compiled) as a constant constructor.

Record types

The type definition typeparams ident ={ label-decl_{1} ; …; label-decl_{n} } defines a record type. The label declarations label-decl_{1}, …, label-decl_{n} describe the labels associated to this record type. The label declaration ident : typexpr declares the local name ident in the module being compiled as a label, whose argument has type typexpr. The label declaration mutable ident : typexpr behaves similarly; in addition, it allows physical modification over the argument to this label.

Type abbreviations

The type definition typeparams ident == typexpr defines the type constructor ident as an abbreviation for the type expression typexpr.

Abstract types

The type definition typeparams ident defines ident as an abstract type. When appearing in a module interface, this definition allows exporting a type constructor while hiding how it is represented in the module implementation.

2.8.2 Exception definitions

exception-definition ::= exception constr-decl {and constr-decl}

Exception definitions add new constructors to the built-in variant type exn of exception values. The constructors are declared as for a definition of a variant type.

2.9 Directives

directive ::= #open string
| #close string
| # ident string

Directives control the behavior of the compiler. They apply to the remainder of the current compilation unit.

The two directives #open and #close modify the list of opened modules, that the compiler uses to complete unqualified identifiers, as described in section 2.2. The directive #open string adds the module whose name is given by the string constant string to the list of opened modules, in first position. The directive #close string removes the first occurrence of the module whose name is given by the string constant string from the list of opened modules.

Implementations can provide other directives, provided they follow the syntax # ident string, where ident is the name of the directive, and the string constant string is the argument to the directive. The behavior of these additional directives is implementation-dependent.

2.10 Module implementations

implementation ::= {impl-phrase ;;}
impl-phrase ::= expr
| value-definition
| type-definition
| exception-definition
| directive
value-definition ::= let [rec] let-binding {and let-binding}

A module implementation consists in a sequence of implementation phrases, terminated by double semicolons. An implementation phrase is either an expression, a value definition, a type or exception definition, or a directive. At run-time, implementation phrases are evaluated sequentially, in the order in which they appear in the module implementation.

Implementation phrases consisting in an expression are evaluated for their side-effects.

Value definitions bind global value variables in the same way as a let … in … expression binds local variables. The expressions are evaluated, and their values are matched against the left-hand sides of the = sides, as explained in section 2.7.1. If the matching succeeds, the bindings of identifiers to values performed during matching are interpreted as bindings to the global value variables whose local name is the identifier, and whose module name is the name of the module. If the matching fails, the exception Match_failure is raised. The scope of these bindings is the phrases that follow the value definition in the module implementation.

Type and exception definitions introduce type constructors, variant constructors and record labels as described in sections 2.8.1 and 2.8.2. The scope of these definitions is the phrases that follow the value definition in the module implementation. The evaluation of an implementation phrase consisting in a type or exception definition produces no effect at run-time.

Directives modify the behavior of the compiler on the subsequent phrases of the module implementation, as described in section 2.9. The evaluation of an implementation phrase consisting in a directive produces no effect at run-time. Directives apply only to the module currently being compiled; in particular, they have no effect on other modules that refer to globals exported by the module being compiled.

2.11 Module interfaces

interface ::= {intf-phrase ;;}
intf-phrase ::= value-declaration
| type-definition
| exception-definition
| directive
value-declaration ::= value ident : typexpr {and ident : typexpr}

Module interfaces declare the global objects (value variables, type constructors, variant constructors, record labels) that a module exports, that is, makes available to other modules. Other modules can refer to these globals using qualified identifiers or the #open directive, as explained in section 2.2.

A module interface consists in a sequence of interface phrases, terminated by double semicolons. An interface phrase is either a value declaration, a type definition, an exception definition, or a directive.

Value declarations declare global value variables that are exported by the module implementation, and the types with which they are exported. The module implementation must define these variables, with types at least as general as the types declared in the interface. The scope of the bindings for these global variables extends from the module implementation itself to all modules that refer to those variables.

Type or exception definitions introduce type constructors, variant constructors and record labels as described in sections 2.8.1 and 2.8.2. Exception definitions and type definitions that are not abstract type declarations also take effect in the module implementation; that is, the type constructors, variant constructors and record labels they define are considered bound on entrance to the module implementation, and can be referred to by the implementation phrases. Type definitions that are not abstract type declarations must not be redefined in the module implementation. In contrast, the type constructors that are declared abstract in a module interface must be defined in the module implementation, with the same names.

Directives modify the behavior of the compiler on the subsequent phrases of the module interface, as described in section 2.9. Directives apply only to the interface currently being compiled; in particular, they have no effect on other modules that refer to globals exported by the interface being compiled.

34

Language extensions

This chapter describes the language features that are implemented in Caml Light, but not described in the Caml Light reference manual. In contrast with the fairly stable kernel language that is described in the reference manual, the extensions presented here are still experimental, and may be removed or changed in the future.

3.1 Streams, parsers, and printers

Caml Light comprises a built-in type for streams (possibly infinite sequences of elements, that are evaluated on demand), and associated stream expressions, to build streams, and stream patterns, to destructure streams. Streams and stream patterns provide a natural approach to the writing of recursive-descent parsers.

Streams are presented by the following extensions to the syntactic classes of expressions:

expr ::= …

                            | [<>]

| [< stream-component {; stream-component} >]
| function stream-matching
| match expr with stream-matching
stream-component ::= ' expr
| expr
stream-matching ::= stream-pattern -> expr {| stream-pattern -> expr}
stream-pattern ::= [<>]
| [< stream-comp-pat {; stream-comp-pat} >]
stream-comp-pat ::= ' pattern
| expr pattern
| ident

Stream expressions are bracketed by [< and >]. They represent the concatenation of their components. The component ' expr represents the one-element stream whose element is the value of expr. The component expr represents a sub-stream. For instance, if both s and t are streams of integers, then [<'1; s; t; '2>] is a stream of integers containing the element 1, then the elements of s, then those of t, and finally 2. The empty stream is denoted by [< >].

Unlike any other kind of expressions in the language, stream expressions are submitted to lazy evaluation: the components are not evaluated when the stream is built, but only when they are accessed during stream matching. The components are evaluated once, the first time they are accessed; the following accesses reuse the value computed the first time.

Stream patterns, also bracketed by [< and >], describe initial segments of streams. In particular, the stream pattern [< >] matches all streams. Stream pattern components are matched against the corresponding elements of a stream. The component ' pattern matches the corresponding stream element against the pattern. The component expr pattern applies the function denoted by expr to the current stream, then matches the result of the function against pattern. Finally, the component ident simply binds the identifier to the stream being matched. (The current implementation limits ident to appear last in a stream pattern.)

Stream matching proceeds destructively: once a component has been matched, it is discarded from the stream (by in-place modification).

Stream matching proceeds in two steps: first, a pattern is selected by matching the stream against the first components of the stream patterns; then, the following components of the selected pattern are checked against the stream. If the following components do not match, the exception Parse_error is raised. There is no backtracking here: stream matching commits to the pattern selected according to the first element. If none of the first components of the stream patterns match, the exception Parse_failure is raised. The Parse_failure exception causes the next alternative to be tried, if it occurs during the matching of the first element of a stream, before matching has committed to one pattern.

See Functional programming using Caml Light for a more gentle introductions to streams, and for some examples of their use in writing parsers. A more formal presentation of streams, and a discussion of alternate semantics, can be found in Parsers in ML by Michel Mauny and Daniel de Rauglaudre, in the proceedings of the 1992 ACM conference on Lisp and Functional Programming.

3.2 Guards

Cases of a pattern matching can include guard expressions, which are arbitrary boolean expressions that must evaluate to true for the match case to be selected. Guards occur just before the -> token and are introduced by the when keyword:

                        match expr

with pattern_{1}[whencond_{1}] -> expr_{1}
| …
| pattern_{n}[whencond_{n}] -> expr_{n}

(Same syntax for the fun, function, and try … with constructs.) During matching, if the value of expr matches some pattern pattern_{i} which has a guard cond_{i}, then the expression cond_{i} is evaluated (in an environment enriched by the bindings performed during matching). If cond_{i} evaluates to true, then expr_{i} is evaluated and its value returned as the result of the matching, as usual. But if cond_{i} evaluates to false, the matching is resumed against the patterns following pattern_{i}.

3.3 Range patterns

In patterns, Caml Light recognizes the form ` c ` .. ` d ` (two character constants separated by ..) as a shorthand for the pattern

                       ` c `|` c_{1} `|` c_{2} `| . . . |` c_{n} `|` d `

where c_{1}, c_{2}, …, c_{n} are the characters that occur between c and d in the ASCII character set. For instance, the pattern `0`..`9` matches all characters that are digits.

3.4 Recursive definitions of values

Besides let rec definitions of functional values, as described in the reference manual, Caml Light supports a certain class of recursive definitions of non-functional values. For instance, the following definition is accepted:

        let rec x = 1 :: y and y = 2 :: x;;

and correctly binds x to the cyclic list 1::2::1::2::…, and y to the cyclic list 2::1::2::1::… Informally, the class of accepted definitions consists of those definitions where the defined variables occur only inside function bodies or as a field of a data structure. Moreover, the patterns in the left-hand sides must be identifiers, nothing more complex.

3.5 Local definitions using where

A postfix syntax for local definitions is provided:

expr ::= …
| expr where [rec] let-binding

The expression expr where let-binding behaves exactly as let let-binding in expr, and similarly

for where rec and let rec.

3.6 Mutable variant types

The argument of a value constructor can be declared “mutable” when the variant type is defined:

        type foo = A of mutable int
                 | B of mutable int * int
                 | ...

This allows in-place modification of the argument part of a constructed value. Modification is performed by a new kind of expressions, written ident <- expr, where ident is an identifier bound by pattern-matching to the argument of a mutable constructor, and expr denotes the value that must be stored in place of that argument. Continuing the example above:

        let x = A 1 in
          begin match x with A y -> y <- 2 | _-> () end;
          x

returns the value A 2. The notation ident <- expr works also if ident is an identifier bound by pattern-matching to the value of a mutable field in a record. For instance,

        type bar = {mutable lbl : int};;
        let x = {lbl = 1} in
          begin match x with {lbl = y} -> y <- 2 end;
          x

returns the value {lbl = 2}.

3.7 String access

Extra syntactic constructs are provided to access and modify characters in strings:

expr ::= …
| expr .[ expr ]
| expr .[ expr ]<- expr

The expression expr_{1} .[ expr_{2} ] is equivalent to the application nth_char expr_{1} expr_{2}. In the initial environment, the identifier nth_char resolves to a built-in function that returns the character number expr_{2} in the string denoted by expr_{1}. The first element has number 0; the last element has number n − 1, where n is the length of the string. The exception Invalid_argument is raised if the access is out of bounds.

The expression expr_{1} .[ expr_{2} ]<- expr_{3} is equivalent to set_nth_char expr_{1} expr_{2} expr_{3}. In the initial environment, the identifier set_nth_char resolves to a built-in function that modifies in-place the string denoted by expr_{1}, replacing character number expr_{2} by the value of expr_{3}. The exception Invalid_argument is raised if the access is out of bounds. The built-in function returns

().

3.8 Alternate syntax

The syntax of some constructs has been slightly relaxed:

• An optional ; may terminate a sequence, list expression, or record expression. For instance, begin e_{1} ; e_{2} ;end is syntactically correct and synonymous with begin e_{1} ; e_{2} end.

• Similarly, an optional | may begin a pattern-matching expression. For instance, function | pat_{1} -> expr_{1} | … is syntactically correct and synonymous with function pat_{1} -> expr_{1} | . . ..

• The tokens && and || are recognized as synonymous for & (sequential “and”) and or (sequential “or”), respectively.

3.9 Infix symbols

Sequences of “operator characters”, such as <=> or !!, are read as a single token from the infix-symbol or prefix-symbol class:

infix-symbol ::= (= | < | > | @ | ^| | | & | ~ | + | - | * | / | $ | %) {operator-char}
prefix-symbol ::= (! | ?) {operator-char}
operator-char ::= ! | $ | % | & | * | + | - | . | / | : | ; | < | = | > | ? | @ | ^| | | ~

Tokens from these two classes generalize the built-in infix and prefix operators described in chapter 2:

expr ::= …
| prefix-symbol expr
| expr infix-symbol expr
variable ::= …
| prefix prefix-symbol
| prefix infix-symbol

No #infix directive (section 3.10) is needed to give infix symbols their infix status. The precedences and associativities of infix symbols in expressions are determined by their first character(s): symbols beginning with ** have highest precedence (exponentiation), followed by symbols beginning with *, / or % (multiplication), then + and - (addition), then @ and ^(concatenation), then all others symbols (comparisons). The updated precedence table for expressions is shown below. We write “*… ” to mean “any infix symbol starting with *”.

Construction or operator  Associativity
!… ?…  –

.   .(   .[  –

function application  left
constructor application  –
-   -. (prefix)  –
**…  right
*… /… %… mod  left
+… -…  left
::  right
@… ^…  right
comparisons (=  ==  < etc.), all other infix symbols left

not  –

&   &&  left

or  ||  left
,  –

<-   :=  right

if  –

;  right

let  match  fun  function  try  –

Some infix and prefix symbols are predefined in the default environment (see chapters 2 and 13 for a description of their behavior). The others are initially unbound and must be bound before use, with a letprefix infix-symbol = expr or letprefix prefix-symbol = expr binding.

3.10 Directives

In addition to the standard #open and #close directives, Caml Light provides three additional directives.

#infix" id "

Change the lexical status of the identifier id: in the remainder of the compilation unit, id is recognized as an infix operator, just like +. The notation prefix id can be used to refer to the identifier id itself. Expressions of the form expr_{1} id expr_{2} are parsed as the application prefix id expr_{1} expr_{2}. The argument to the #infix directive must be an identifier, that is, a sequence of letters, digits and underscores starting with a letter; otherwise, the #infix declaration has no effect. Example:

 #infix "union";;
 let prefix union = fun x y -> ... ;;
 [1,2] union [3,4];;

#uninfix" id "

Remove the infix status attached to the identifier id by a previous #infix" id " directive.

#directory" dir-name "

Add the named directory to the path of directories searched for compiled module interface files. This is equivalent to the -I command-line option to the batch compiler and the toplevel system.

42

Part III

The Caml Light commands Batch compilation (camlc)

This chapter describes how Caml Light programs can be compiled non-interactively, and turned into standalone executable files. This is achieved by the command camlc, which compiles and links Caml Light source files.

Mac: This command is not a standalone Macintosh application. To run camlc, you need the Macintosh Programmer’s Workshop (MPW) programming environment. The programs generated by camlc are also MPW tools, not standalone Macintosh applications.

4.1 Overview of the compiler

The camlc command has a command-line interface similar to the one of most C compilers. It accepts several types of arguments: source files for module implementations; source files for module interfaces; and compiled module implementations.

• Arguments ending in .mli are taken to be source files for module interfaces. Module interfaces declare exported global identifiers, define public data types, and so on. From the file x.mli, the camlc compiler produces a compiled interface in the file x.zi.

• Arguments ending in .ml are taken to be source files for module implementation. Module implementations bind global identifiers to values, define private data types, and contain expressions to be evaluated for their side-effects. From the file x.ml, the camlc compiler produces compiled object code in the file x.zo. If the interface file x.mli exists, the module implementation x.ml is checked against the corresponding compiled interface x.zi, which is assumed to exist. If no interface x.mli is provided, the compilation of x.ml produces a compiled interface file x.zi in addition to the compiled object code file x.zo. The file x.zi produced corresponds to an interface that exports everything that is defined in the implementation x.ml.

• Arguments ending in .zo are taken to be compiled object code. These files are linked together, along with the object code files obtained by compiling .ml arguments (if any), and the Caml Light standard library, to produce a standalone executable program. The order in which .zo and .ml arguments are presented on the command line is relevant: global identifiers are initialized in that order at run-time, and it is a link-time error to use a global identifier before having initialized it. Hence, a given x.zo file must come before all .zo files that refer to identifiers defined in the file x.zo.

The output of the linking phase is a file containing compiled code that can be executed by the Caml Light runtime system: the command named camlrun. If caml.out is the name of the file produced by the linking phase, the command

   camlrun caml.out arg_{1} arg_{2} . . . arg_{n}

executes the compiled code contained in caml.out, passing it as arguments the character strings arg_{1} to arg_{n}. (See chapter 6 for more details.)

Unix: On most Unix systems, the file produced by the linking phase can be run directly, as in:

./caml.out arg_{1} arg_{2} … arg_{n}

The produced file has the executable bit set, and it manages to launch the bytecode interpreter by itself.

PC: The output file produced by the linking phase is directly executable, provided it is given extension .EXE. Hence, if the output file is named caml_out.exe, you can execute it with the command

caml_out arg_{1} arg_{2} … arg_{n}

Actually, the produced file caml_out.exe is a tiny executable file prepended to the bytecode file. The executable simply runs the camlrun runtime system on the remainder of the file. (As a consequence, this is not a standalone executable: it still requires camlrun.exe to reside in one of the directories in the path.)

4.2 Options

The following command-line options are recognized by camlc.

-c Compile only. Suppress the linking phase of the compilation. Source code files are turned into compiled files, but no executable file is produced. This option is useful to compile modules separately.

-ccopt option

Pass the given option to the C compiler and linker, when linking in “custom runtime” mode (see the -custom option). For instance, -ccopt -Ldir causes the C linker to search for C libraries in directory dir.

-custom

Link in “custom runtime” mode. In the default linking mode, the linker produces bytecode that is intended to be executed with the shared runtime system, camlrun. In the custom runtime mode, the linker produces an output file that contains both the runtime system and the bytecode for the program. The resulting file is considerably larger, but it can be executed directly, even if the camlrun command is not installed. Moreover, the “custom runtime” mode enables linking Caml Light code with user-defined C functions, as described in chapter 12.

Unix: Never strip an executable produced with the -custom option.

PC: This option requires the DJGPP port of the GNU C compiler to be installed.

-g Cause the compiler to produce additional debugging information. During the linking phase, this option add information at the end of the executable bytecode file produced. This information is required by the debugger camldebug and also by the catch-all exception handler from the standard library module printexc.

During the compilation of an implementation file (.ml file), when the -g option is set, the compiler adds debugging information to the .zo file. It also writes a .zix file that describes the full interface of the .ml file, that is, all types and values defined in the .ml file, including those that are local to the .ml file (i.e. not declared in the .mli interface file). Used in conjunction with the -g option to the toplevel system (chapter 5), the .zix file gives access to the local values of the module, making it possible to print or “trace” them. The .zix file is not produced if the implementation file has no explicit interface, since, in this case, the module has no local values.

-i Cause the compiler to print the declared types, exceptions, and global variables (with their inferred types) when compiling an implementation (.ml file). This can be useful to check the types inferred by the compiler. Also, since the output follows the syntax of module interfaces, it can help in writing an explicit interface (.mli file) for a file: just redirect the standard output of the compiler to a .mli file, and edit that file to remove all declarations of unexported globals.

-I directory Add the given directory to the list of directories searched for compiled interface files (.zi) and compiled object code files (.zo). By default, the current directory is searched first, then the standard library directory. Directories added with -I are searched after the current directory, but before the standard library directory. When several directories are added with several -I options on the command line, these directories are searched from right to left (the rightmost directory is searched first, the leftmost is searched last). (Directories can also be added to the search path from inside the programs with the #directory directive; see chapter 3.)

-lang language-code Translate the compiler messages to the specified language. The language-code is fr for French, es for Spanish, de for German, … (See the file camlmsgs.txt in the Caml Light standard library directory for a list of available languages.) When an unknown language is specified, or no translation is available for a message, American English is used by default.

-o exec-file Specify the name of the output file produced by the linker.

Unix: The default output name is a.out, in keeping with the tradition.

PC: The default output name is caml_out.exe.

Mac: The default output name is Caml.Out.

-O module-set Specify which set of standard modules is to be implicitly “opened” at the beginning of a compilation. There are three module sets currently available:

cautious provides the standard operations on integers, floating-point numbers, characters, strings, arrays, …, as well as exception handling, basic input/output, etc. Operations from the cautious set perform range and bound checking on string and array operations, as well as various sanity checks on their arguments.

fast provides the same operations as the cautious set, but without sanity checks on their arguments. Programs compiled with -O fast are therefore slightly faster, but unsafe.

none suppresses all automatic opening of modules. Compilation starts in an almost empty environment. This option is not of general use, except to compile the standard library itself.

The default compilation mode is -O cautious. See chapter 13 for a complete listing of the modules in the cautious and fast sets.

-p Compile and link in profiling mode. See the description of the profiler camlpro in chapter 10.

-v Print the version number of the compiler.

-W Print extra warning messages for the following events:

• A #open directive is useless (no identifier in the opened module is ever referenced).

• A variable name in a pattern matching is capitalized (often corresponds to a misspelled constant constructor).

Unix: The following environment variable is also consulted:

LANG

When set, control which language is used to print the compiler messages (see the -lang command-line option).

PC: The following option is also supported:

@response-file Process the files whose names are listed in file response-file, just as if these names appeared on the command line. File names in response-file are separated by blanks (spaces, tabs, newlines). This option allows to overcome silly limitations on the length of the command line.

The following environment variables are also consulted:

CAMLLIB

Contain the path to the standard library directory.

      LANG

When set, control which language is used to print the compiler messages (see the -lang command-line option).

4.3 Modules and the file system

This short section is intended to clarify the relationship between the names of the modules and the names of the files that contain their compiled interface and compiled implementation.

The compiler always derives the name of the compiled module by taking the base name of the source file (.ml or .mli file). That is, it strips the leading directory name, if any, as well as the .ml or .mli suffix. The produced .zi and .zo files have the same base name as the source file; hence, the compiled files produced by the compiler always have their base name equal to the name of the module they describe (for .zi files) or implement (for .zo files).

For compiled interface files (.zi files), this invariant must be preserved at all times, since the compiler relies on it to load the compiled interface file for the modules that are used from the module being compiled. Hence, it is risky and generally incorrect to rename .zi files. It is admissible to move them to another directory, if their base name is preserved, and the correct -I options are given to the compiler.

Compiled bytecode files (.zo files), on the other hand, can be freely renamed once created. That’s because 1- .zo files contain the true name of the module they define, so there is no need to derive that name from the file name; 2- the linker never attempts to find by itself the .zo file that implements a module of a given name: it relies on the user providing the list of .zo files by hand.

4.4 Common errors

This section describes and explains the most frequently encountered error messages.

Cannot find file filename The named file could not be found in the current directory, nor in the directories of the search path. The filename is either a compiled interface file (.zi file), or a compiled bytecode file (.zo file). If filename has the format mod.zi, this means you are trying to compile a file that references identifiers from module mod, but you have not yet compiled an interface for module mod. Fix: compile mod.mli or mod.ml first, to create the compiled interface mod.zi.

If filename has the format mod.zo, this means you are trying to link a bytecode object file that does not exist yet. Fix: compile mod.ml first.

If your program spans several directories, this error can also appear because you haven’t specified the directories to look into. Fix: add the correct -I options to the command line.

Corrupted compiled interface file filename The compiler produces this error when it tries to read a compiled interface file (.zi file) that has the wrong structure. This means something went wrong when this .zi file was written: the disk was full, the compiler was interrupted in the middle of the file creation, and so on. This error can also appear if a .zi file is modified after its creation by the compiler. Fix: remove the corrupted .zi file, and rebuild it.

This expression has type t_{1}, but is used with type t_{2} This is by far the most common type error in programs. Type t_{1} is the type inferred for the expression (the part of the program that is displayed in the error message), by looking at the expression itself. Type t_{2} is the type expected by the context of the expression; it is deduced by looking at how the value of this expression is used in the rest of the program. If the two types t_{1} and t_{2} are not compatible, then the error above is produced.

In some cases, it is hard to understand why the two types t_{1} and t_{2} are incompatible. For instance, the compiler can report that “expression of type foo cannot be used with type foo”, and it really seems that the two types foo are compatible. This is not always true. Two type constructors can have the same name, but actually represent different types. This can happen if a type constructor is redefined. Example:

        type foo = A | B;;
        let f = function A -> 0 | B -> 1;;
        type foo = C | D;;
        f C;;

This result in the error message “expression C of type foo cannot be used with type foo”.

Incompatible types with the same names can also appear when a module is changed and recompiled, but some of its clients are not recompiled. That’s because type constructors in .zi files are not represented by their name (that would not suffice to identify them, because of type redefinitions), but by unique stamps that are assigned when the type declaration is compiled. Consider the three modules:

        mod1.ml:    type t = A | B;;
                    let f = function A -> 0 | B -> 1;;
        mod2.ml:    let g x = 1 + mod1__f(x);;
        mod3.ml:    mod2__g mod1__A;;

Now, assume mod1.ml is changed and recompiled, but mod2.ml is not recompiled. The recompilation of mod1.ml can change the stamp assigned to type t. But the interface mod2.zi will still use the old stamp for mod1__t in the type of mod2__g. Hence, when compiling mod3.ml, the system complains that the argument type of mod2__g (that is, mod1__t with the old stamp) is not compatible with the type of mod1__A (that is, mod1__t with the new stamp). Fix: use make or a similar tool to ensure that all clients of a module mod are recompiled when the interface mod.zi changes. To check that the Makefile contains the right dependencies, remove all .zi files and rebuild the whole program; if no “Cannot find file” error appears, you’re all set.

The type inferred for name, that is, t, contains non-generalizable type variables Type variables ('a, 'b, …) in a type t can be in either of two states: generalized (which means that the type t is valid for all possible instantiations of the variables) and not generalized (which means that the type t is valid only for one instantiation of the variables). In a let binding let name = expr, the type-checker normally generalizes as many type variables as possible in the type of expr. However, this leads to unsoundness (a well-typed program can crash) in conjunction with polymorphic mutable data structures. To avoid this, generalization is performed at let bindings only if the bound expression expr belongs to the class of “syntactic values”, which includes constants, identifiers, functions, tuples of syntactic values, etc. In all other cases (for instance, expr is a function application), a polymorphic mutable could have been created and generalization is therefore turned off.

Non-generalized type variables in a type cause no difficulties inside a given compilation unit (the contents of a .ml file, or an interactive session), but they cannot be allowed in types written in a .zi compiled interface file, because they could be used inconsistently in other compilation units. Therefore, the compiler flags an error when a .ml implementation without a .mli interface defines a global variable name whose type contains non-generalized type variables. There are two solutions to this problem:

• Add a type constraint or a .mli interface to give a monomorphic type (without type variables) to name. For instance, instead of writing

        let sort_int_list = sort (prefix <);;
        (* inferred type 'a list -> 'a list, with 'a not generalized *)

write

        let sort_int_list = (sort (prefix <) : int list -> int list);;

• If you really need name to have a polymorphic type, turn its defining expression into a function by adding an extra parameter. For instance, instead of writing

        let map_length = map vect_length;;
        (* inferred type 'a vect list -> int list, with 'a not generalized *)

write

            let map_length lv = map vect_length lv;;

mod__name is referenced before being defined This error appears when trying to link an incomplete or incorrectly ordered set of files. Either you have forgotten to provide an implementation for the module named mod on the command line (typically, the file named mod.zo, or a library containing that file). Fix: add the missing .ml or .zo file to the command line. Or, you have provided an implementation for the module named mod, but it comes too late on the command line: the implementation of mod must come before all bytecode object files that reference one of the global variables defined in module mod. Fix: change the order of .ml and .zo files on the command line.

Of course, you will always encounter this error if you have mutually recursive functions across modules. That is, function mod1__f calls function mod2__g, and function mod2__g calls function mod1__f. In this case, no matter what permutations you perform on the command line, the program will be rejected at link-time. Fixes:

• Put f and g in the same module.

• Parameterize one function by the other. That is, instead of having

mod1.ml:    let f x = ... mod2__g ... ;;
mod2.ml:    let g y = ... mod1__f ... ;;

define

mod1.ml:    let f g x = ... g ... ;;
mod2.ml:    let rec g y = ... mod1__f g ... ;;

and link mod1 before mod2.

• Use a reference to hold one of the two functions, as in :

mod1.ml:    let forward_g =
                ref((fun x -> failwith "forward_g") : <type>);;
            let f x = ... !forward_g ... ;;
mod2.ml:    let g y = ... mod1__f ... ;;
            mod1__forward_g := g;;

Unavailable C primitive f This error appears when trying to link code that calls external functions written in C in “default runtime” mode. As explained in chapter 12, such code must be linked in “custom runtime” mode. Fix: add the -custom option, as well as the (native code) libraries and (native code) object files that implement the required external functions.

The toplevel system (camllight)

This chapter describes the toplevel system for Caml Light, that permits interactive use of the Caml Light system, through a read-eval-print loop. In this mode, the system repeatedly reads Caml Light phrases from the input, then typechecks, compile and evaluate them, then prints the inferred type and result value, if any. The system prints a # (sharp) prompt before reading each phrase. A phrase can span several lines. Phrases are delimited by ;; (the final double-semicolon).

From the standpoint of the module system, all phrases entered at toplevel are treated as the implementation of a module named top. Hence, all toplevel definitions are entered in the module

top.

Unix: The toplevel system is started by the command camllight. Phrases are read on standard input, results are printed on standard output, errors on standard error. End-of-file on standard input terminates camllight (see also the quit system function below). The toplevel system does not perform line editing, but it can easily be used in conjunction with an external line editor such as fep; just run fep -emacs camllight or fep -vi camllight. Another option is to use camllight under Gnu Emacs, which gives the full editing power of Emacs (see the directory contrib/camlmode in the distribution). At any point, the parsing, compilation or evaluation of the current phrase can be interrupted by pressing ctrl-C (or, more precisely, by sending the intr signal to the camllight process). This goes back to the # prompt.

Mac: The toplevel system is presented as the standalone Macintosh application Caml Light. This application does not require the Macintosh Programmer’s Workshop to run. Once launched from the Finder, the application opens two windows, “Caml Light Input” and “Caml Light Output”. Phrases are entered in the “Caml Light Input” window. The “Caml Light Output” window displays a copy of the input phrases as they are processed by the Caml Light toplevel, interspersed with the toplevel responses. The “Return” key sends the contents of the Input window to the Caml Light toplevel. The “Enter” key inserts a newline without sending the contents of the Input window. (This can be configured with the “Preferences” menu item.) The contents of the input window can be edited at all times, with the standard Macintosh interface. An history of previously entered phrases is maintained, and can be accessed with the “Previous entry” (command-P) and “Next entry” (command-N) menu items.

To quit the Caml Light application, either select “Quit” from the “Files” menu, or use the quit function described below.

At any point, the parsing, compilation or evaluation of the current phrase can be interrupted by pressing “command-period”, or by selecting the item “Interrupt Caml Light” in the “Caml Light” menu. This goes back to the # prompt.

PC: The toplevel system is presented as a Windows application named Camlwin.exe. It should be launched from the Windows file manager or program manager.

The “Terminal” windows is split in two panes. Phrases are entered and edited in the bottom pane. The top pane displays a copy of the input phrases as they are processed by the Caml Light toplevel, interspersed with the toplevel responses. The “Return” key sends the contents of the bottom pane to the Caml Light toplevel. The “Enter” key inserts a newline without sending the contents of the Input window. (This can be configured with the “Preferences” menu item.)

The contents of the input window can be edited at all times, with the standard Windows interface. An history of previously entered phrases is maintained and displayed in a separate window.

To quit the Camlwin application, either select “Quit” from the “File” menu, or use the quit function described below.

At any point, the parsing, compilation or evaluation of the current phrase can be interrupted by selecting the “Interrupt Caml Light” menu item. This goes back to the # prompt.

A text-only version of the toplevel system is available under the name caml.exe. It runs under MSDOS as well as under Windows in a DOS window. No editing facilities are provided.

5.1 Options

The following command-line options are recognized by the caml or camllight commands.

-g Start the toplevel system in debugging mode. This mode gives access to values and types that are local to a module, that is, not exported by the interface of the module. When debugging mode is off, these local objects are not accessible (attempts to access them produce an “Unbound identifier” error). When debugging mode is on, these objects become visible, just like the objects that are exported in the module interface. In particular, values of abstract types are printed using their concrete representations, and the functions local to a module can be “traced” (see the trace function in section 5.2). This applies only to the modules that have been compiled in debugging mode (either by the batch compiler with the -g option, or by the toplevel system in debugging mode), that is, those modules that have an associated

.zix file.

-I directory Add the given directory to the list of directories searched for compiled interface files (.zi) and compiled object code files (.zo). By default, the current directory is searched first, then the standard library directory. Directories added with -I are searched after the current directory, but before the standard library directory. When several directories are added with several -I options on the command line, these directories are searched from right to left (the rightmost directory is searched first, the leftmost is searched last). Directories can also be added to the search path once the toplevel is running with the #directory directive; see chapter 3.

-lang language-code Translate the toplevel messages to the specified language. The language-code is fr for French, es for Spanish, de for German, … (See the file camlmsgs.txt in the Caml Light standard library directory for a list of available languages.) When an unknown language is specified, or no translation is available for a message, American English is used by default.

-O module-set Specify which set of standard modules is to be implicitly “opened” when the toplevel starts. There are three module sets currently available:

cautious provides the standard operations on integers, floating-point numbers, characters, strings, arrays, …, as well as exception handling, basic input/output, … Operations from the cautious set perform range and bound checking on string and vector operations, as well as various sanity checks on their arguments.

fast provides the same operations as the cautious set, but without sanity checks on their arguments. Programs compiled with -O fast are therefore slightly faster, but unsafe.

none suppresses all automatic opening of modules. Compilation starts in an almost empty environment. This option is not of general use.

The default compilation mode is -O cautious. See chapter 13 for a complete listing of the modules in the cautious and fast sets.

Unix: The following environment variables are also consulted:

 LANG

When set, control which language is used to print the compiler messages (see the -lang command-line option).

 LC_CTYPE

If set to iso_8859_1, accented characters (from the ISO Latin-1 character set) in string and character literals are printed as is; otherwise, they are printed as decimal escape sequences (\ddd).

5.2 Toplevel control functions

The standard library module toplevel, opened by default when the toplevel system is launched, provides a number of functions that control the toplevel behavior, load files in memory, and trace program execution.

value quit : unit -> unit

Exit the toplevel loop and terminate the camllight command.

value include : string -> unit

Read, compile and execute source phrases from the given file. The .ml extension is automatically added to the file name, if not present. This is textual inclusion: phrases are processed just as if they were typed on standard input. In particular, global identifiers defined by these phrases are entered in the module named top, not in a new module.

value load : string -> unit

Load in memory the source code for a module implementation. Read, compile and execute source phrases from the given file. The .ml extension is automatically added if not present. The load function behaves much like include, except that a new module is created, with name the base name of the source file name. Global identifiers defined in the source file are entered in this module, instead of the top module as in the case of include. For instance, assuming file foo.ml contains the single phrase

           let bar = 1;;

executing load "foo" defines the identifier foo__bar with value 1. Caveat: the loaded module is not automatically opened: the identifier bar does not automatically complete to foo__bar. To achieve this, you must execute the directive #open "foo" afterwards.

value compile : string -> unit

Compile the source code for a module implementation or interface (.ml or .mli file). Compilation proceeds as with the batch compiler, and produces the same results as camlc -c. If the toplevel system is in debugging mode (option -g or function debug_mode below), the compilation is also performed in debugging mode, as when giving the -g option to the batch compiler. The result of the compilation is left in files (.zo, .zi, .zix). The compiled code is not loaded in memory. Use load_object to load a .zo file produced by

compile.
value load_object : string -> unit

Load in memory the compiled bytecode contained in the given file. The .zo extension is automatically added to the file name, if not present. The bytecode file has been produced either by the standalone compiler camlc or by the compile function. Global identifiers defined in the file being loaded are entered in their own module, not in the top module, just as with the load function.

value trace : string -> unit

After the execution of trace "foo", all calls to the global function named foo will be “traced”. That is, the argument and the result are displayed for each call, as well as the exceptions escaping out of foo, either raised by foo itself, or raised by one of the functions called from foo. If foo is a curried function, each argument is printed as it is passed to the function. Only functions implemented in ML can be traced; system primitives such as string_length or user-supplied C functions cannot.

Chapter 5. The toplevel system (camllight) 57

value untrace : string -> unit

Executing untrace "foo" stops all tracing over the global function named foo.

value verbose_mode: bool -> unit

verbose_mode true causes the compile function to print the inferred types and other information. verbose_mode false reverts to the default silent behavior.

value install_printer : string -> unit
install_printer "printername" registers the function named printername as a printer

for objects whose types match its argument type. That is, the toplevel loop will call printername when it has such an object to print. The printing function printername must use the format library module to produce its output, otherwise the output of printername will not be correctly located in the values printed by the toplevel loop.

value remove_printer : string -> unit
remove_printer "printername" removes the function named printername from the table

of toplevel printers.

value set_print_depth : int -> unit

set_print_depth n limits the printing of values to a maximal depth of n. The parts of values whose depth exceeds n are printed as ... (ellipsis).

value set_print_length : int -> unit

set_print_length n limits the number of value nodes printed to at most n. Remaining parts of values are printed as ... (ellipsis).

value debug_mode: bool -> unit

Set whether extended module interfaces must be used debug_mode true or not debug_mode false. Extended module interfaces are .zix files that describe the actual implementation of a module, including private types and variables. They are generated when compiling with camlc -g, or with the compile function above when debug_mode is true. When debug_mode is true, toplevel phrases can refer to private types and variables of modules, and private functions can be traced with trace. Setting debug_mode true is equivalent to starting the toplevel with the -g option.

value cd : string -> unit

Change the current working directory.

value directory : string -> unit

Add the given directory to the search path for files. Same behavior as the -I option or the

#directory directive.

5.3 The toplevel and the module system

Toplevel phrases can refer to identifiers defined in modules other than the top module with the same mechanisms as for separately compiled modules: either by using qualified identifiers (modulename__localname), or by using unqualified identifiers that are automatically completed by searching the list of opened modules. (See section 2.2.) The modules opened at startup are given in the documentation for the standard library. Other modules can be opened with the #open directive.

However, before referencing a global variable from a module other than the top module, a definition of that global variable must have been entered in memory. At start-up, the toplevel system contains the definitions for all the identifiers in the standard library. The definitions for user modules can be entered with the load or load_object functions described above. Referencing a global variable for which no definition has been provided by load or load_object results in

the error “Identifier foo__bar is referenced before being defined”. Since this is a tricky

point, let us consider some examples.

1. The library function sub_string is defined in module string. This module is part of the standard library, and is one of the modules automatically opened at start-up. Hence, both phrases

             sub_string "qwerty" 1 3;;
             string__sub_string "qwerty" 1 3;;

are correct, without having to use #open, load, or load_object.

2. The library function printf is defined in module printf. This module is part of the standard library, but it is not automatically opened at start-up. Hence, the phrase

             printf__printf "%s %s" "hello" "world";;

is correctly executed, while

             printf "%s %s" "hello" "world";;

causes the error “Variable printf is unbound”, since none of the currently opened modules define a global with local name printf. However,

             #open "printf";;
             printf "%s %s" "hello" "world";;

executes correctly.

3. Assume the file foo.ml resides in the current directory, and contains the single phrase

             let x = 1;;

When the toplevel starts, references to x will cause the error “Variable x is unbound”. References to foo__x will cause the error “Cannot find file foo.zi”, since the typechecker is attempting to load the compiled interface for module foo in order to find the type of x. To load in memory the module foo, just do:

        load "foo";;

Then, references to foo__x typecheck and evaluate correctly. Since load does not open the module it loads, references to x will still fail with the error “Variable x is unbound”. You will have to do

        #open "foo";;

explicitly, for x to complete automatically into foo__x.

4. Finally, assume the file foo.ml above has been previously compiled with the camlc -c command. The current directory therefore contains a compiled interface foo.zi, claiming that foo__x is a global variable with type int, and a compiled bytecode file foo.zo, defining foo__x to have the value 1. When the toplevel starts, references to foo__x will cause the

error “foo__x is referenced before being defined”. In contrast with case 3 above, the

typechecker has succeeded in finding the compiled interface for module foo. But execution cannot proceed, because no definition for foo__x has been entered in memory. To do so, execute:

        load_object "foo";;

This loads the file foo.zo in memory, therefore defining foo__x. Then, references to foo__x evaluate correctly. As in case 3 above, references to x still fail, because load_object does not open the module it loads. Again, you will have to do

        #open "foo";;

explicitly, for x to complete automatically into foo__x.

5.4 Common errors

This section describes and explains the most frequently encountered error messages.

Cannot find file filename The named file could not be found in the current directory, nor in the directories of the search path.

If filename has the format mod.zi, this means the current phrase references identifiers from module mod, but you have not yet compiled an interface for module mod. Fix: either load the file mod.ml, which will also create in memory the compiled interface for module mod; or use camlc to compile mod.mli or mod.ml, creating the compiled interface mod.zi, before you start the toplevel.

If filename has the format mod.zo, this means you are trying to load with load_object a bytecode object file that does not exist yet. Fix: compile mod.ml with camlc before you start the toplevel. Or, use load instead of load_object to load the source code instead of a compiled object file.

If filename has the format mod.ml, this means load or include could not find the specified source file. Fix: check the spelling of the file name, or write it if it does not exist.

mod__name is referenced before being defined You have neglected to load in memory an implementation for a module, with load or load_object. This is explained in full detail in section 5.3 above.

Corrupted compiled interface file filename See section 4.4.

Expression of type t_{1} cannot be used with type t_{2} See section 4.4.

The type inferred for the value name, that is, t, contains type variables that cannot be generalized

See section 4.4.

5.5 Building custom toplevel systems: camlmktop

The camlmktop command builds Caml Light toplevels that contain user code preloaded at start-up.

Mac: This command is not available in the Macintosh version.

The camlmktop command takes as argument a set of .zo files, and links them with the object files that implement the Caml Light toplevel. The typical use is:

        camlmktop -o mytoplevel foo.zo bar.zo gee.zo

This creates the bytecode file mytoplevel, containing the Caml Light toplevel system, plus the code from the three .zo files. To run this toplevel, give it as argument to the camllight command:

        camllight mytoplevel

This starts a regular toplevel loop, except that the code from foo.zo, bar.zo and gee.zo is already loaded in memory, just as if you had typed:

        load_object "foo";;
        load_object "bar";;
        load_object "gee";;

on entrance to the toplevel. The modules foo, bar and gee are not opened, though; you still have to do

        #open "foo";;

yourself, if this is what you wish.

5.6 Options

The following command-line options are recognized by camlmktop.

-ccopt option

Pass the given option to the C compiler and linker, when linking in “custom runtime” mode. See the corresponding option for camlc, in chapter 4.

-custom

Link in “custom runtime” mode. See the corresponding option for camlc, in chapter 4.

-g Add debugging information to the toplevel file produced, which can then be debugged with camldebug (chapter 9).

-I directory Add the given directory to the list of directories searched for compiled object code files (.zo).

-o exec-file Specify the name of the toplevel file produced by the linker.

Unix: The default is camltop.out.

PC: The default is camltop.exe. The name must have .exe extension.

62

The runtime system (camlrun)

The camlrun command executes bytecode files produced by the linking phase of the camlc command.

Mac: This command is a MPW tool, not a standalone Macintosh application.

6.1 Overview

The camlrun command comprises three main parts: the bytecode interpreter, that actually executes bytecode files; the memory allocator and garbage collector; and a set of C functions that implement primitive operations such as input/output.

The usage for camlrun is:

camlrun options bytecode-executable arg_{1} ... arg_{n}

The first non-option argument is taken to be the name of the file containing the executable bytecode. (That file is searched in the executable path as well as in the current directory.) The remaining arguments are passed to the Caml Light program, in the string array sys__command_line. Element 0 of this array is the name of the bytecode executable file; elements 1 to n are the remaining arguments arg_{1} to arg_{n}.

As mentioned in chapter 4, in most cases, the bytecode executable files produced by the camlc command are self-executable, and manage to launch the camlrun command on themselves automatically. That is, assuming caml.out is a bytecode executable file,

        caml.out arg_{1} ... arg_{n}

works exactly as

        camlrun caml.out arg_{1} ... arg_{n}

Notice that it is not possible to pass options to camlrun when invoking caml.out directly.

6.2 Options

The following command-line option is recognized by camlrun.

-V Print out the camlrun version number. Exit immediately without executing any byte-code file.

The following environment variable are also consulted:

CAMLRUNPARAM

Set the garbage collection parameters. This variable must be a sequence of parameter specifications. A parameter specification is an option letter followed by an = sign and a decimal number. There are four options, corresponding to the four fields of the control record documented in section 14.5:

s (minor_heap_size) Size of the minor heap. i (major_heap_increment) Minimum size increment for the major heap. o (space_overhead) The major GC speed setting. v (verbose) Whether to print GC messages or not. 0 is false; 1 is true; other values may

give unexpected results.

For example, under csh the command

        setenv CAMLRUNPARAM 's=250000 v=1'

tells a subsequent camlrun to set its initial minor heap size to about 1 megabyte (on a 32-bit machine) and to print its GC messages.

PATH

List of directories searched to find the bytecode executable file.

6.3 Common errors

This section describes and explains the most frequently encountered error messages.

filename: no such file or directory

If filename is the name of a self-executable bytecode file, this means that either that file does not exist, or that it failed to run the camlrun bytecode interpreter on itself. The second possibility indicates that Caml Light has not been properly installed on your system.

Cannot exec camlrun

(When launching a self-executable bytecode file.) The camlrun command could not be found in the executable path. Check that Caml Light has been properly installed on your system.

Cannot find the bytecode file

The file that camlrun is trying to execute (e.g. the file given as first non-option argument to camlrun) either does not exist, or is not a valid executable bytecode file.

Truncated bytecode file

The file that camlrun is trying to execute is not a valid executable bytecode file. Probably it has been truncated or mangled since created. Erase and rebuild it.

Uncaught exception

The program being executed contains a “stray” exception. That is, it raises an exception at some point, and this exception is never caught. This causes immediate termination of the program. If you wish to know which exception thus escapes, use the printexc__f function from the standard library (and don’t forget to link your program with the -g option).

Out of memory

The program being executed requires more memory than available. Either the program builds too large data structures; or the program contains too many nested function calls, and the stack overflows. In some cases, your program is perfectly correct, it just requires more memory than your machine provides. (This happens quite frequently on small microcomputers, but is unlikely on Unix machines.) In other cases, the “out of memory” message reveals an error in your program: non-terminating recursive function, allocation of an excessively large array or string, attempts to build an infinite list or other data structure, …

To help you diagnose this error, run your program with the -v option to camlrun. If it displays lots of “Growing stack… ” messages, this is probably a looping recursive function. If it displays lots of “Growing heap… ” messages, with the heap size growing slowly, this is probably an attempt to construct a data structure with too many (infinitely many?) cells. If it displays few “Growing heap… ” messages, but with a huge increment in the heap size, this is probably an attempt to build an excessively large array or string.

66

The librarian (camllibr)

Mac: This command is a MPW tool, not a standalone Macintosh application.

7.1 Overview

The camllibr program packs in one single file a set of bytecode object files (.zo files). The resulting file is also a bytecode object file and also has the .zo extension. It can be passed to the link phase of the camlc compiler in replacement of the original set of bytecode object files. That is, after running

        camllibr -o library.zo mod1.zo mod2.zo mod3.zi mod4.zo

all calls to the linker with the form

        camlc ... library.zo ...

are exactly equivalent to

        camlc ... mod1.zo mod2.zo mod3.zi mod4.zo ...

The typical use of camllibr is to build a library composed of several modules: this way, users of the library have only one .zo file to specify on the command line to camlc, instead of a bunch of .zo files, one per module contained in the library.

The linking phase of camlc is clever enough to discard the code corresponding to useless phrases: in particular, definitions for global variables that are never used after their definitions. Hence, there is no problem with putting many modules, even rarely used ones, into one single library: this will not result in bigger executables.

The usage for camllibr is:

        camllibr options file_{1}.zo ... file_{n}.zo

where file_{1}.zo through file_{n}.zo are the object files to pack together. The order in which these file names are presented on the command line is relevant: the compiled phrases contained in the library will be executed in that order. (Remember that it is a link-time error to refer to a global variable that has not yet been defined.)

7.2 Options

The following command-line options are recognized by camllibr.

-I directory Add the given directory to the list of directories searched for the input .zo files. By default, the current directory is searched first, then the standard library directory. Directories added with -I are searched after the current directory, but before the standard library directory. When several directories are added with several -I options on the command line, these directories are searched from right to left (the rightmost directory is searched first, the leftmost is searched last).

-o library-name Specify the name of the output file. The default is library.zo.

PC: The following option is also supported:

@response-file Process the files whose names are listed in file response-file, just as if these names appeared on the command line. File names in response-file are separated by blanks (spaces, tabs, newlines). This option allows to overcome silly limitations on the length of the command line.

7.3 Turning code into a library

To develop a library, it is usually more convenient to split it into several modules, that reflect the internal structure of the library. From the standpoint of the library users, however, it is preferable to view the library as a single module, with only one interface file (.zi file) and one implementation file (.zo file): linking is easier, and there is no need to put a bunch of #open directives, nor to have to remember the internal structure of the library.

The camllibr command allows having a single .zo file for the whole library. Here is how the Caml Light module system can be used (some say “abused”) to have a single .zi file for the whole library. To be more concrete, assume that the library comprises three modules, windows, images and buttons. The idea is to add a fourth module, mylib, that re-exports the public parts of windows, images and buttons. The interface mylib.mli contains definitions for those types that are public (exported with their definitions), declarations for those types that are abstract (exported without their definitions), and declarations for the functions that can be called from the user’s code:

(* File mylib.mli *)
type 'a option = None | Some of 'a;;    (* a public type *)
type window and image and button;;      (* three abstract types *)
value new_window : int -> int -> window (* the public functions *)
  and draw_image : image -> window -> int -> int -> unit
  and ...

The implementation of the mylib module simply equates the abstract types and the public functions to the corresponding types and functions in the modules windows, images and buttons:

(* File mylib.ml *)
type window == windows__win
 and image  == images__pixmap
 and button == buttons__t;;
let new_window = windows__open_window
and draw_image = images__draw
and ...

The files windows.ml, images.ml and buttons.ml can open the mylib module, to access the public types defined in the interface mylib.mli, such as the option type. Of course, these modules must not reference the abstract types nor the public functions, to avoid circularities.

Types such as windows__win in the example above can be exported by the windows module either abstractly or concretely (with their definition). Often, it is necessary to export them concretely, because the other modules in the library (images, buttons) need to build or destructure directly values of that type. Even if windows__win is exported concretely by the windows module, that type will remain abstract to the library user, since it is abstracted by the public interface

mylib.

The actual building of the library mylib proceeds as follows:

camlc -c mylib.mli              # create mylib.zi
camlc -c windows.mli windows.ml images.mli images.ml
camlc -c buttons.mli buttons.ml
camlc -c mylib.ml               # create mylib.zo
mv mylib.zo tmplib.zo           # renaming to avoid overwriting mylib.zo
camllibr -o mylib.zo windows.zo images.zo buttons.zo tmplib.zo

Then, copy mylib.zi and mylib.zo to a place accessible to the library users. The other .zi and .zo files need not be copied.

70

Lexer and parser generators (camllex, camlyacc)

This chapter describes two program generators: camllex, that produces a lexical analyzer from a set of regular expressions with associated semantic actions, and camlyacc, that produces a parser from a grammar with associated semantic actions.

These program generators are very close to the well-known lex and yacc commands that can be found in most C programming environments. This chapter assumes a working knowledge of lex and yacc: while it describes the input syntax for camllex and camlyacc and the main differences with lex and yacc, it does not explain the basics of writing a lexer or parser description in lex and yacc. Readers unfamiliar with lex and yacc are referred to “Compilers: principles, techniques, and tools” by Aho, Sethi and Ullman (Addison-Wesley, 1986), “Compiler design in C” by Holub (Prentice-Hall, 1990), or “Lex & Yacc”, by Mason and Brown (O’Reilly, 1990).

Streams and stream matching, as described in section 3.1, provide an alternative way to write lexers and parsers. The stream matching technique is more powerful than the combination of camllex and camlyacc in some cases (higher-order parsers), but less powerful in other cases (precedences). Choose whichever approach is more adapted to your parsing problem.

Mac: These commands are MPW tool, not standalone Macintosh applications.

8.1 Overview of camllex

The camllex command produces a lexical analyzer from a set of regular expressions with attached semantic actions, in the style of lex. Assuming the input file is lexer.mll, executing

        camllex lexer.mll

produces Caml Light code for a lexical analyzer in file lexer.ml. This file defines one lexing function per entry point in the lexer definition. These functions have the same names as the entry points. Lexing functions take as argument a lexer buffer, and return the semantic attribute of the corresponding entry point.

Lexer buffers are an abstract data type implemented in the standard library module lexing.

The functions create_lexer_channel, create_lexer_string and create_lexer from module

lexing create lexer buffers that read from an input channel, a character string, or any reading function, respectively. (See the description of module lexing in chapter 13.)

When used in conjunction with a parser generated by camlyacc, the semantic actions compute a value belonging to the type token defined by the generated parsing module. (See the description of camlyacc below.)

8.2 Syntax of lexer definitions

The format of lexer definitions is as follows:

{ header } rule entrypoint =

parse regexp { action } | … | regexp { action }

and entrypoint =

  parse . . .
and . . .
;;

Comments are delimited by (* and *), as in Caml Light.

8.2.1 Header

The header section is arbitrary Caml Light text enclosed in curly braces. It can be omitted. If it is present, the enclosed text is copied as is at the beginning of the output file. Typically, the header section contains the #open directives required by the actions, and possibly some auxiliary functions used in the actions.

8.2.2 Entry points

The names of the entry points must be valid Caml Light identifiers.

8.2.3 Regular expressions

The regular expressions are in the style of lex, with a more Caml-like syntax.

` char `

A character constant, with the same syntax as Caml Light character constants. Match the denoted character.

_Match any character.

eof Match the end of the lexer input.

" string "

A string constant, with the same syntax as Caml Light string constants. Match the corresponding sequence of characters.

[ character-set ]

Match any single character belonging to the given character set. Valid character sets are: single character constants ` c `; ranges of characters ` c_{1} ` - ` c_{2} ` (all characters between c_{1} and c_{2}, inclusive); and the union of two or more character sets, denoted by concatenation.

[^character-set ]

Match any single character not belonging to the given character set.

regexp *

(Repetition.) Match the concatenation of zero or more strings that match regexp.

regexp +

(Strict repetition.) Match the concatenation of one or more strings that match regexp.

regexp ?

(Option.) Match either the empty string, or a string matching regexp.

regexp_{1} | regexp_{2}

(Alternative.) Match any string that matches either regexp_{1} or regexp_{2}

regexp_{1} regexp_{2}

(Concatenation.) Match the concatenation of two strings, the first matching regexp_{1}, the second matching regexp_{2}.

( regexp )

Match the same strings as regexp.

Concerning the precedences of operators, * and + have highest precedence, followed by ?, then concatenation, then | (alternation).

8.2.4 Actions

The actions are arbitrary Caml Light expressions. They are evaluated in a context where the identifier lexbuf is bound to the current lexer buffer. Some typical uses for lexbuf, in conjunction with the operations on lexer buffers provided by the lexing standard library module, are listed below.

lexing__get_lexeme lexbuf

Return the matched string.

lexing__get_lexeme_char lexbuf n

Return the n^{th}character in the matched string. The first character corresponds to n = 0.

lexing__get_lexeme_start lexbuf

Return the absolute position in the input text of the beginning of the matched string. The first character read from the input text has position 0.

lexing__get_lexeme_end lexbuf

Return the absolute position in the input text of the end of the matched string. The first character read from the input text has position 0. entrypoint lexbuf

(Where entrypoint is the name of another entry point in the same lexer definition.) Recursively call the lexer on the given entry point. Useful for lexing nested comments, for example.

8.3 Overview of camlyacc

The camlyacc command produces a parser from a context-free grammar specification with attached semantic actions, in the style of yacc. Assuming the input file is grammar.mly, executing

camlyacc options grammar.mly

produces Caml Light code for a parser in the file grammar.ml, and its interface in file grammar.mli.

The generated module defines one parsing function per entry point in the grammar. These functions have the same names as the entry points. Parsing functions take as arguments a lexical analyzer (a function from lexer buffers to tokens) and a lexer buffer, and return the semantic attribute of the corresponding entry point. Lexical analyzer functions are usually generated from a lexer specification by the camllex program. Lexer buffers are an abstract data type implemented in the standard library module lexing. Tokens are values from the concrete type token, defined in the interface file grammar.mli produced by camlyacc.

8.4 Syntax of grammar definitions

Grammar definitions have the following format:

%{

header

%}

declarations

%%

rules

%%

trailer

Comments are enclosed between /* and */ (as in C) in the “declarations” and “rules” sections, and between (* and *) (as in Caml) in the “header” and “trailer” sections.

8.4.1 Header and trailer

The header and the trailer sections are Caml Light code that is copied as is into file grammar.ml. Both sections are optional. The header goes at the beginning of the output file; it usually contains #open directives required by the semantic actions of the rules. The trailer goes at the end of the output file.

8.4.2 Declarations

Declarations are given one per line. They all start with a % sign.

%token symbol … symbol

Declare the given symbols as tokens (terminal symbols). These symbols are added as constant constructors for the token concrete type.

%token< type > symbol … symbol

Declare the given symbols as tokens with an attached attribute of the given type. These symbols are added as constructors with arguments of the given type for the token concrete type. The type part is an arbitrary Caml Light type expression, except that all type constructor names must be fully qualified (e.g. modname__typename) for all types except standard built-in types, even if the proper #open directives (e.g. #open "modname") were given in the header section. That’s because the header is copied only to the .ml output file, but not to the .mli output file, while the type part of a %token declaration is copied to both.

%start symbol … symbol

Declare the given symbols as entry points for the grammar. For each entry point, a parsing function with the same name is defined in the output module. Non-terminals that are not declared as entry points have no such parsing function. Start symbols must be given a type with the %type directive below.

%type< type > symbol … symbol

Specify the type of the semantic attributes for the given symbols. This is mandatory for start symbols only. Other nonterminal symbols need not be given types by hand: these types will be inferred when running the output files through the Caml Light compiler (unless the -s option is in effect). The type part is an arbitrary Caml Light type expression, except that all type constructor names must be fully qualified (e.g. modname__typename) for all types except standard built-in types, even if the proper #open directives (e.g. #open "modname") were given in the header section. That’s because the header is copied only to the .ml output file, but not to the .mli output file, while the type part of a %token declaration is copied to both.

%left symbol … symbol

%right symbol … symbol

%nonassoc symbol … symbol

Associate precedences and associativities to the given symbols. All symbols on the same line are given the same precedence. They have higher precedence than symbols declared before in a %left, %right or %nonassoc line. They have lower precedence than symbols declared after in a %left, %right or %nonassoc line. The symbols are declared to associate to the left (%left), to the right (%right), or to be non-associative (%nonassoc). The symbols are usually tokens. They can also be dummy nonterminals, for use with the %prec directive inside the rules.

8.4.3 Rules

The syntax for rules is as usual:

nonterminal :

symbol … symbol { semantic-action } | … | symbol … symbol { semantic-action }

;

Rules can also contain the %prec symbol directive in the right-hand side part, to override the default precedence and associativity of the rule with the precedence and associativity of the given symbol.

Semantic actions are arbitrary Caml Light expressions, that are evaluated to produce the semantic attribute attached to the defined nonterminal. The semantic actions can access the semantic attributes of the symbols in the right-hand side of the rule with the $ notation: $1 is the attribute for the first (leftmost) symbol, $2 is the attribute for the second symbol, etc.

Actions occurring in the middle of rules are not supported. Error recovery is not implemented.

8.5 Options

The camlyacc command recognizes the following options:

-v Generate a description of the parsing tables and a report on conflicts resulting from ambiguities in the grammar. The description is put in file grammar.output.

-s Generate a grammar.ml file with smaller phrases. Semantic actions are presented in the grammar.ml output file as one large vector of functions. By default, this vector is built by a single phrase. When the grammar is large, or contains complicated semantic actions, the resulting phrase may require large amounts of memory to be compiled by Caml Light. With the -s option, the vector of actions is constructed incrementally, one phrase per action. This lowers the memory requirements for the compiler, but it is no longer possible to infer the types of nonterminal symbols: typechecking is turned off on symbols that do not have a type specified by a %type directive.

-bprefix Name the output files prefix.ml, prefix.mli, prefix.output, instead of the default naming convention.

8.6 A complete example

The all-time favorite: a desk calculator. This program reads arithmetic expressions on standard input, one per line, and prints their values. Here is the grammar definition:

        /* File parser.mly */
        %token <int> INT
        %token PLUS MINUS TIMES DIV

%token LPAREN RPAREN
%token EOL
%left PLUS MINUS        /* lowest precedence */
%left TIMES DIV         /* medium precedence */
%nonassoc UMINUS        /* highest precedence */
%start Main             /* the entry point */
%type <int> Main
%%
Main:
    Expr EOL                { $1 }
;
Expr:
    INT                     { $1 }
  | LPAREN Expr RPAREN      { $2 }
  | Expr PLUS Expr          { $1 + $3 }
  | Expr MINUS Expr         { $1 - $3 }
  | Expr TIMES Expr         { $1 * $3 }
  | Expr DIV Expr           { $1 / $3 }
  | MINUS Expr %prec UMINUS { - $2 }
;

Here is the definition for the corresponding lexer:

(* File lexer.mll *)
{
#open "parser";;        (* The type token is defined in parser.mli *)
exception Eof;;
}
rule Token = parse
    [` ` `\t`]     { Token lexbuf }     (* skip blanks *)
  | [`\n` ]        { EOL }
  | [`0`-`9`]+     { INT(int_of_string (get_lexeme lexbuf)) }
  | `+`            { PLUS }
  | `-`            { MINUS }
  | `*`            { TIMES }
  | `/`            { DIV }
  | `(`            { LPAREN }
  | `)`            { RPAREN }
  | eof            { raise Eof }
;;

Here is the main program, that combines the parser with the lexer:

(* File calc.ml *)
try
  let lexbuf = lexing__create_lexer_channel std_in in
  while true do

    let result = parser__Main lexer__Token lexbuf in
      print_int result; print_newline(); flush std_out
  done
with Eof ->
  ()
;;

To compile everything, execute:

camllex lexer.mll       # generates lexer.ml
camlyacc parser.mly     # generates parser.ml and parser.mli
camlc -c parser.mli
camlc -c lexer.ml
camlc -c parser.ml
camlc -c calc.ml
camlc -o calc lexer.zo parser.zo calc.zo

The debugger (camldebug)

This chapter describes the Caml Light source-level replay debugger camldebug.

Unix: The debugger resides in the directory contrib/debugger in the distribution. It requires a Unix system that provides BSD sockets.

Mac: The debugger is not available.

PC: The debugger is not available.

9.1 Compiling for debugging

Before the debugger can be used, the program must be compiled and linked with the -g option: all .zo files that are part of the program should have been created with camlc -g, and they must be linked together with camlc -g.

Compiling with -g entails no penalty on the running time of programs: .zo files and bytecode executable files are bigger and take slightly longer to produce, but the executable files run at exactly the same speed as if they had been compiled without -g. It is therefore perfectly acceptable to compile always in -g mode.

9.2 Invocation

9.2.1 Starting the debugger

The Caml Light debugger is invoked by running the program camldebug with the name of the bytecode executable file as argument:

        camldebug program

The following command-line options are recognized:

-stdlib directory Look for the standard library files in directory instead of in the default directory.

-s socket Use socket for communicating with the debugged program. See the description of the command set socket (section 9.8.7) for the format of socket.

-c count Set the maximum number of checkpoints to count.

-cd directory Run the debugger program from the working directory directory, instead of the current directory.

-emacs

Tell the debugger it is executing under Emacs. (See section 11.4 for information on how to run the debugger under Emacs.)

9.2.2 Quitting the debugger

The command quit exits the debugger. You can also exit the debugger by typing an end-of-file character (usually ctrl-D).

Typing an interrupt character (usually ctrl-C) will not exit the debugger, but will terminate the action of any debugger command that is in progress and return to the debugger command level.

9.3 Commands

A debugger command is a single line of input. It starts with a command name, which is followed by arguments depending on this name. Examples:

        run
        goto 1000
        set arguments arg1 arg2

A command name can be truncated as long as there is no ambiguity. For instance, go 1000 is understood as goto 1000, since there are no other commands whose name starts with go. For the most frequently used commands, ambiguous abbreviations are allowed. For instance, r stands for run even though there are others commands starting with r. You can test the validity of an abbreviation using the help command.

If the previous command has been successful, a blank line (typing just RET) will repeat it.

9.3.1 Getting help

The Caml Light debugger has a simple on-line help system, which gives a brief description of each command and variable.

help

Print the list of commands.

help command Give help about the command command. help set variable, help show variable

Give help about the variable variable. The list of all debugger variables can be obtained with

help set.
help info topic

Give help about topic. Use help info to get a list of known topics.

9.3.2 Accessing the debugger state

set variable value Set the debugger variable variable to the value value.

show variable Print the value of the debugger variable variable.

info subject Give information about the given subject. For instance, info breakpoints will print the list of all breakpoints.

9.4 Executing a program

9.4.1 Events

Events are “interesting” locations in the source code, corresponding to the beginning or end of evaluation of “interesting” sub-expressions. Events are the unit of single-stepping (stepping goes to the next or previous event encountered in the program execution). Also, breakpoints can only be set at events. Thus, events play the role of line numbers in debuggers for conventional languages.

During program execution, a counter is incremented at each event encountered. The value of this counter is referred as the current time. Thanks to reverse execution, it is possible to jump back and forth to any time of the execution.

Here is where the debugger events (written ▷◁) are located in the source code:

• Following a function application:

(f arg)▷◁

• After receiving an argument to a function:

fun x▷◁ y▷◁ z -> ▷◁ ...

If a curried function is defined by pattern-matching with several cases, events corresponding to the passing of arguments are displayed on the first case of the function, because patternmatching has not yet determined which case to select:

fun pat1▷◁ pat2▷◁ pat3 -> ▷◁ ...
  | ...

• On each case of a pattern-matching definition (function, match… with construct, try… with construct):

     function pat1 -> ▷◁ expr1
            | ...
            | patN -> ▷◁ exprN

• Between subexpressions of a sequence:

     expr1; ▷◁ expr2; ▷◁ ...; ▷◁ exprN

• In the two branches of a conditional expression:

     if cond then ▷◁ expr1 else ▷◁ expr2

• At the beginning of each iteration of a loop:

     while cond do ▷◁ body done
     for i = a to b do ▷◁ body done

Exceptions: A function application followed by a function return is replaced by the compiler by a jump (tail-call optimization). In this case, no event is put after the function application. Also, no event is put after a function application when the function is a primitive function (written in C). Finally, several events may correspond to the same location in the compiled program. Then, the debugger cannot distinguish them, and selects one of the events to associate with the given code location. The event chosen is a “function application” event if there is one at that location, or otherwise the event which appears last in the source. This heuristic generally picks the “most interesting” event associated with the code location.

9.4.2 Starting the debugged program

The debugger starts executing the debugged program only when needed. This allows setting breapoints or assigning debugger variables before execution starts. There are several ways to start execution:

run Run the program until a breakpoint is hit, or the program terminates.

step 0

Load the program and stop on the first event.

goto time

Load the program and execute it until the given time. Useful when you already know approximately at what time the problem appears. Also useful to set breakpoints on function values that have not been computed at time 0 (see section 9.5).

The execution of a program is affected by certain information it receives when the debugger starts it, such as the command-line arguments to the program and its working directory. The debugger provides commands to specify this information (set arguments and cd). These commands must be used before program execution starts. If you try to change the arguments or the working directory after starting your program, the debugger will kill the program (after asking for confirmation).

9.4.3 Running the program

The following commands execute the program forward or backward, starting at the current time. The execution will stop either when specified by the command or when a breakpoint is encountered.

run Execute the program forward from current time. Stops at next breakpoint or when the program terminates.

reverse

Execute the program backward from current time. Mostly useful to go to the last breakpoint encountered before the current time.

step [count] Run the program and stop at the next event. With an argument, do it count times.

backstep [count]

Run the program backward and stop at the previous event. With an argument, do it count times.

next [count] Run the program and stop at the next event, skipping over function calls. With an argument, do it count times.

finish

Run the program until the current function returns.

9.4.4 Time travel

You can jump directly to a given time, without stopping on breakpoints, using the goto command.

As you move through the program, the debugger maintains an history of the successive times you stop at. The last command can be used to revisit these times: each last command moves one step back through the history. That is useful mainly to undo commands such as step and next.

goto time

Jump to the given time.

last [count] Go back to the latest time recorded in the execution history. With an argument, do it count times.

set history size

Set the size of the execution history.

9.4.5 Killing the program

kill

Kill the program being executed. This command is mainly useful if you wish to recompile the program without leaving the debugger.

9.5 Breakpoints

A breakpoint causes the program to stop whenever a certain point in the program is reached. It can be set in several ways using the break command. Breakpoints are assigned numbers when set, for further reference.

break

Set a breakpoint at the current position in the program execution. The current position must be on an event (i.e., neither at the beginning, nor at the end of the program).

break function Set a breakpoint at the beginning of function. This works only when the functional value of the identifier function has been computed and assigned to the identifier. Hence this command cannot be used at the very beginning of the program execution, when all identifiers are still undefined. Moreover, C functions are not recognized by the debugger.

break @ [module] line Set a breakpoint in module module (or in the current module if module is not given), at the first event of line line.

break @ [module] line column Set a breakpoint in module module (or in the current module if module is not given), at the event closest to line line, column column.

break @ [module] # character Set a breakpoint in module module at the event closest to character number character.

break address Set a breakpoint at the code address address.

delete [breakpoint-numbers] Delete the specified breakpoints. Without argument, all breakpoints are deleted (after asking for confirmation).

info breakpoints

Print the list of all breakpoints.

9.6 The call stack

Each time the program performs a function application, it saves the location of the application (the return address) in a block of data called a stack frame. The frame also contains the local variables of the caller function. All the frames are allocated in a region of memory called the call stack. The command backtrace (or bt) displays parts of the call stack.

At any time, one of the stack frames is “selected” by the debugger; several debugger commands refer implicitly to the selected frame. In particular, whenever you ask the debugger for the value of a local variable, the value is found in the selected frame. The commands frame, up and down select whichever frame you are interested in.

When the program stops, the debugger automatically selects the currently executing frame and describes it briefly as the frame command does.

frame

Describe the currently selected stack frame.

frame frame-number Select a stack frame by number and describe it. The frame currently executing when the program stopped has number 0; its caller has number 1; and so on up the call stack.

backtrace [count], bt [count] Print the call stack. This is useful to see which sequence of function calls led to the currently executing frame. With a positive argument, print only the innermost count frames. With a negative argument, print only the outermost -count frames.

up [count] Select and display the stack frame just “above” the selected frame, that is, the frame that called the selected frame. An argument says how many frames to go up.

down [count] Select and display the stack frame just “below” the selected frame, that is, the frame that was called by the selected frame. An argument says how many frames to go down.

9.7 Examining variable values

The debugger can print the current value of a program variable (either a global variable or a local variable relative to the selected stack frame). It can also print selected parts of a value by matching it against a pattern.

Variable names can be specified either fully qualified (module-name__var-name) or unqualified (var-name). Unqualified names either correspond to local variables, or are completed into fully qualified global names by looking at a list of “opened” modules that define the same name (see section 9.8.5 for how to open modules in the debugger.) The completion follows the same rules as in the Caml Light language (see section 2.2).

print variables Print the values of the given variables.

match variable pattern Match the value of the given variable against a pattern, and print the values bound to the identifiers in the pattern.

The syntax of patterns for the match command extends the one for Caml Light patterns:

pattern ::= ident

                             | _

| ( pattern )
| ncconstr pattern
| pattern , pattern {, pattern}
| { label = pattern {; label = pattern} }

                             | []

| [ pattern {; pattern} ]
| pattern :: pattern
| # integer-literal pattern
| > pattern

The pattern ident, where ident is an identifier, matches any value, and binds the identifier to this value. The pattern # n pattern matches a list, a vector or a tuple whose n-th element matches pattern. The pattern > pattern matches any constructed value whose argument matches pattern, regardless of the constructor; it is a shortcut for skipping a constructor.

Example: assuming the value of a is Constr{x = [1;2;3;4]}, the command match a > {x =

# 2 k} prints k = 3.
set print_depth d

Limit the printing of values to a maximal depth of d.

set print_length l

Limit the printing of values to at most l nodes printed.

9.8 Controlling the debugger

9.8.1 Setting the program name and arguments

set program file

Set the program name to file.

set arguments arguments

Give arguments as command-line arguments for the program.

A shell is used to pass the arguments to the debugged program. You can therefore use wildcards, shell variables, and file redirections inside the arguments. To debug programs that read from standard input, it is recommended to redirect their input from a file (using set arguments < input-file), otherwise input to the program and input to the debugger are not properly separated.

9.8.2 How programs are loaded

The loadingmode variable controls how the program is executed.

set loadingmode direct

The program is run directly by the debugger. This is the default mode.

set loadingmode runtime

The debugger execute the Caml Light runtime camlrun on the program. Rarely useful; moreover it prevents the debugging of programs compiled in “custom runtime” mode.

set loadingmode manual

The user starts manually the program, when asked by the debugger. Allows remote debugging (see section 9.8.7).

9.8.3 Search path for files

The debugger searches for source files and compiled interface files in a list of directories, the search path. The search path initially contains the current directory . and the standard library directory. The directory command adds directories to the path.

Whenever the search path is modified, the debugger will clear any information it may have cached about the files.

directory directorynames Add the given directories to the search path. These directories are added at the front, and will therefore be searched first.

directory

Reset the search path. This requires confirmation.

9.8.4 Working directory

Each time a program is started in the debugger, it inherits its working directory from the current working directory of the debugger. This working directory is initially whatever it inherited from its parent process (typically the shell), but you can specify a new working directory in the debugger with the cd command or the -cd command-line option.

cd directory Set the working directory for camldebug to directory.

pwd Print the working directory for camldebug.

9.8.5 Module management

Like the Caml Light compiler, the debugger maintains a list of opened modules in order to resolves variable name ambiguities. The opened modules also affect the printing of values: whether fully qualified names or short names are used for constructors and record labels.

When a program is executed, the debugger automatically opens the modules of the standard library it uses.

open modules Open the given modules.

close modules Close the given modules.

info modules

List the modules used by the program, and the open modules.

9.8.6 Turning reverse execution on and off

In some cases, you may want to turn reverse execution off. This speeds up the program execution, and is also sometimes useful for interactive programs.

Normally, the debugger takes checkpoints of the program state from time to time. That is, it makes a copy of the current state of the program (using the Unix system call fork). If the variable checkpoints is set to off, the debugger will not take any checkpoints.

set checkpoints on/off

Select whether the debugger makes checkpoints or not.

9.8.7 Communication between the debugger and the program

The debugger communicate with the program being debugged through a Unix socket. You may need to change the socket name, for example if you need to run the debugger on a machine and your program on another.

set socket socket

Use socket for communication with the program. socket can be either a file name, or an Internet port specification host:port, where host is a host name or an Internet address in dot notation, and port is a port number on the host.

On the debugged program side, the socket name is passed either by the -D command line option to camlrun, or through the CAML_DEBUG_SOCKET environment variable.

9.8.8 Fine-tuning the debugger

Several variables enables to fine-tune the debugger. Reasonable defaults are provided, and you should normally not have to change them.

set processcount count

Set the maximum number of checkpoints to count. More checkpoints facilitate going far back in time, but use more memory and create more Unix processes.

As checkpointing is quite expensive, it must not be done too often. On the other hand, backward execution is faster when checkpoints are taken more often. In particular, backward single-stepping is more responsive when many checkpoints have been taken just before the current time. To finetune the checkpointing strategy, the debugger does not take checkpoints at the same frequency for long displacements (e.g. run) and small ones (e.g. step). The two variables bigstep and smallstep contain the number of events between two checkpoints in each case.

set bigstep count

Set the number of events between two checkpoints for long displacements.

set smallstep count

Set the number of events between two checkpoints for small displacements.

The following commands display information on checkpoints and events:

info checkpoints

Print a list of checkpoints.

info events [module]

Print the list of events in the given module (the current module, by default).

9.9 Miscellaneous commands

list [module] [beginning] [end] List the source of module module, from line number beginning to line number end. By default, 20 lines of the current module are displayed, starting 10 lines before the current position.

source filename Read debugger commands from the script filename.

90

Profiling (camlpro)

This chapter describes how the execution of Caml Light programs can be profiled, by recording how many times functions are called, branches of conditionals are taken, …

Mac: This command is not available.

PC: This command is not available.

10.1 Compiling for profiling

Before profiling an execution, the program must be compiled in profiling mode, using the -p option to the batch compiler camlc (see chapter 4). When compiling modules separately, the -p option must be given both when compiling the modules (production of .zo files) and when linking them together.

The amount of profiling information can be controlled by adding one or several letters after the -p option, indicating which parts of the program should be profiled:

a all options

f function calls : a count point is set at the beginning of function bodies

i if … then … else …: count points are set in both then branch and else branch

l while, for loops: a count point is set at the beginning of the loop body

m match branches: a count point is set at the beginning of the body of each branch

t try … with … branches: a count point is set at the beginning of the body of each branch

For instance, compiling with -pfilm profiles function calls, if… then … else…, loops and pattern matching.

The -p option without additional letters defaults to -pfm, meaning that only function calls and pattern matching are profiled.

10.2 Profiling an execution

Running a bytecode executable file that has been compiled and linked with -p records the execution counts for the specified parts of the program and saves them in a file called camlpro.dump in the current directory.

More precisely, the dump file camlpro.dump is written when the io__exit function is called. The linker, called with the -p option, adds io__exit 0 as the last phrase of the bytecode executable, in case the original program never calls io__exit. However, if the program terminates with an uncaught exception, the dump file will not be produced.

If a compatible dump file already exists in the current directory, then the profiling information is accumulated in this dump file. This allows, for instance, the profiling of several executions of a program on different inputs.

10.3 Printing profiling information

The camlpro command produces a source listing of the program modules where execution counts have been inserted as comments. For instance,

        camlpro foo.ml

prints the source code for the foo module, with comments indicating how many times the functions in this module have been called. Naturally, this information is accurate only if the source file has not been modified since the profiling execution took place.

The following options are recognized by camlpro:

compiler options -stdlib, -I, -include, -O, -open, -i, -lang

See chapter 4 for the detailed usage.

-f dumpfile Specifies an alternate dump file of profiling information

-F string Specifies an additional string to be output with profiling information. By default, camlpro will annotate progams with comments of the form (* n *) where n is the counter value for a profiling point. With option -F s, the annotation will be (* sn *).

An additional argument specifies the output file. For instance

camlpro -f ../test/camlpro.dump foo.ml foo_profiled.ml

will save the annotated program in file foo_profiled.ml. Otherwise, the annotated program is written on the standard output.

10.4 Known bugs

The following situation (file x.ml)

let a = 1;;
x__a ;;

will break the profiler. More precisely, one should avoid to refer to symbols of the current module with the qualified symbol syntax.

94

Using Caml Light under Emacs

This chapter describes how Caml Light can be used in conjunction with Gnu Emacs version 19 (version 18 is also partially supported).

Unix: The Emacs Lisp files implementing the Caml/Emacs interface are in contrib/camlmode in the distribution.

Mac: The Caml/Emacs interface is not available.

PC: The Caml/Emacs interface is not available.

11.1 Updating your .emacs

The following initializations must be added to your .emacs file:

(setq auto-mode-alist (cons '("\\.ml[iylp]?" . caml-mode) auto-mode-alist))
(autoload 'caml-mode "caml" "Major mode for editing Caml code." t)
(autoload 'run-caml "inf-caml" "Run an inferior Caml process." t)
(autoload 'camldebug "camldebug" "Run the Caml debugger." t)

11.2 The caml editing mode

The caml-mode function is a major editing mode for Caml source files. It provides the correct syntax tables, comment syntax, … for the Caml language. An extremely crude indentation facility is provided, as well as a slightly enhanced next-error command (to display the location of a compilation error). The following key bindings are performed:

TAB (function caml-indent-command)

At the beginning of a line, indent that line like the line above. Successive TABs increase the indentation level by 2 spaces (by default; can be set with the caml-mode-indentation variable).

M-TAB (function caml-unindent-command)

Decrease the indentation level of the current phrase.

C-x ` (function caml-next-error)

Display the next compilation error, just as next-error does. In addition, it puts the point and the mark around the exact location of the error (the subexpression that caused the error). Under Emacs 19, that subexpression is also highlighted.

M-C-h (function caml-mark-phrase)

Mark the Caml phrase that contains the point: the point is put at the beginning of the phrase and the mark at the end. Phrases are delimited by ;; (the final double-semicolon). This function does not properly ignore ;; inside string literals or comments.

C-x SPC

When the Caml debugger is running as an inferior process (section 11.4 below), set a breakpoint at the current position of the point.

M-C-x or C-c C-e (function caml-eval-phrase)

When a Caml toplevel is running as an inferior process (section 11.3 below), send it the the Caml phrase that contains the point. The phrase will then be evaluated by the inferior toplevel as usual. The phrase is delimited by ;; as described for the caml-mark-phrase command.

C-c C-r (function caml-eval-region)

Send the region to a Caml toplevel running in an inferior process.

11.3 Running the toplevel as an inferior process

M-x run-caml starts a Caml toplevel with input and output in an Emacs buffer named *inferior-caml*. This gives you the full power of Emacs to edit the input to the Caml toplevel. An history of input lines is maintained, as in Shell mode. This includes the following commands (see the function comint-mode for a complete description):

RET Send the current line to the toplevel.

M-n and M-p

Move to the next or previous line in the history.

M-r and M-s

Regexp search in the history.

C-c C-c

Send a break (interrupt signal) to the Caml toplevel.

Phrases can also be sent to the Caml toplevel for evaluation from any buffer in Caml mode,

using M-C-x, C-c C-e or C-c C-r.

11.4 Running the debugger as an inferior process

The Caml debugger is started by the command M-x camldebug, with argument the name of the executable file progname to debug. Communication with the debugger takes place in an Emacs buffer named *camldebug-progname*. The editing and history facilities of Shell mode are available for interacting with the debugger.

In addition, Emacs displays the source files containing the current event (the current position in the program execution) and highlights the location of the event. This display is updated synchronously with the debugger action.

The following bindings for the most common debugger commands are available in the *camldebug-progname* buffer (see section 9.3 for a full explanation of the commands):

M-r run command: execute the program forward.

M-s step command: execute the program one step forward.

M-b back command: execute the program one step backward.

M-l last command: go back one step in the command history.

C-c >

down command: select the stack frame below the current frame.

C-c <

up command: select the stack frame above the current frame.

C-c C-f

finish command: run till the current function returns.

In a buffer in Caml editing mode, C-x SPC sets a breakpoint at the current position of the point.

98

Interfacing C with Caml Light

This chapter describes how user-defined primitives, written in C, can be added to the Caml Light runtime system and called from Caml Light code.

12.1 Overview and compilation information

12.1.1 Declaring primitives

User primitives are declared in a module interface (a .mli file), in the same way as a regular ML value, except that the declaration is followed by the = sign, the function arity (number of arguments), and the name of the corresponding C function. For instance, here is how the input primitive is declared in the interface for the standard library module io:

        value input : in_channel -> string -> int -> int -> int
                    = 4 "input"

Primitives with several arguments are always curried. The C function does not necessarily have the same name as the ML function.

Values thus declared primitive in a module interface must not be implemented in the module implementation (the .ml file). They can be used inside the module implementation.

12.1.2 Implementing primitives

User primitives with arity n ≤ 5 are implemented by C functions that take n arguments of type value, and return a result of type value. The type value is the type of the representations for Caml Light values. It encodes objects of several base types (integers, floating-point numbers, strings, …), as well as Caml Light data structures. The type value and the associated conversion functions and macros are described in details below. For instance, here is the declaration for the C function implementing the input primitive:

        value input(channel, buffer, offset, length)
                value channel, buffer, offset, length;
        {
         ...
        }

When the primitive function is applied in a Caml Light program, the C function is called with the values of the expressions to which the primitive is applied as arguments. The value returned by the function is passed back to the Caml Light program as the result of the function application.

User primitives with arity greater than 5 are implemented by C functions that receive two arguments: a pointer to an array of Caml Light values (the values for the arguments), and an integer which is the number of arguments provided:

        value prim_with_lots_of_args(argv, argn)
                value * argv;
                int argn;
        {
          ... argv[0] ...;              /* The first argument */
          ... argv[6] ...;              /* The seventh argument */
        }

Implementing a user primitive is actually two separate tasks: on the one hand, decoding the arguments to extract C values from the given Caml Light values, and encoding the return value as a Caml Light value; on the other hand, actually computing the result from the arguments. Except for very simple primitives, it is often preferable to have two distinct C functions to implement these two tasks. The first function actually implements the primitive, taking native C values as arguments and returning a native C value. The second function, often called the “stub code”, is a simple wrapper around the first function that converts its arguments from Caml Light values to C values, call the first function, and convert the returned C value to Caml Light value. For instance, here is the stub code for the input primitive:

        value input(channel, buffer, offset, length)
                value channel, buffer, offset, length;
        {
          return Val_long(getblock((struct channel *) channel,
                                   &Byte(buffer, Long_val(offset)),
                                   Long_val(length)));
        }

(Here, Val_long, Long_val and so on are conversion macros for the type value, that will be described later.) The hard work is performed by the function getblock, which is declared as:

        long getblock(channel, p, n)
             struct channel * channel;
             char * p;
             long n;
        {
          ...
        }

To write C code that operates on Caml Light values, the following include files are provided:

Include file Provides

mlvalues.h definition of the value type, and conversion macros

alloc.h allocation functions (to create structured Caml Light objects)

memory.h miscellaneous memory-related functions (for in-place modification of structures, etc).

These files reside in the Caml Light standard library directory (usually /usr/local/lib/caml-light).

12.1.3 Linking C code with Caml Light code

The Caml Light runtime system comprises three main parts: the bytecode interpreter, the memory manager, and a set of C functions that implement the primitive operations. Some bytecode instructions are provided to call these C functions, designated by their offset in a table of functions (the table of primitives).

In the default mode, the Caml Light linker produces bytecode for the standard runtime system, with a standard set of primitives. References to primitives that are not in this standard set result in the “unavailable C primitive” error.

In the “custom runtime” mode, the Caml Light linker scans the bytecode object files (.zo files) and determines the set of required primitives. Then, it builds a suitable runtime system, by calling the native code linker with:

• the table of the required primitives

• a library that provides the bytecode interpreter, the memory manager, and the standard primitives

• libraries and object code files (.o files) mentioned on the command line for the Caml Light linker, that provide implementations for the user’s primitives.

This builds a runtime system with the required primitives. The Caml Light linker generates bytecode for this custom runtime system. The bytecode is appended to the end of the custom runtime system, so that it will be automatically executed when the output file (custom runtime + bytecode) is launched.

To link in “custom runtime” mode, execute the camlc command with:

• the -custom option

• the names of the desired Caml Light object files (.zo files)

• the names of the C object files and libraries (.o and .a files) that implement the required primitives. (Libraries can also be specified with the usual -l syntax.)

12.2 The value type

All Caml Light objects are represented by the C type value, defined in the include file mlvalues.h, along with macros to manipulate values of that type. An object of type value is either:

• an unboxed integer • a pointer to a block inside the heap (such as the blocks allocated through one of the alloc_*

functions below)

• a pointer to an object outside the heap (e.g., a pointer to a block allocated by malloc, or to a C variable).

12.2.1 Integer values

Integer values encode 31-bit signed integers. They are unboxed (unallocated).

12.2.2 Blocks

Blocks in the heap are garbage-collected, and therefore have strict structure constraints. Each block includes a header containing the size of the block (in words), and the tag of the block. The tag governs how the contents of the blocks are structured. A tag lower than No_scan_tag indicates a structured block, containing well-formed values, which is recursively traversed by the garbage collector. A tag greater than or equal to No_scan_tag indicates a raw block, whose contents are not scanned by the garbage collector. For the benefits of ad-hoc polymorphic primitives such as equality and structured input-output, structured and raw blocks are further classified according to their tags as follows:

Tag Contents of the block

0 to No_scan_tag − 1 A structured block (an array of Caml Light objects). Each

field is a value.

Closure_tag A closure representing a functional value. The first word is a pointer to a piece of bytecode, the second word is a value containing the environment.

String_tag A character string.

Double_tag A double-precision floating-point number. Abstract_tag A block representing an abstract datatype. Final_tag A block representing an abstract datatype with a “finaliza-

tion” function, to be called when the block is deallocated.

12.2.3 Pointers to outside the heap

Any pointer to outside the heap can be safely cast to and from the type value. This includes pointers returned by malloc, and pointers to C variables obtained with the & operator.

12.3 Representation of Caml Light data types

This section describes how Caml Light data types are encoded in the value type.

12.3.1 Atomic types

Caml type Encoding
int  Unboxed integer values.
char  Unboxed integer values (ASCII code).

                    float  Blocks with tag Double_tag.
                    string  Blocks with tag String_tag.

12.3.2 Product types

Tuples and arrays are represented by pointers to blocks, with tag 0.

Records are also represented by zero-tagged blocks. The ordering of labels in the record type declaration determines the layout of the record fields: the value associated to the label declared first is stored in field 0 of the block, the value associated to the label declared next goes in field 1, and so on.

12.3.3 Concrete types

Constructed terms are represented by blocks whose tag encode the constructor. The constructors for a given concrete type are numbered from 0 to the number of constructors minus one, following the order in which they appear in the concrete type declaration. Constant constructors are represented by zero-sized blocks (atoms), tagged with the constructor number. Non-constant constructors declared with a n-tuple as argument are represented by a block of size n, tagged with the constructor number; the n fields contain the components of its tuple argument. Other non-constant constructors are represented by a block of size 1, tagged with the constructor number; the field 0 contains the value of the constructor argument. Example:

Constructed term Representation
()  Size = 0, tag = 0
false  Size = 0, tag = 0
true  Size = 0, tag = 1
[]  Size = 0, tag = 0
h::t  Size = 2, tag = 1, first field = h, second field = t

12.4 Operations on values

12.4.1 Kind tests

• Is_int(v) is true if value v is an immediate integer, false otherwise

• Is_block(v) is true if value v is a pointer to a block, and false if it is an immediate integer.

12.4.2 Operations on integers

• Val_long(l) returns the value encoding the long int l

• Long_val(v) returns the long int encoded in value v • Val_int(i) returns the value encoding the int i

• Int_val(v) returns the int encoded in value v

12.4.3 Accessing blocks

• Wosize_val(v) returns the size of value v, in words, excluding the header.

• Tag_val(v) returns the tag of value v.

• Field(v, n) returns the value contained in the n^{th} field of the structured block v. Fields are numbered from 0 to Wosize_val(v) − 1.

• Code_val(v) returns the code part of the closure v.

• Env_val(v) returns the environment part of the closure v.

• string_length(v) returns the length (number of characters) of the string v.

• Byte(v, n) returns the n^{th} character of the string v, with type char. Characters are numbered

  from 0 to string_length(v) − 1.

• Byte_u(v, n) returns the n^{th} character of the string v, with type unsigned char. Characters are numbered from 0 to string_length(v) − 1.

• String_val(v) returns a pointer to the first byte of the string v, with type char *. This pointer is a valid C string: there is a null character after the last character in the string. However, Caml Light strings can contain embedded null characters, that will confuse the usual C functions over strings.

• Double_val(v) returns the floating-point number contained in value v, with type double.

The expressions Field(v, n), Code_val(v), Env_val(v), Byte(v, n), Byte_u(v, n) and Double_val(v) are valid l-values. Hence, they can be assigned to, resulting in an in-place modification of value v. Assigning directly to Field(v, n) must be done with care to avoid confusing the garbage collector (see below).

12.4.4 Allocating blocks

From the standpoint of the allocation functions, blocks are divided according to their size as zerosized blocks, small blocks (with size less than or equal to Max_young_wosize), and large blocks (with size greater than to Max_young_wosize). The constant Max_young_wosize is declared in the include file mlvalues.h. It is guaranteed to be at least 64 (words), so that any block with constant size less than or equal to 64 can be assumed to be small. For blocks whose size is computed at run-time, the size must be compared against Max_young_wosize to determine the correct allocation procedure.

• Atom(t) returns an “atom” (zero-sized block) with tag t. Zero-sized blocks are preallocated outside of the heap. It is incorrect to try and allocate a zero-sized block using the functions below. For instance, Atom(0) represents (), false and []; Atom(1) represents true. (As a

  convenience, mlvalues.h defines the macros Val_unit, Val_false and Val_true.)

• alloc(n, t) returns a fresh small block of size n ≤ Max_young_wosize words, with tag t. If this block is a structured block (i.e. if t < No_scan_tag), then the fields of the block (initially containing garbage) must be initialized with legal values (using direct assignment to the fields of the block) before the next allocation.

• alloc_tuple(n) returns a fresh small block of size n ≤ Max_young_wosize words, with tag 0. The fields of this block must be filled with legal values before the next allocation or modification.

• alloc_shr(n, t) returns a fresh block of size n, with tag t. The size of the block can be greater than Max_young_wosize. (It can also be smaller, but in this case it is more efficient to call alloc instead of alloc_shr.) If this block is a structured block (i.e. if t < No_scan_tag), then the fields of the block (initially containing garbage) must be initialized with legal values (using the initialize function described below) before the next allocation.

• alloc_string(n) returns a string value of length n characters. The string initially contains garbage.

• copy_string(s) returns a string value containing a copy of the null-terminated C string s (a

char *).

• copy_double(d) returns a floating-point value initialized with the double d.

• alloc_array(f, a) allocates an array of values, calling function f over each element of the input array a to transform it into a value. The array a is an array of pointers terminated by the null pointer. The function f receives each pointer as argument, and returns a value. The zero-tagged block returned by alloc_array(f, a) is filled with the values returned by the successive calls to f.

• copy_string_array(p) allocates an array of strings, copied from the pointer to a string array

p (a char **).

12.4.5 Raising exceptions

C functions cannot raise arbitrary exceptions. However, two functions are provided to raise two standard exceptions:

• failwith(s), where s is a null-terminated C string (with type char *), raises exception Failure with argument s.

• invalid_argument(s), where s is a null-terminated C string (with type char *), raises exception Invalid_argument with argument s.

12.5 Living in harmony with the garbage collector

Unused blocks in the heap are automatically reclaimed by the garbage collector. This requires some cooperation from C code that manipulates heap-allocated blocks.

Rule 1 After a structured block (a block with tag less than No_scan_tag) is allocated, all fields of this block must be filled with well-formed values before the next allocation operation. If the block has been allocated with alloc or alloc_tuple, filling is performed by direct assignment to the fields of the block:

        Field(v, n) = v_{n};

If the block has been allocated with alloc_shr, filling is performed through the initialize function:

        initialize(&Field(v, n), v_{n});

The next allocation can trigger a garbage collection. The garbage collector assumes that all structured blocks contain well-formed values. Newly created blocks contain random data, which generally do not represent well-formed values.

If you really need to allocate before the fields can receive their final value, first initialize with a constant value (e.g. Val_long(0)), then allocate, then modify the fields with the correct value (see rule 3).

Rule 2 Local variables containing values must be registered with the garbage collector (using the Push_roots and Pop_roots macros), if they are to survive a call to an allocation function.

Registration is performed with the Push_roots and Pop_roots macros. Push_roots(r,n) declares an array r of n values and registers them with the garbage collector. The values contained in r[0] to r[n-1] are treated like roots by the garbage collector. A root value has the following properties: if it points to a heap-allocated block, this block (and its contents) will not be reclaimed; moreover, if this block is relocated by the garbage collector, the root value is updated to point to the new location for the block. Push_roots(r,n) must occur in a C block exactly between the last local variable declaration and the first statement in the block. To un-register the roots, Pop_roots() must be called before the C block containing Push_roots(r,n) is exited. (Roots are automatically un-registered if a Caml exception is raised.)

Rule 3 Direct assignment to a field of a block, as in

        Field(v, n) = w;

is safe only if v is a block newly allocated by alloc or alloc_tuple; that is, if no allocation took place between the allocation of v and the assignment to the field. In all other cases, never assign directly. If the block has just been allocated by alloc_shr, use initialize to assign a value to a field for the first time:

        initialize(&Field(v, n), w);

Otherwise, you are updating a field that previously contained a well-formed value; then, call the modify function:

        modify(&Field(v, n), w);

To illustrate the rules above, here is a C function that builds and returns a list containing the two integers given as parameters:

value alloc_list_int(i1, i2)
      int i1, i2;
{
value result;
Push_roots(r, 1);
r[0] = alloc(2, 1);                     /* Allocate a cons cell */
Field(r[0], 0) = Val_int(i2);           /* car = the integer i2 */
Field(r[0], 1) = Atom(0);               /* cdr = the empty list [] */
result = alloc(2, 1);                   /* Allocate the other cons cell */
Field(result, 0) = Val_int(i1);         /* car = the integer i1 */
Field(result, 1) = r[0];                /* cdr = the first cons cell */
Pop_roots();
return result;
}

The “cons” cell allocated first needs to survive the allocation of the other cons cell; hence, the value returned by the first call to alloc must be stored in a registered root. The value returned by the second call to alloc can reside in the un-registered local variable result, since we won’t do any further allocation in this function.

In the example above, the list is built bottom-up. Here is an alternate way, that proceeds top-down. It is less efficient, but illustrates the use of modify.

value alloc_list_int(i1, i2)
      int i1, i2;
{
value tail;
Push_roots(r, 1);
r[0] = alloc(2, 1);                     /* Allocate a cons cell */
Field(r[0], 0) = Val_int(i1);           /* car = the integer i1 */
Field(r[0], 1) = Val_int(0);            /* A dummy value
tail = alloc(2, 1);                     /* Allocate the other cons cell */
Field(tail, 0) = Val_int(i2);           /* car = the integer i2 */
Field(tail, 1) = Atom(0);               /* cdr = the empty list [] */
modify(&Field(r[0], 1), tail);          /* cdr of the result = tail */
Pop_roots();
return r[0];
}

It would be incorrect to perform Field(r[0], 1) = tail directly, because the allocation of tail has taken place since r[0] was allocated.

12.6 A complete example

This section outlines how the functions from the Unix curses library can be made available to Caml Light programs. First of all, here is the interface curses.mli that declares the curses primitives and data types:

type window;;                   (* The type "window" remains abstract *)
value initscr: unit -> window = 1 "curses_initscr"
  and endwin: unit -> unit = 1 "curses_endwin"
  and refresh: unit -> unit = 1 "curses_refresh"
  and wrefresh : window -> unit = 1 "curses_wrefresh"
  and newwin: int -> int -> int -> int -> window = 4 "curses_newwin"
  and mvwin: window -> int -> int -> unit = 3 "curses_mvwin"
  and addch: char -> unit = 1 "curses_addch"
  and mvwaddch: window -> int -> int -> char -> unit = 4 "curses_mvwaddch"
  and addstr: string -> unit = 1 "curses_addstr"
  and mvwaddstr: window -> int -> int -> string -> unit = 4 "curses_mvwaddstr"
;; (* lots more omitted *)

To compile this interface:

        camlc -c curses.mli

To implement these functions, we just have to provide the stub code; the core functions are already implemented in the curses library. The stub code file, curses.o, looks like:

#include <curses.h>
#include <mlvalues.h>
value curses_initscr(unit)
        value unit;
{
  return (value) initscr();     /* OK to coerce directly from WINDOW * to value
                                   since that's a block created by malloc() */
}
value curses_wrefresh(win)
        value win;
{
  wrefresh((WINDOW *) win);
  return Val_unit;
}
value curses_newwin(nlines, ncols, x0, y0)
        value nlines, ncols, x0, y0;
{
  return (value) newwin(Int_val(nlines), Int_val(ncols),
                        Int_val(x0), Int_val(y0));
}
value curses_addch(c)
        value c;
{

  addch(Int_val(c));            /* Characters are encoded like integers */
  return Val_unit;
}
value curses_addstr(s)
        value s;
{
  addstr(String_val(s));
  return Val_unit;
}
/* This goes on for pages. */

(Actually, it would be better to create a library for the stub code, with each stub code function in a separate file, so that linking would pick only those functions from the curses library that are actually used.)

The file curses.c can be compiled with:

        cc -c -I/usr/local/lib/caml-light curses.c

or, even simpler,

        camlc -c curses.c

(When passed a .c file, the camlc command simply calls cc on that file, with the right -I option.) Now, here is a sample Caml Light program test.ml that uses the curses module:

#open "curses";;
let main_window = initscr () in
let small_window = newwin 10 5 20 10 in
  mvwaddstr main_window 10 2 "Hello";
  mvwaddstr small_window 4 3 "world";
  refresh();
  for i = 1 to 100000 do () done;
  endwin()
;;

To compile this program, run:

        camlc -c test.ml

Finally, to link everything together:

        camlc -custom -o test test.zo curses.o -lcurses

110

Part IV

The Caml Light library The core library

This chapter describes the functions provided by the Caml Light core library. This library is special in two ways:

• It is automatically linked with the user’s object code files by the camlc command (chapter 4). Hence, the globals defined by these libraries can be used in standalone programs without having to add any .zo file on the command line for the linking phase. Similarly, in interactive use, these globals can be used in toplevel phrases without having to load any .zo file in memory.

• The interfaces for the modules below are automatically “opened” when a compilation starts, or when the toplevel system is launched. Hence, it is possible to use unqualified identifiers to refer to the functions provided by these modules, without adding #open directives. Actually, the list of automatically opened modules depend on the -O option given to the compiler or to the toplevel system:

-O option Opened modules (reverse opening order)

        -O cautious (default) io, eq, int, float, ref, pair, list, vect, char,
                             string, bool, exc, stream, builtin
        -O fast              io, eq, int, float, ref, pair, list, fvect,
                             fchar, fstring, bool, exc, stream, builtin

        -O none              builtin

Conventions

For easy reference, the modules are listed below in alphabetical order of module names. For each module, the declarations from its interface file are printed one by one in typewriter font, followed by a short comment. All modules and the identifiers they export are indexed at the end of this report.

13.1 bool: boolean operations

value prefix &  : bool -> bool -> bool

value prefix && : bool -> bool -> bool
value prefix or : bool -> bool -> bool
value prefix || : bool -> bool -> bool

The boolean and is written e1 & e2 or e1 && e2. The boolean or is written e1 or e2 or e1 || e2. Both constructs are sequential, left-to-right: e2 is evaluated only if needed. Actually, e1 & e2 is equivalent to if e1 then e2 else false, and e1 or e2 is equivalent

to if e1 then true else e2.
value prefix not : bool -> bool

The boolean negation.

value string_of_bool : bool -> string

Return a string representing the given boolean.

13.2 builtin: base types and constructors

This module defines some types and exceptions for which the language provides special syntax, and are therefore treated specially by the compiler.

type int
type float
type string
type char

The types of integers, floating-point numbers, character strings, and characters, respectively.

type exn

The type of exception values.

type bool = false | true

The type of boolean values.

type 'a vect

The type of arrays whose elements have type 'a.

type unit = ()

The type of the unit value.

type 'a list = [] | prefix :: of 'a * 'a list

The type of lists.

type 'a option = None | Some of 'a

The type of optional values.

exception Match_failure of string * int * int

The exception raised when a pattern-matching fails. The argument indicates the position in the source code of the pattern-matching (source file name, position of the first character of the matching, position of the last character.

Chapter 13. The core library 115

13.3 char: character operations

value int_of_char : char -> int

Return the ASCII code of the argument.

value char_of_int : int -> char

Return the character with the given ASCII code. Raise Invalid_argument "char_of_int" if the argument is outside the range 0–255.

value string_of_char : char -> string

Return a string representing the given character.

value char_for_read : char -> string

Return a string representing the given character, with special characters escaped following the lexical conventions of Caml Light.

13.4 eq: generic comparisons

value prefix = : 'a -> 'a -> bool

e1 = e2 tests for structural equality of e1 and e2. Mutable structures (e.g. references and arrays) are equal if and only if their current contents are structurally equal, even if the two mutable objects are not the same physical object. Equality between functional values raises Invalid_argument. Equality between cyclic data structures may not terminate.

value prefix <> : 'a -> 'a -> bool

Negation of prefix =.

value prefix < : 'a -> 'a -> bool
value prefix <= : 'a -> 'a -> bool
value prefix > : 'a -> 'a -> bool
value prefix >= : 'a -> 'a -> bool

Structural ordering functions. These functions coincide with the usual orderings over integer, string and floating-point numbers, and extend them to a total ordering over all types. The ordering is compatible with prefix =. As in the case of prefix =, mutable structures are compared by contents. Comparison between functional values raises Invalid_argument. Comparison between cyclic structures may not terminate.

value compare: 'a -> 'a -> int

compare x y returns 0 if x=y, a negative integer if x<y, and a positive integer if x>y. The same restrictions as for = apply. compare can be used as the comparison function required by the set and map modules.

value min: 'a -> 'a -> 'a

Return the smaller of the two arguments.

value max: 'a -> 'a -> 'a

Return the greater of the two arguments.

value prefix == : 'a -> 'a -> bool

e1 == e2 tests for physical equality of e1 and e2. On integers and characters, it is the same as structural equality. On mutable structures, e1 == e2 is true if and only if physical modification of e1 also affects e2. On non-mutable structures, the behavior of prefix == is implementation-dependent, except that e1 == e2 implies e1 = e2.

value prefix != : 'a -> 'a -> bool

Negation of prefix ==.

13.5 exc: exceptions

value raise : exn -> 'a

Raise the given exception value.

A few general-purpose predefined exceptions.

exception Out_of_memory

Raised by the garbage collector, when there is insufficient memory to complete the computation.

exception Invalid_argument of string

Raised by library functions to signal that the given arguments do not make sense.

exception Failure of string

Raised by library functions to signal that they are undefined on the given arguments.

exception Not_found

Raised by search functions when the desired object could not be found.

exception Exit

This exception is not raised by any library function. It is provided for use in your programs.

value failwith : string -> 'a

Raise exception Failure with the given string.

value invalid_arg : string -> 'a

Raise exception Invalid_argument with the given string.

13.6 fchar: character operations, without sanity checks

This module implements the same functions as the char module, but does not perform bound checks on the arguments of the functions. The functions are therefore faster than those in the char module, but calling these functions with incorrect parameters (that is, parameters that would cause the Invalid_argument exception to be raised by the corresponding functions in the char module) can crash the program.

13.7 float: operations on floating-point numbers

value int_of_float : float -> int

Truncate the given float to an integer value. The result is unspecified if it falls outside the range of representable integers.

value float_of_int : int -> float

Convert an integer to floating-point.

value minus : float -> float
value minus_float : float -> float

Unary negation.

value prefix + : float -> float -> float
value prefix +. : float -> float -> float
value add_float : float -> float -> float

Addition.

value prefix - : float -> float -> float
value prefix -. : float -> float -> float
value sub_float : float -> float -> float

Subtraction.

value prefix * : float -> float -> float
value prefix *. : float -> float -> float
value mult_float : float -> float -> float

Product.

value prefix / : float -> float -> float
value prefix /. : float -> float -> float
value div_float : float -> float -> float

Division.

value prefix ** : float -> float -> float
value prefix **. : float -> float -> float
value power : float -> float -> float

Exponentiation.

value eq_float : float -> float -> bool
value prefix =. : float -> float -> bool

Floating-point equality. Equivalent to generic equality, just faster.

value neq_float : float -> float -> bool
value prefix <>. : float -> float -> bool

Negation of eq_float.

value prefix <. : float -> float -> bool
value lt_float : float -> float -> bool
value prefix >. : float -> float -> bool
value gt_float : float -> float -> bool
value prefix <=. : float -> float -> bool
value le_float : float -> float -> bool
value prefix >=. : float -> float -> bool
value ge_float : float -> float -> bool

Usual comparisons between floating-point numbers.

value acos : float -> float
value asin : float -> float
value atan : float -> float
value atan2 : float -> float -> float
value cos : float -> float
value cosh : float -> float
value exp : float -> float
value log : float -> float
value log10 : float -> float
value sin : float -> float
value sinh : float -> float
value sqrt : float -> float
value tan : float -> float
value tanh : float -> float

Usual transcendental functions on floating-point numbers.

value ceil : float -> float
value floor : float -> float

Round the given float to an integer value. floor f returns the greatest integer value less than or equal to f. ceil f returns the least integer value greater than or equal to f.

Chapter 13. The core library 119

value abs_float : float -> float

Return the absolute value of the argument.

value mod_float : float -> float -> float

fmod a b returns the remainder of a with respect to b.

value frexp : float -> float * int

frexp f returns the pair of the significant and the exponent of f (when f is zero, the significant x and the exponent n of f are equal to zero; when f is non-zero, they are defined

by f = x *. 2 ** n).
value ldexp : float -> int -> float
ldexp x n returns x *. 2 ** n.
value modf : float -> float * float

modf f returns the pair of the fractional and integral part of f.

value string_of_float : float -> string

Convert the given float to its decimal representation.

value float_of_string : string -> float

Convert the given string to a float, in decimal. The result is unspecified if the given string is not a valid representation of a float.

13.8 fstring: string operations, without sanity checks

This module implements the same functions as the string module, but does not perform bound checks on the arguments of the functions. The functions are therefore faster than those in the string module, but calling these functions with incorrect parameters (that is, parameters that would cause the Invalid_argument exception to be raised by the corresponding functions in the string module) can crash the program.

13.9 fvect: operations on vectors, without sanity checks

This module implements the same functions as the vect module, but does not perform bound checks on the arguments of the functions. The functions are therefore faster than those in the vect module, but calling these functions with incorrect parameters (that is, parameters that would cause the Invalid_argument exception to be raised by the corresponding functions in the vect module) can crash the program.

13.10 int: operations on integers

Integers are 31 bits wide (or 63 bits on 64-bit processors). All operations are taken modulo 2^{31} (or 2^{63}). They do not fail on overflow.

exception Division_by_zero
value minus : int -> int
value minus_int : int -> int

Unary negation. You can write -e instead of minus e.

value succ : int -> int
     succ x is x+1.
value pred : int -> int
     pred x is x-1.
value prefix + : int -> int -> int
value add_int : int -> int -> int

Addition.

value prefix - : int -> int -> int
value sub_int : int -> int -> int

Subtraction.

value prefix * : int -> int -> int
value mult_int : int -> int -> int

Multiplication.

value prefix / : int -> int -> int
value div_int : int -> int -> int
value prefix quo : int -> int -> int

Integer division. Raise Division_by_zero if the second argument is 0. Give unpredictable results if either argument is negative.

value prefix mod : int -> int -> int

Remainder. Raise Division_by_zero if the second argument is 0. Give unpredictable results if either argument is negative.

value eq_int : int -> int -> bool

Integer equality. Equivalent to generic equality, just faster.

value neq_int : int -> int -> bool

Negation of eq_int.

Chapter 13. The core library 121

value lt_int : int -> int -> bool
value gt_int : int -> int -> bool
value le_int : int -> int -> bool
value ge_int : int -> int -> bool

Usual comparisons between integers.

value abs : int -> int

Return the absolute value of the argument.

value max_int : int
value min_int : int

The greatest and smallest integer values.

Bitwise operations

value prefix land : int -> int -> int

Bitwise logical and.

value prefix lor : int -> int -> int

Bitwise logical or.

value prefix lxor : int -> int -> int

Bitwise logical exclusive or.

value lnot : int -> int

Bitwise complement

value prefix lsl : int -> int -> int
value lshift_left : int -> int -> int

n lsl m, or equivalently lshift_left n m, shifts n to the left by m bits.

value prefix lsr : int -> int -> int

n lsr m shifts n to the right by m bits. This is a logical shift: zeroes are inserted regardless of sign.

value prefix asr : int -> int -> int
value lshift_right : int -> int -> int

n asr m, or equivalently lshift_right n m, shifts n to the right by m bits. This is an arithmetic shift: the sign bit is replicated.

Conversion functions

value string_of_int : int -> string

Convert the given integer to its decimal representation.

value int_of_string : string -> int

Convert the given string to an integer, in decimal (by default) or in hexadecimal, octal or binary if the string begins with 0x, 0o or 0b. Raise Failure "int_of_string" if the given string is not a valid representation of an integer.

13.11 io: buffered input and output

type in_channel
type out_channel

The abstract types of input channels and output channels.

exception End_of_file

Raised when an operation cannot complete, because the end of the file has been reached.

value stdin : in_channel
value std_in : in_channel
value stdout : out_channel
value std_out : out_channel
value stderr : out_channel
value std_err : out_channel

The standard input, standard output, and standard error output for the process. std_in, std_out and std_err are respectively synonymous with stdin, stdout and stderr.

value exit : int -> 'a

Flush all pending writes on std_out and std_err, and terminate the process, returning the given status code to the operating system (usually 0 to indicate no errors, and a small positive integer to indicate failure.) This function should be called at the end of all standalone programs that output results on std_out or std_err; otherwise, the program may appear to produce no output, or its output may be truncated.

Output functions on standard output

value print_char : char -> unit

Print the character on standard output.

Chapter 13. The core library 123

value print_string : string -> unit

Print the string on standard output.

value print_int : int -> unit

Print the integer, in decimal, on standard output.

value print_float : float -> unit

Print the floating-point number, in decimal, on standard output.

value print_endline : string -> unit

Print the string, followed by a newline character, on standard output.

value print_newline : unit -> unit

Print a newline character on standard output, and flush standard output. This can be used to simulate line buffering of standard output.

Output functions on standard error

value prerr_char : char -> unit

Print the character on standard error.

value prerr_string : string -> unit

Print the string on standard error.

value prerr_int : int -> unit

Print the integer, in decimal, on standard error.

value prerr_float : float -> unit

Print the floating-point number, in decimal, on standard error.

value prerr_endline : string -> unit

Print the string, followed by a newline character on standard error and flush standard error.

Input functions on standard input

value read_line : unit -> string

Flush standard output, then read characters from standard input until a newline character is encountered. Return the string of all characters read, without the newline character at the end.

value read_int : unit -> int

Flush standard output, then read one line from standard input and convert it to an integer. Raise Failure "int_of_string" if the line read is not a valid representation of an integer.

value read_float : unit -> float

Flush standard output, then read one line from standard input and convert it to a floating-point number. The result is unspecified if the line read is not a valid representation of a floating-point number.

General output functions

value open_out : string -> out_channel

Open the named file for writing, and return a new output channel on that file, positionned at the beginning of the file. The file is truncated to zero length if it already exists. It is created if it does not already exists. Raise sys__Sys_error if the file could not be opened.

value open_out_bin : string -> out_channel

Same as open_out, but the file is opened in binary mode, so that no translation takes place during writes. On operating systems that do not distinguish between text mode and binary mode, this function behaves like open_out.

value open_out_gen : sys__open_flag list -> int -> string -> out_channel
open_out_gen mode rights filename opens the file named filename for writing, as

above. The extra argument mode specify the opening mode (see sys__open). The extra argument rights specifies the file permissions, in case the file must be created (see sys__open). open_out and open_out_bin are special cases of this function.

value open_descriptor_out : int -> out_channel

open_descriptor_out fd returns a buffered output channel writing to the file descriptor fd. The file descriptor fd must have been previously opened for writing, else the behavior is undefined.

value flush : out_channel -> unit

Flush the buffer associated with the given output channel, performing all pending writes on that channel. Interactive programs must be careful about flushing std_out and std_err at the right time.

Chapter 13. The core library 125

value output_char : out_channel -> char -> unit

Write the character on the given output channel.

value output_string : out_channel -> string -> unit

Write the string on the given output channel.

value output : out_channel -> string -> int -> int -> unit

output chan buff ofs len writes len characters from string buff, starting at offset ofs,

to the output channel chan. Raise Invalid_argument "output" if ofs and len do not

designate a valid substring of buff.

value output_byte : out_channel -> int -> unit

Write one 8-bit integer (as the single character with that code) on the given output channel. The given integer is taken modulo 256.

value output_binary_int : out_channel -> int -> unit

Write one integer in binary format on the given output channel. The only reliable way to read it back is through the input_binary_int function. The format is compatible across all machines for a given version of Caml Light.

value output_value : out_channel -> 'a -> unit

Write the representation of a structured value of any type to a channel. Circularities and sharing inside the value are detected and preserved. The object can be read back, by the function input_value. The format is compatible across all machines for a given version of Caml Light.

value output_compact_value : out_channel -> 'a -> unit

Same as output_value, but uses a different format, which occupies less space on the file, but takes more time to generate and read back.

value seek_out : out_channel -> int -> unit

seek_out chan pos sets the current writing position to pos for channel chan. This works only for regular files. On files of other kinds (such as terminals, pipes and sockets), the behavior is unspecified.

value pos_out : out_channel -> int

Return the current writing position for the given channel.

value out_channel_length : out_channel -> int

Return the total length (number of characters) of the given channel. This works only for regular files. On files of other kinds, the result is meaningless.

value close_out : out_channel -> unit

Close the given channel, flushing all buffered write operations. The behavior is unspecified if any of the functions above is called on a closed channel.

General input functions

value open_in : string -> in_channel

Open the named file for reading, and return a new input channel on that file, positionned at the beginning of the file. Raise sys__Sys_error if the file could not be opened.

value open_in_bin : string -> in_channel

Same as open_in, but the file is opened in binary mode, so that no translation takes place during reads. On operating systems that do not distinguish between text mode and binary mode, this function behaves like open_in.

value open_in_gen : sys__open_flag list -> int -> string -> in_channel
open_in_gen mode rights filename opens the file named filename for reading, as above.

The extra arguments mode and rights specify the opening mode and file permissions (see sys__open). open_in and open_in_bin are special cases of this function.

value open_descriptor_in : int -> in_channel

open_descriptor_in fd returns a buffered input channel reading from the file descriptor fd. The file descriptor fd must have been previously opened for reading, else the behavior is undefined.

value input_char : in_channel -> char

Read one character from the given input channel. Raise End_of_file if there are no more characters to read.

value input_line : in_channel -> string

Read characters from the given input channel, until a newline character is encountered. Return the string of all characters read, without the newline character at the end. Raise End_of_file if the end of the file is reached at the beginning of line.

value input : in_channel -> string -> int -> int -> int

input chan buff ofs len attempts to read len characters from channel chan, storing them in string buff, starting at character number ofs. It returns the actual number of characters read, between 0 and len (inclusive). A return value of 0 means that the end of file was reached. A return value between 0 and len exclusive means that no more characters were available at that time; input must be called again to read the remaining characters, if desired. Exception Invalid_argument "input" is raised if ofs and len do not designate a valid substring of buff.

value really_input : in_channel -> string -> int -> int -> unit

really_input chan buff ofs len reads len characters from channel chan, storing them in string buff, starting at character number ofs. Raise End_of_file if the end of file is reached before len characters have been read. Raise Invalid_argument "really_input" if ofs and len do not designate a valid substring of buff.

Chapter 13. The core library 127

value input_byte : in_channel -> int

Same as input_char, but return the 8-bit integer representing the character. Raise End_of_file if an end of file was reached.

value input_binary_int : in_channel -> int

Read an integer encoded in binary format from the given input channel. See output_binary_int. Raise End_of_file if an end of file was reached while reading the integer.

value input_value : in_channel -> 'a

Read the representation of a structured value, as produced by output_value or output_compact_value, and return the corresponding value. This is not type-safe. The type of the returned object is not 'a properly speaking: the returned object has one unique type, which cannot be determined at compile-time. The programmer should explicitly give the expected type of the returned value, using the following syntax: (input_value chan : type). The behavior is unspecified if the object in the file does not belong to the given type.

value seek_in : in_channel -> int -> unit

seek_in chan pos sets the current reading position to pos for channel chan. This works only for regular files. On files of other kinds, the behavior is unspecified.

value pos_in : in_channel -> int

Return the current reading position for the given channel.

value in_channel_length : in_channel -> int

Return the total length (number of characters) of the given channel. This works only for regular files. On files of other kinds, the result is meaningless.

value close_in : in_channel -> unit

Close the given channel. Anything can happen if any of the functions above is called on a closed channel.

13.12 list: operations on lists

value list_length : 'a list -> int

Return the length (number of elements) of the given list.

value prefix @ : 'a list -> 'a list -> 'a list

List concatenation.

value hd : 'a list -> 'a

Return the first element of the given list. Raise Failure "hd" if the list is empty.

value tl : 'a list -> 'a list

Return the given list without its first element. Raise Failure "tl" if the list is empty.

value rev : 'a list -> 'a list

List reversal.

value map : ('a -> 'b) -> 'a list -> 'b list

map f [a1; ...; an] applies function f to a1, ..., an, and builds the list [f a1; ...; f an] with the results returned by f.

value do_list : ('a -> unit) -> 'a list -> unit

do_list f [a1; ...; an] applies function f in turn to a1; ...; an, discarding all the

results. It is equivalent to begin f a1; f a2; ...; f an; () end.
value it_list : ('a -> 'b -> 'a) -> 'a -> 'b list -> 'a
it_list f a [b1; ...; bn] is f (... (f (f a b1) b2) ...) bn.
value list_it : ('a -> 'b -> 'b) -> 'a list -> 'b -> 'b
list_it f [a1; ...; an] b is f a1 (f a2 (... (f an b) ...)).
value map2 : ('a -> 'b -> 'c) -> 'a list -> 'b list -> 'c list
map2 f [a1; ...; an] [b1; ...; bn] is [f a1 b1; ...; f an bn]. Raise

Invalid_argument "map2" if the two lists have different lengths.

value do_list2 : ('a -> 'b -> unit) -> 'a list -> 'b list -> unit
do_list2 f [a1; ...; an] [b1; ...; bn] calls in turn f a1 b1; ...; f an bn,

discarding the results. Raise Invalid_argument "do_list2" if the two lists have different lengths.

value it_list2 : ('a -> 'b -> 'c -> 'a) -> 'a -> 'b list -> 'c list -> 'a
it_list2 f a [b1; ...; bn] [c1; ...; cn] is
f (... (f (f a b1 c1) b2 c2) ...) bn cn. Raise Invalid_argument "it_list2" if

the two lists have different lengths.

value list_it2 : ('a -> 'b -> 'c -> 'c) -> 'a list -> 'b list -> 'c -> 'c
list_it2 f [a1; ...; an] [b1; ...; bn] c is
f a1 b1 (f a2 b2 (... (f an bn c) ...)). Raise Invalid_argument "list_it2" if

the two lists have different lengths.

value flat_map : ('a -> 'b list) -> 'a list -> 'b list
flat_map f [l1; ...; ln] is (f l1) @ (f l2) @ ... @ (f ln).
value for_all : ('a -> bool) -> 'a list -> bool
for_all p [a1; ...; an] is (p a1) & (p a2) & ... & (p an).
value exists : ('a -> bool) -> 'a list -> bool
exists p [a1; ...; an] is (p a1) or (p a2) or ... or (p an).
value mem : 'a -> 'a list -> bool

mem a l is true if and only if a is structurally equal (see module eq) to an element of l.

value memq : 'a -> 'a list -> bool

memq a l is true if and only if a is physically equal (see module eq) to an element of l.

value except : 'a -> 'a list -> 'a list

except a l returns the list l where the first element structurally equal to a has been removed. The list l is returned unchanged if it does not contain a.

value exceptq : 'a -> 'a list -> 'a list

Same as except, with physical equality instead of structural equality.

value subtract : 'a list -> 'a list -> 'a list

subtract l1 l2 returns the list l1 where all elements structurally equal to one of the elements of l2 have been removed.

value union : 'a list -> 'a list -> 'a list

union l1 l2 appends before list l2 all the elements of list l1 that are not structurally equal to an element of l2.

value intersect : 'a list -> 'a list -> 'a list

intersect l1 l2 returns the list of the elements of l1 that are structurally equal to an element of l2.

value index : 'a -> 'a list -> int

index a l returns the position of the first element of list l that is structurally equal to a. The head of the list has position 0. Raise Not_found if a is not present in l.

value assoc : 'a -> ('a * 'b) list -> 'b

assoc a l returns the value associated with key a in the list of pairs l. That is, assoc a [ ...; (a,b); ...] = b if (a,b) is the leftmost binding of a in list l. Raise Not_found if there is no value associated with a in the list l.

value assq :  'a -> ('a * 'b) list -> 'b

Same as assoc, but use physical equality instead of structural equality to compare keys.

value mem_assoc : 'a -> ('a * 'b) list -> bool

Same as assoc, but simply return true if a binding exists, and false if no bindings exist for the given key.

13.13 pair: operations on pairs

value fst : 'a * 'b -> 'a

Return the first component of a pair.

value snd : 'a * 'b -> 'b

Return the second component of a pair.

value split : ('a * 'b) list -> 'a list * 'b list

Transform a list of pairs into a pair of lists: split [(a1,b1); ...; (an,bn)] is

([a1; ...; an], [b1; ...; bn])
value combine : 'a list * 'b list -> ('a * 'b) list

Transform a pair of lists into a list of pairs: combine ([a1; ...; an], [b1; ...; bn]) is

[(a1,b1); ...; (an,bn)]. Raise Invalid_argument "combine" if the two lists have

different lengths.

value map_combine : ('a * 'b -> 'c) -> 'a list * 'b list -> 'c list
map_combine f ([a1; ...; an], [b1; ...; bn]) is [f (a1, b1); ...; f (an, bn)].

Raise invalid_argument "map_combine" if the two lists have different lengths.

value do_list_combine : ('a * 'b -> unit) -> 'a list * 'b list -> unit
do_list_combine f ([a1; ...; an], [b1; ...; bn]) calls in turn

f (a1, b1); ...; f (an, bn), discarding the results. Raise

Invalid_argument "do_list_combine" if the two lists have different lengths.

13.14 ref: operations on references

type 'a ref = ref of mutable 'a

The type of references (mutable indirection cells) containing a value of type 'a.

value prefix ! : 'a ref -> 'a

!r returns the current contents of reference r. Could be defined as fun (ref x) -> x.

value prefix := : 'a ref -> 'a -> unit

r := a stores the value of a in reference r.

value incr : int ref -> unit

Increment the integer contained in the given reference. Could be defined as

fun r -> r := succ !r.
value decr : int ref -> unit

Decrement the integer contained in the given reference. Could be defined as

fun r -> r := pred !r.

Chapter 13. The core library 131

13.15 stream: operations on streams

type 'a stream

The type of streams containing values of type 'a.

exception Parse_failure

Raised by parsers when none of the first component of the stream patterns is accepted

exception Parse_error

Raised by parsers when the first component of a stream pattern is accepted, but one of the following components is rejected

value stream_next : 'a stream -> 'a

stream_next s returns the first element of stream s, and removes it from the stream. Raise Parse_failure if the stream is empty.

value stream_from : (unit -> 'a) -> 'a stream

stream_from f returns the stream which fetches its terminals using the function f. This function could be defined as:

    let rec stream_from f = [< 'f(); stream_from f >]

but is implemented more efficiently.

value stream_of_string : string -> char stream

stream_of_string s returns the stream of the characters in string s.

value stream_of_channel : in_channel -> char stream

stream_of_channel ic returns the stream of characters read on channel ic.

value do_stream : ('a -> unit) -> 'a stream -> unit

do_stream f s scans the whole stream s, applying the function f in turn to each terminal encountered

value stream_check : ('a -> bool) -> 'a stream -> 'a

stream_check p returns the parser which returns the first terminal of the stream if the predicate p returns true on this terminal, and raises Parse_failure otherwise.

value end_of_stream : 'a stream -> unit

Return () iff the stream is empty, and raise Parse_failure otherwise.

value stream_get : 'a stream -> 'a * 'a stream

stream_get s return the first element of the stream s, and a stream containing the remaining elements of s. Raise Parse_failure if the stream is empty. The stream s is not modified. This function makes it possible to access a stream non-destructively.

13.16 string: string operations

value string_length : string -> int

Return the length (number of characters) of the given string.

value nth_char : string -> int -> char

nth_char s n returns character number n in string s. The first character is character number 0. The last character is character number string_length s - 1. Raise

Invalid_argument "nth_char" if n is ouside the range 0 to (string_length s - 1). You

can also write s.[n] instead of nth_char s n.

value set_nth_char : string -> int -> char -> unit

set_nth_char s n c modifies string s in place, replacing the character number n by c.

Raise Invalid_argument "set_nth_char" if n is ouside the range 0 to
(string_length s - 1). You can also write s.[n] <- c instead of set_nth_char s n c.
value prefix ^: string -> string -> string

s1 ^s2 returns a fresh string containing the concatenation of the strings s1 and s2.

value concat : string list -> string

Return a fresh string containing the concatenation of all the strings in the argument list.

value sub_string : string -> int -> int -> string

sub_string s start len returns a fresh string of length len, containing the characters

number start to start + len - 1 of string s. Raise Invalid_argument "sub_string" if

start and len do not designate a valid substring of s; that is, if start < 0, or len < 0, or

start + len > string_length s.
value create_string : int -> string

create_string n returns a fresh string of length n. The string initially contains arbitrary characters.

value make_string : int -> char -> string

make_string n c returns a fresh string of length n, filled with the character c.

value fill_string : string -> int -> int -> char -> unit

fill_string s start len c modifies string s in place, replacing the characters number

start to start + len - 1 by c. Raise Invalid_argument "fill_string" if start and

len do not designate a valid substring of s.

value blit_string : string -> int -> string -> int -> int -> unit

blit_string s1 o1 s2 o2 len copies len characters from string s1, starting at character number o1, to string s2, starting at character number o2. It works correctly even if s1 and s2 are the same string, and the source and destination chunks overlap. Raise Invalid_argument "blit_string" if o1 and len do not designate a valid substring of s1, or if o2 and len do not designate a valid substring of s2.

value replace_string : string -> string -> int -> unit

replace_string dest src start copies all characters from the string src into the string dst, starting at character number start in dst. Raise

Invalid_argument "replace_string" if copying would overflow string dest.
value eq_string : string -> string -> bool
value neq_string : string -> string -> bool
value le_string : string -> string -> bool
value lt_string : string -> string -> bool
value ge_string : string -> string -> bool
value gt_string : string -> string -> bool

Comparison functions (lexicographic ordering) between strings.

value compare_strings : string -> string -> int

General comparison between strings. compare_strings s1 s2 returns 0 if s1 and s2 are equal, or else -2 if s1 is a prefix of s2, or 2 if s2 is a prefix of s1, or else -1 if s1 is lexicographically before s2, or 1 if s2 is lexicographically before s1.

value string_for_read : string -> string

Return a copy of the argument, with special characters represented by escape sequences, following the lexical conventions of Caml Light.

value index_char: string -> char -> int

index_char s c returns the position of the leftmost occurrence of character c in string s. Raise Not_found if c does not occur in s.

value rindex_char: string -> char -> int

rindex_char s c returns the position of the rightmost occurrence of character c in string s. Raise Not_found if c does not occur in s.

value index_char_from: string -> int -> char -> int
value rindex_char_from: string -> int -> char -> int

Same as index_char and rindex_char, but start searching at the character position given as second argument. index_char s c is equivalent to index_char_from s 0 c, and

rindex_char s c to rindex_char_from s (string_length s - 1) c.

13.17 vect: operations on vectors

value vect_length : 'a vect -> int

Return the length (number of elements) of the given vector.

value vect_item : 'a vect -> int -> 'a

vect_item v n returns the element number n of vector v. The first element has number 0.

The last element has number vect_length v - 1. Raise Invalid_argument "vect_item"

if n is outside the range 0 to (vect_length v - 1). You can also write v.(n) instead of

vect_item v n.
value vect_assign : 'a vect -> int -> 'a -> unit

vect_assign v n x modifies vector v in place, replacing element number n with x. Raise

Invalid_argument "vect_assign" if n is outside the range 0 to vect_length v - 1. You
can also write v.(n) <- x instead of vect_assign v n x.
value make_vect : int -> 'a -> 'a vect

make_vect n x returns a fresh vector of length n, initialized with x. All the elements of this new vector are initially physically equal to x (see module eq). Consequently, if x is mutable, it is shared among all elements of the vector, and modifying x through one of the vector entries will modify all other entries at the same time.

value make_matrix : int -> int -> 'a -> 'a vect vect

make_matrix dimx dimy e returns a two-dimensional array (a vector of vectors) with first dimension dimx and second dimension dimy. All the elements of this new matrix are initially physically equal to e. The element (x,y) of a matrix m is accessed with the notation

m.(x).(y).
value init_vect : int -> (int -> 'a) -> 'a vect

init_vect n f returns a fresh array of length n, with element number i equal to f i.

value concat_vect : 'a vect -> 'a vect -> 'a vect

concat_vect v1 v2 returns a fresh vector containing the concatenation of vectors v1 and

v2.
value sub_vect : 'a vect -> int -> int -> 'a vect

sub_vect v start len returns a fresh vector of length len, containing the elements

number start to start + len - 1 of vector v. Raise Invalid_argument "sub_vect" if

start and len do not designate a valid subvector of v; that is, if start < 0, or len < 0, or

start + len > vect_length v.
value copy_vect : 'a vect -> 'a vect

copy_vect v returns a copy of v, that is, a fresh vector containing the same elements as v.

Chapter 13. The core library 135

value fill_vect : 'a vect -> int -> int -> 'a -> unit

fill_vect v ofs len x modifies the vector v in place, storing x in elements number ofs

to ofs + len - 1. Raise Invalid_argument "fill_vect" if ofs and len do not designate

a valid subvector of v.

value blit_vect : 'a vect -> int -> 'a vect -> int -> int -> unit

blit_vect v1 o1 v2 o2 len copies len elements from vector v1, starting at element number o1, to vector v2, starting at element number o2. It works correctly even if v1 and v2 are the same vector, and the source and destination chunks overlap. Raise Invalid_argument "blit_vect" if o1 and len do not designate a valid subvector of v1, or if o2 and len do not designate a valid subvector of v2.

value list_of_vect : 'a vect -> 'a list

list_of_vect v returns the list of all the elements of v, that is:

[v.(0); v.(1); ...; v.(vect_length v - 1)].
value vect_of_list : 'a list -> 'a vect

vect_of_list l returns a fresh vector containing the elements of l.

value map_vect : ('a -> 'b) -> 'a vect -> 'b vect

map_vect f v applies function f to all the elements of v, and builds a vector with the

results returned by f: [| f v.(0); f v.(1); ...; f v.(vect_length v - 1) |].
value map_vect_list : ('a -> 'b) -> 'a vect -> 'b list

map_vect_list f v applies function f to all the elements of v, and builds a list with the

results returned by f: [ f v.(0); f v.(1); ...; f v.(vect_length v - 1) ].
value do_vect : ('a -> unit) -> 'a vect -> unit

do_vect f v applies function f in turn to all the elements of v, discarding all the results:

f v.(0); f v.(1); ...; f v.(vect_length v - 1); ().

136

The standard library

This chapter describes the functions provided by the Caml Light standard library. Just as the modules from the core library, the modules from the standard library are automatically linked with the user’s object code files by the camlc command. Hence, the globals defined by these libraries can be used in standalone programs without having to add any .zo file on the command line for the linking phase. Similarly, in interactive use, these globals can be used in toplevel phrases without having to load any .zo file in memory.

Unlike the modules from the core library, the modules from the standard library are not automatically “opened” when a compilation starts, or when the toplevel system is launched. Hence it is necessary to use qualified identifiers to refer to the functions provided by these modules, or to add #open directives.

Conventions

For easy reference, the modules are listed below in alphabetical order of module names. For each module, the declarations from its interface file are printed one by one in typewriter font, followed by a short comment. All modules and the identifiers they export are indexed at the end of this report.

14.1 arg: parsing of command line arguments

This module provides a general mechanism for extracting options and arguments from the command line to the program.

Syntax of command lines: A keyword is a character string starting with a -. An option is a keyword alone or followed by an argument. There are four types of keywords: Unit, String, Int, and Float. Unit keywords do not take an argument. String, Int, and Float keywords take the following word on the command line as an argument. Arguments not preceded by a keyword are called anonymous arguments.

Examples (cmd is assumed to be the command name):

cmd -flag (a unit option) cmd -int 1 (an int option with argument 1) cmd -string foobar (a string option with argument "foobar") cmd -float 12.34 (a float option with argument 12.34) cmd 1 2 3 (three anonymous arguments: "1", "2", and "3")

cmd 1 2 -flag 3 -string bar 4

(four anonymous arguments, a unit option, and a string option with argument "bar")

type spec =
String of (string -> unit)
| Int of (int -> unit)
| Unit of (unit -> unit)
| Float of (float -> unit)

The concrete type describing the behavior associated with a keyword.

value parse : (string * spec) list -> (string -> unit) -> unit

parse speclist anonfun parses the command line, calling the functions in speclist whenever appropriate, and anonfun on anonymous arguments. The functions are called in the same order as they appear on the command line. The strings in the (string * spec) list are keywords and must start with a -, else they are ignored. For the user to be able to specify anonymous arguments starting with a -, include for example

("--", String anonfun) in speclist.
exception Bad of string

Functions in speclist or anonfun can raise Bad with an error message to reject invalid arguments.

14.2 baltree: basic balanced binary trees

This module implements balanced ordered binary trees. All operations over binary trees are applicative (no side-effects). The set and map modules are based on this module. This modules gives a more direct access to the internals of the binary tree implementation than the set and map abstractions, but is more delicate to use and not as safe. For advanced users only.

type 'a t = Empty | Node of 'a t * 'a * 'a t * int

The type of trees containing elements of type 'a. Empty is the empty tree (containing no elements).

type 'a contents = Nothing | Something of 'a

Used with the functions modify and split, to represent the presence or the absence of an element in a tree.

value add: ('a -> int) -> 'a -> 'a t -> 'a t

add f x t inserts the element x into the tree t. f is an ordering function: f y must return 0 if x and y are equal (or equivalent), a negative integer if x is smaller than y, and a positive integer if x is greater than y. The tree t is returned unchanged if it already contains an element equivalent to x (that is, an element y such that f y is 0). The ordering f must be consistent with the orderings used to build t with add, remove, modify or split operations.

value contains: ('a -> int) -> 'a t -> bool

contains f t checks whether t contains an element satisfying f, that is, an element x such that f x is 0. f is an ordering function with the same constraints as for add. It can be coarser (identify more elements) than the orderings used to build t, but must be consistent with them.

value find: ('a -> int) -> 'a t -> 'a

Same as contains, except that find f t returns the element x such that f x is 0, or raises Not_found if none has been found.

value remove: ('a -> int) -> 'a t -> 'a t

remove f t removes one element x of t such that f x is 0. f is an ordering function with the same constraints as for add. t is returned unchanged if it does not contain any element satisfying f. If several elements of t satisfy f, only one is removed.

value modify: ('a -> int) -> ('a contents -> 'a contents) -> 'a t -> 'a t

General insertion/modification/deletion function. modify f g t searchs t for an element x satisfying the ordering function f. If one is found, g is applied to Something x; if g returns Nothing, the element x is removed; if g returns Something y, the element y replaces x in the tree. (It is assumed that x and y are equivalent, in particular, that f y is 0.) If the tree does not contain any x satisfying f, g is applied to Nothing; if it returns Nothing, the tree is returned unchanged; if it returns Something x, the element x is inserted in the tree. (It is assumed that f x is 0.) The functions add and remove are special cases of modify, slightly more efficient.

value split: ('a -> int) -> 'a t -> 'a t * 'a contents * 'a t

split f t returns a triple (less, elt, greater) where less is a tree containing all elements x of t such that f x is negative, greater is a tree containing all elements x of t such that f x is positive, and elt is Something x if t contains an element x such that f x is 0, and Nothing otherwise.

value compare: ('a -> 'a -> int) -> 'a t -> 'a t -> int

Compare two trees. The first argument f is a comparison function over the tree elements: f e1 e2 is zero if the elements e1 and e2 are equal, negative if e1 is smaller than e2, and positive if e1 is greater than e2. compare f t1 t2 compares the fringes of t1 and t2 by lexicographic extension of f.

14.3 filename: operations on file names

value current_dir_name : string

The conventional name for the current directory (e.g. . in Unix).

value concat : string -> string -> string

concat dir file returns a file name that designates file file in directory dir.

value is_absolute : string -> bool

Return true if the file name is absolute or starts with an explicit reference to the current directory (./ or ../ in Unix), and false if it is relative to the current directory.

value check_suffix : string -> string -> bool

check_suffix name suff returns true if the filename name ends with the suffix suff.

value chop_suffix : string -> string -> string

chop_suffix name suff removes the suffix suff from the filename name. The behavior is undefined if name does not end with the suffix suff.

value basename : string -> string
value dirname : string -> string

Split a file name into directory name / base file name.

concat (dirname name) (basename name) returns a file name which is equivalent to name.

Moreover, after setting the current directory to dirname name (with sys__chdir), references to basename name (which is a relative file name) designate the same file as name before the call to chdir.

14.4 format: pretty printing

This module implements a pretty-printing facility to format text within “pretty-printing boxes”. The pretty-printer breaks lines at specified break hints, and indents lines according to the box structure.

Rule of thumb for casual users: use simple boxes (as obtained by open_box 0); use simple break hints (as obtained by print_cut () that outputs a simple break hint, or by print_space () that ouputs a space indicating a break hint); once a box is opened, display its material with basic printing functions (e. g. print_int

and print_string);

when the material for a box has been printed, call close_box () to close the box; at the end of your routine, evaluate print_newline () to close all remaining boxes and flush the pretty-printer.

You may alternatively consider this module as providing an extension to the printf facility: you can simply add pretty-printing annotations to your regular printf formats, as explained below in the documentation of the function fprintf. The behaviour of pretty-printing commands is unspecified if there is no opened pretty-printing box. Each box opened via one of the open_functions below must be closed using close_box for proper formatting. Otherwise, some of the material printed in the boxes may not be output, or may be formatted incorrectly. In case of interactive use, the system closes all opened boxes and flushes all pending text (as with the print_newline function) after each phrase. Each phrase is therefore executed in the initial state of the pretty-printer.

Boxes

value open_box : int -> unit

open_box d opens a new pretty-printing box with offset d. This box is the general purpose pretty-printing box. Material in this box is displayed “horizontal or vertical”: break hints inside the box may lead to a new line, if there is no more room on the line to print the remainder of the box, or if a new line may lead to a new indentation (demonstrating the indentation of the box). When a new line is printed in the box, d is added to the current indentation.

value close_box : unit -> unit

Close the most recently opened pretty-printing box.

Formatting functions

value print_string : string -> unit

print_string str prints str in the current box.

value print_as : int -> string -> unit

print_as len str prints str in the current box. The pretty-printer formats str as if it were of length len.

value print_int : int -> unit

Print an integer in the current box.

value print_float : float -> unit

Print a floating point number in the current box.

value print_char : char -> unit

Print a character in the current box.

value print_bool : bool -> unit

Print an boolean in the current box.

Break hints

value print_space : unit -> unit

print_space () is used to separate items (typically to print a space between two words). It indicates that the line may be split at this point. It either prints one space or splits the line. It is equivalent to print_break 1 0.

value print_cut : unit -> unit

print_cut () is used to mark a good break position. It indicates that the line may be split at this point. It either prints nothing or splits the line. This allows line splitting at the current point, without printing spaces or adding indentation. It is equivalent to

print_break 0 0.
value print_break : int -> int -> unit

Insert a break hint in a pretty-printing box. print_break nspaces offset indicates that the line may be split (a newline character is printed) at this point, if the contents of the current box does not fit on one line. If the line is split at that point, offset is added to the current indentation. If the line is not split, nspaces spaces are printed.

value print_flush : unit -> unit

Flush the pretty printer: all opened boxes are closed, and all pending text is displayed.

value print_newline : unit -> unit

Equivalent to print_flush followed by a new line.

value force_newline : unit -> unit

Force a newline in the current box. Not the normal way of pretty-printing, you should prefer break hints.

value print_if_newline : unit -> unit

Execute the next formatting command if the preceding line has just been split. Otherwise, ignore the next formatting command.

Margin

value set_margin : int -> unit

set_margin d sets the value of the right margin to d (in characters): this value is used to detect line overflows that leads to split lines. Nothing happens if d is smaller than 2 or bigger than 999999999.

value get_margin : unit -> int

Return the position of the right margin.

Maximum indentation limit

value set_max_indent : int -> unit

set_max_indent d sets the value of the maximum indentation limit to d (in characters): once this limit is reached, boxes are rejected to the left, if they do not fit on the current line. Nothing happens if d is smaller than 2 or bigger than 999999999.

value get_max_indent : unit -> int

Return the value of the maximum indentation limit (in characters).

Formatting depth: maximum number of boxes allowed before ellipsis

value set_max_boxes : int -> unit

set_max_boxes max sets the maximum number of boxes simultaneously opened. Material inside boxes nested deeper is printed as an ellipsis (more precisely as the text returned by get_ellipsis_text ()). Nothing happens if max is not greater than 1.

value get_max_boxes : unit -> int

Return the maximum number of boxes allowed before ellipsis.

value over_max_boxes : unit -> bool

Test the maximum number of boxes allowed have already been opened.

Advanced formatting

value open_hbox : unit -> unit

open_hbox () opens a new pretty-printing box. This box is “horizontal”: the line is not split in this box (new lines may still occur inside boxes nested deeper).

value open_vbox : int -> unit

open_vbox d opens a new pretty-printing box with offset d. This box is “vertical”: every break hint inside this box leads to a new line. When a new line is printed in the box, d is added to the current indentation.

value open_hvbox : int -> unit

open_hvbox d opens a new pretty-printing box with offset d. This box is “horizontal-vertical”: it behaves as an “horizontal” box if it fits on a single line, otherwise it behaves as a “vertical” box. When a new line is printed in the box, d is added to the current indentation.

value open_hovbox : int -> unit

open_hovbox d opens a new pretty-printing box with offset d. This box is “horizontal or vertical”: break hints inside this box may lead to a new line, if there is no more room on the line to print the remainder of the box. When a new line is printed in the box, d is added to the current indentation.

Tabulations

value open_tbox : unit -> unit

Open a tabulation box.

value close_tbox : unit -> unit

Close the most recently opened tabulation box.

value print_tbreak : int -> int -> unit

Break hint in a tabulation box. print_tbreak spaces offset moves the insertion point to the next tabulation (spaces being added to this position). Nothing occurs if insertion point is already on a tabulation mark. If there is no next tabulation on the line, then a newline is printed and the insertion point moves to the first tabulation of the box. If a new line is printed, offset is added to the current indentation.

value set_tab : unit -> unit

Set a tabulation mark at the current insertion point.

value print_tab : unit -> unit
print_tab () is equivalent to print_tbreak (0,0).

Ellipsis

value set_ellipsis_text : string -> unit

Set the text of the ellipsis printed when too many boxes are opened (a single dot, ., by default).

value get_ellipsis_text : unit -> string

Return the text of the ellipsis.

Redirecting formatter output

value set_formatter_out_channel : out_channel -> unit

Redirect the pretty-printer output to the given channel.

value set_formatter_output_functions :
 (string -> int -> int -> unit) -> (unit -> unit) -> unit
set_formatter_output_functions out flush redirects the pretty-printer output to the

functions out and flush. The out function performs the pretty-printer output. It is called with a string s, a start position p, and a number of characters n; it is supposed to output characters p to p+n-1 of s. The flush function is called whenever the pretty-printer is

flushed using print_flush or print_newline.
value get_formatter_output_functions :
   unit -> (string -> int -> int -> unit) * (unit -> unit)

Return the current output functions of the pretty-printer.

Multiple formatted output

type formatter

Abstract data type corresponding to a pretty-printer and all its machinery. Defining new pretty-printers permits the output of material in parallel on several channels. Parameters of the pretty-printer are local to the pretty-printer: margin, maximum indentation limit, maximum number of boxes simultaneously opened, ellipsis, and so on, are specific to each pretty-printer and may be fixed independently. A new formatter is obtained by calling the

     make_formatter function.
value std_formatter : formatter

The standard formatter used by the formatting functions above. It is defined using make_formatter with output function output stdout and flushing function

     fun () -> flush stdout.
value err_formatter : formatter

A formatter to use with formatting functions below for output to standard error. It is defined using make_formatter with output function output stderr and flushing function

     fun () -> flush stderr.
value make_formatter :
        (string -> int -> int -> unit) -> (unit -> unit) -> formatter

make_formatter out flush returns a new formatter that writes according to the output function out, and flushing function flush. Hence, a formatter to out channel oc is returned

     by make_formatter (output oc) (fun () -> flush oc).
value pp_open_hbox : formatter -> unit -> unit
value pp_open_vbox : formatter -> int -> unit
value pp_open_hvbox : formatter -> int -> unit
value pp_open_hovbox : formatter -> int -> unit
value pp_open_box : formatter -> int -> unit
value pp_close_box : formatter -> unit -> unit
value pp_print_string : formatter -> string -> unit
value pp_print_as : formatter -> int -> string -> unit
value pp_print_int : formatter -> int -> unit
value pp_print_float : formatter -> float -> unit
value pp_print_char : formatter -> char -> unit
value pp_print_bool : formatter -> bool -> unit
value pp_print_break : formatter -> int -> int -> unit
value pp_print_cut : formatter -> unit -> unit
value pp_print_space : formatter -> unit -> unit
value pp_force_newline : formatter -> unit -> unit
value pp_print_flush : formatter -> unit -> unit
value pp_print_newline : formatter -> unit -> unit
value pp_print_if_newline : formatter -> unit -> unit

value pp_open_tbox : formatter -> unit -> unit
value pp_close_tbox : formatter -> unit -> unit
value pp_print_tbreak : formatter -> int -> int -> unit
value pp_set_tab : formatter -> unit -> unit
value pp_print_tab : formatter -> unit -> unit
value pp_set_margin : formatter -> int -> unit
value pp_get_margin : formatter -> unit -> int
value pp_set_max_indent : formatter -> int -> unit
value pp_get_max_indent : formatter -> unit -> int
value pp_set_max_boxes : formatter -> int -> unit
value pp_get_max_boxes : formatter -> unit -> int
value pp_over_max_boxes : formatter -> unit -> bool
value pp_set_ellipsis_text : formatter -> string -> unit
value pp_get_ellipsis_text : formatter -> unit -> string
value pp_set_formatter_out_channel : formatter -> out_channel -> unit
value pp_set_formatter_output_functions : formatter ->
   (string -> int -> int -> unit) -> (unit -> unit) -> unit
value pp_get_formatter_output_functions :
   formatter -> unit -> (string -> int -> int -> unit) * (unit -> unit)

The basic functions to use with formatters. These functions are the basic ones: usual functions operating on the standard formatter are defined via partial evaluation of these

primitives. For instance, print_string is equal to pp_print_string std_formatter.
value fprintf : formatter -> ('a, formatter, unit) format -> 'a

fprintf ff format arg1 ... argN formats the arguments arg1 to argN according to the format string format, and outputs the resulting string on the formatter ff. The format is a character string which contains three types of objects: plain characters and conversion specifications as specified in the printf module, and pretty-printing indications. The pretty-printing indication characters are introduced by a @ character, and their meanings are: [: open a pretty-printing box. The type and offset of the box may be optionally specified with the following syntax: the < character, followed by an optional box type indication, then an optional integer offset, and the closing > character. Box type is one of h, v, hv, or hov, which stand respectively for an horizontal, vertical, “horizontal-vertical” and “horizontal or vertical” box. ]: close the most recently opened pretty-printing box. ,: output a good break as with print_cut (). : output a space, as with print_space (). \n: force a newline, as with force_newline (). ;: output a good break as with print_break. The nspaces and offset parameters of the break may be optionally specified with the following syntax: the < character, followed by an integer nspaces value, then an integer offset, and a closing > character. .: flush the pretty printer as with print_newline (). @: a plain @ character.

value printf : ('a, formatter, unit) format -> 'a
Same as fprintf, but output on std_formatter.
value eprintf: ('a, formatter, unit) format -> 'a
Same as fprintf, but output on err_formatter.

14.5 gc: memory management control and statistics

type stat = {
minor_words : int;
promoted_words : int;
major_words : int;
minor_collections : int;
major_collections : int;
heap_words : int;
heap_chunks : int;
live_words : int;
live_blocks : int;
free_words : int;
free_blocks : int;
largest_words : int;
fragments : int
}

The memory management counters are returned in a stat record. All the numbers are computed since the start of the program. The fields of this record are: minor_words Number of words allocated in the minor heap. promoted_words Number of words allocated in the minor heap that survived a minor collection and were moved to the major heap. major_words Number of words allocated in the major heap, including the promoted words. minor_collections Number of minor collections. major_collections Number of major collection cycles, not counting the current cycle. heap_words Total size of the major heap, in words. heap_chunks Number of times the major heap size was increased. live_words Number of words of live data in the major heap, including the header words. live_blocks Number of live objects in the major heap. free_words Number of words in the free list. free_blocks Number of objects in the free list. largest_words Size (in words) of the largest object in the free list. fragments Number of wasted words due to fragmentation. These are 1-words free blocks placed between two live objects. They cannot be inserted in the free list, thus they are not available for allocation.

The total amount of memory allocated by the program is (in words)

minor_words + major_words - promoted_words. Multiply by the word size (4 on a 32-bit

machine, 8 on a 64-bit machine) to get the number of bytes.

type control = {
mutable minor_heap_size : int;
mutable major_heap_increment : int;
mutable space_overhead : int;
mutable verbose : bool
}

The GC parameters are given as a control record. The fields are: minor_heap_size The size (in words) of the minor heap. Changing this parameter will trigger a minor collection. major_heap_increment The minimum number of words to add to the major heap when increasing it. space_overhead The major GC speed is computed from this parameter. This is the percentage of heap space that will be ”wasted” because the GC does not immediatly collect unreachable objects. The GC will work more (use more CPU time and collect objects more eagerly) if space_overhead is smaller. The computation of the GC speed assumes that the amount of live data is constant. verbose This flag controls the GC messages on standard error output.

value stat : unit -> stat

Return the current values of the memory management counters in a stat record.

value print_stat : io__out_channel -> unit

Print the current values of the memory management counters (in human-readable form) into the channel argument.

value get : unit -> control

Return the current values of the GC parameters in a control record.

value set : control -> unit

set r changes the GC parameters according to the control record r. The normal usage is:

       let r = gc__get () in    (* Get the current parameters. *)
         r.verbose <- true;     (* Change some of them. *)
         gc__set r              (* Set the new values. *)
value minor : unit -> unit

Trigger a minor collection.

value major : unit -> unit

Finish the current major collection cycle.

value full_major : unit -> unit

Finish the current major collection cycle and perform a complete new cycle. This will collect all currently unreachable objects.

14.6 genlex: a generic lexical analyzer

This module implements a simple “standard” lexical analyzer, presented as a function from character streams to token streams. It implements roughly the lexical conventions of Caml, but is parameterized by the set of keywords of your language.

type token =
Kwd of string
| Ident of string
| Int of int
| Float of float
| String of string
| Char of char

The type of tokens. The lexical classes are: Int and Float for integer and floating-point numbers; String for string literals, enclosed in double quotes; Char for character literals, enclosed in backquotes; Ident for identifiers (either sequences of letters, digits, underscores and quotes, or sequences of “operator characters” such as +, *, etc); and Kwd for keywords (either identifiers or single “special characters” such as (, }, etc).

value make_lexer: string list -> (char stream -> token stream)

Construct the lexer function. The first argument is the list of keywords. An identifier s is returned as Kwd s if s belongs to this list, and as Ident s otherwise. A special character s is returned as Kwd s if s belongs to this list, and cause a lexical error (exception Parse_error) otherwise. Blanks and newlines are skipped. Comments delimited by (* and *) are skipped as well, and can be nested.

Example: a lexer suitable for a desk calculator is obtained by

           let lexer = make_lexer ["+";"-";"*";"/";"let";"="; "("; ")"]

The associated parser would be a function from token stream to, for instance, int, and would have rules such as:

           let parse_expr = function
                  [< 'Int n >] -> n
                | [< 'Kwd "("; parse_expr n; 'Kwd ")" >] -> n
                | [< parse_expr n1; (parse_remainder n1) n2 >] -> n2
           and parse_remainder n1 = function
                  [< 'Kwd "+"; parse_expr n2 >] -> n1+n2
                | ...

14.7 hashtbl: hash tables and hash functions

Hash tables are hashed association tables, with in-place modification.

type ('a, 'b) t

The type of hash tables from type 'a to type 'b.

value new : int -> ('a,'b) t

new n creates a new, empty hash table, with initial size n. The table grows as needed, so n is just an initial guess. Better results are said to be achieved when n is a prime number.

Raise Invalid_argument "hashtbl__new" if n is less than 1.
value clear : ('a, 'b) t -> unit

Empty a hash table.

value add : ('a, 'b) t -> 'a -> 'b -> unit

add tbl x y adds a binding of x to y in table tbl. Previous bindings for x are not removed, but simply hidden. That is, after performing remove tbl x, the previous binding for x, if any, is restored. (This is the semantics of association lists.)

value find : ('a, 'b) t -> 'a -> 'b

find tbl x returns the current binding of x in tbl, or raises Not_found if no such binding exists.

value find_all : ('a, 'b) t -> 'a -> 'b list

find_all tbl x returns the list of all data associated with x in tbl. The current binding is returned first, then the previous bindings, in reverse order of introduction in the table.

value remove : ('a, 'b) t -> 'a -> unit

remove tbl x removes the current binding of x in tbl, restoring the previous binding if it exists. It does nothing if x is not bound in tbl.

value do_table : ('a -> 'b -> unit) -> ('a, 'b) t -> unit

do_table f tbl applies f to all bindings in table tbl, discarding all the results. f receives the key as first argument, and the associated value as second argument. Each binding is presented exactly once to f. The order in which the bindings are passed to f is unpredictable, except that successive bindings for the same key are presented in reverse chronological order (most recent first).

value do_table_rev : ('a -> 'b -> unit) -> ('a, 'b) t -> unit

Same as do_table, except that successive bindings for the same key are presented in chronological order (oldest first).

The polymorphic hash primitive

value hash : 'a -> int

hash x associates a positive integer to any value of any type. It is guaranteed that if x = y, then hash x = hash y. Moreover, hash always terminates, even on cyclic structures.

value hash_param : int -> int -> 'a -> int

hash_param n m x computes a hash value for x, with the same properties as for hash. The two extra parameters n and m give more precise control over hashing. Hashing performs a depth-first, right-to-left traversal of the structure x, stopping after n meaningful nodes were encountered, or m nodes, meaningful or not, were encountered. Meaningful nodes are: integers; floating-point numbers; strings; characters; booleans; and constant constructors. Larger values of m and n means that more nodes are taken into account to compute the final hash value, and therefore collisions are less likely to happen. However, hashing takes longer. The parameters m and n govern the tradeoff between accuracy and speed.

14.8 lexing: the run-time library for lexers generated by camllex

Lexer buffers

type lexbuf =
{ refill_buff : lexbuf -> unit;
lex_buffer : string;
mutable lex_abs_pos : int;
mutable lex_start_pos : int;
mutable lex_curr_pos : int;
mutable lex_last_pos : int;
mutable lex_last_action : lexbuf -> obj }

The type of lexer buffers. A lexer buffer is the argument passed to the scanning functions defined by the generated scanners. The lexer buffer holds the current state of the scanner, plus a function to refill the buffer from the input.

value create_lexer_channel : in_channel -> lexbuf

Create a lexer buffer on the given input channel. create_lexer_channel inchan returns a lexer buffer which reads from the input channel inchan, at the current reading position.

value create_lexer_string : string -> lexbuf

Create a lexer buffer which reads from the given string. Reading starts from the first character in the string. An end-of-input condition is generated when the end of the string is reached.

value create_lexer : (string -> int -> int) -> lexbuf

Create a lexer buffer with the given function as its reading method. When the scanner needs more characters, it will call the given function, giving it a character string s and a character count n. The function should put n characters or less in s, starting at character number 0, and return the number of characters provided. A return value of 0 means end of input.

Functions for lexer semantic actions

The following functions can be called from the semantic actions of lexer definitions (the ML code enclosed in braces that computes the value returned by lexing functions). They give access to the character string matched by the regular expression associated with the semantic action. These functions must be applied to the argument lexbuf, which, in the code generated by camllex, is bound to the lexer buffer passed to the parsing function.

value get_lexeme : lexbuf -> string

get_lexeme lexbuf returns the string matched by the regular expression.

value get_lexeme_char : lexbuf -> int -> char

get_lexeme_char lexbuf i returns character number i in the matched string.

value get_lexeme_start : lexbuf -> int

get_lexeme_start lexbuf returns the position in the input stream of the first character of the matched string. The first character of the stream has position 0.

value get_lexeme_end : lexbuf -> int

get_lexeme_end lexbuf returns the position in the input stream of the character following the last character of the matched string. The first character of the stream has position 0.

14.9 map: association tables over ordered types

This module implements applicative association tables, also known as finite maps or dictionaries, given a total ordering function over the keys. All operations over maps are purely applicative (no side-effects). The implementation uses balanced binary trees, and therefore searching and insertion take time logarithmic in the size of the map.

type ('a, 'b) t

The type of maps from type 'a to type 'b.

value empty: ('a -> 'a -> int) -> ('a, 'b) t

The empty map. The argument is a total ordering function over the set elements. This is a two-argument function f such that f e1 e2 is zero if the elements e1 and e2 are equal, f e1 e2 is strictly negative if e1 is smaller than e2, and f e1 e2 is strictly positive if e1 is greater than e2. Examples: a suitable ordering function for type int is prefix -. You can also use the generic structural comparison function eq__compare.

value add: 'a -> 'b -> ('a, 'b) t -> ('a, 'b) t

add x y m returns a map containing the same bindings as m, plus a binding of x to y. Previous bindings for x in m are not removed, but simply hidden: they reappear after performing a remove operation. (This is the semantics of association lists.)

value find:'a -> ('a, 'b) t -> 'b

find x m returns the current binding of x in m, or raises Not_found if no such binding exists.

value remove: 'a -> ('a, 'b) t -> ('a, 'b) t

remove x m returns a map containing the same bindings as m except the current binding for x. The previous binding for x is restored if it exists. m is returned unchanged if x is not bound in m.

value iter: ('a -> 'b -> unit) -> ('a, 'b) t -> unit

iter f m applies f to all bindings in map m, discarding the results. f receives the key as first argument, and the associated value as second argument. The order in which the bindings are passed to f is unspecified. Only current bindings are presented to f: bindings hidden by more recent bindings are not passed to f.

14.10 parsing: the run-time library for parsers generated by camlyacc

value symbol_start : unit -> int
value symbol_end : unit -> int

symbol_start and symbol_end are to be called in the action part of a grammar rule only. They return the position of the string that matches the left-hand side of the rule: symbol_start() returns the position of the first character; symbol_end() returns the position of the last character, plus one. The first character in a file is at position 0.

value rhs_start: int -> int
value rhs_end: int -> int

Same as symbol_start and symbol_end above, but return the position of the string matching the nth item on the right-hand side of the rule, where n is the integer parameter to lhs_start and lhs_end. n is 1 for the leftmost item.

value clear_parser : unit -> unit

Empty the parser stack. Call it just after a parsing function has returned, to remove all pointers from the parser stack to structures that were built by semantic actions during parsing. This is optional, but lowers the memory requirements of the programs.

exception Parse_error

Raised when a parser encounters a syntax error.

14.11 printexc: a catch-all exception handler

value f: ('a -> 'b) -> 'a -> 'b

printexc__f fn x applies fn to x and returns the result. If the evaluation of fn x raises any exception, the name of the exception is printed on standard error output, and the programs aborts with exit code 2. Typical use is printexc__f main (), where main, with type unit->unit, is the entry point of a standalone program, to catch and print stray exceptions. For printexc__f to work properly, the program must have been linked with the -g option.

14.12 printf: formatting printing functions

type ('a, 'b, 'c) format

The type of format strings. 'a is the type of the parameters of the string, 'c is the result type for the printf-style function, and 'b is the type of the first argument given to %a and %t printing functions.

value fprintf: out_channel -> ('a, out_channel, unit) format -> 'a
fprintf outchan format arg1 ... argN formats the arguments arg1 to argN according

to the format string format, and outputs the resulting string on the channel outchan. The format is a character string which contains two types of objects: plain characters, which are simply copied to the output channel, and conversion specifications, each of which causes conversion and printing of one argument. Conversion specifications consist in the % character, followed by optional flags and field widths, followed by one conversion character. The conversion characters and their meanings are: d or i: convert an integer argument to signed decimal u: convert an integer argument to unsigned decimal x: convert an integer argument to unsigned hexadecimal, using lowercase letters. X: convert an integer argument to unsigned hexadecimal, using uppercase letters. s: insert a string argument c: insert a character argument f: convert a floating-point argument to decimal notation, in the style dddd.ddd e or E: convert a floating-point argument to decimal notation, in the style d.ddd e+-dd (mantissa and exponent) g or G: convert a floating-point argument to decimal notation, in style f or e, E (whichever is more compact) b: convert a boolean argument to the string true or false a: user-defined printer. Takes two arguments and apply the first one to outchan (the current output channel) and to the second argument. The first argument must therefore have type out_channel -> 'b -> unit and the second 'b. The output produced by the function is therefore inserted in the output of fprintf at the current point. t: same as %a, but takes only one argument (with type out_channel -> unit) and apply it

to outchan.

Refer to the C library printf function for the meaning of flags and field width specifiers. If too few arguments are provided, printing stops just before converting the first missing argument.

value printf: ('a, out_channel, unit) format -> 'a

Same as fprintf, but output on std_out.

value eprintf: ('a, out_channel, unit) format -> 'a

Same as fprintf, but output on std_err.

value sprintf: ('a, unit, string) format -> 'a

Same as fprintf, except that the result of the formatting is returned as a string instead of being written on a channel.

value fprint: out_channel -> string -> unit

Print the given string on the given output channel, without any formatting. This is the same function as output_string of module io.

value print: string -> unit

Print the given string on std_out, without any formatting. This is the same function as

print_string of module io.
value eprint: string -> unit

Print the given string on std_err, without any formatting. This is the same function as

prerr_string of module io.

14.13 queue: queues

This module implements queues (FIFOs), with in-place modification.

type 'a t

The type of queues containing elements of type 'a.

exception Empty

Raised when take is applied to an empty queue.

value new: unit -> 'a t

Return a new queue, initially empty.

value add: 'a -> 'a t -> unit

add x q adds the element x at the end of the queue q.

value take: 'a t -> 'a

take q removes and returns the first element in queue q, or raises Empty if the queue is empty.

value peek: 'a t -> 'a

peek q returns the first element in queue q, without removing it from the queue, or raises Empty if the queue is empty.

value clear : 'a t -> unit

Discard all elements from a queue.

value length: 'a t -> int

Return the number of elements in a queue.

value iter: ('a -> unit) -> 'a t -> unit

iter f q applies f in turn to all elements of q, from the least recently entered to the most recently entered. The queue itself is unchanged.

Chapter 14. The standard library 157

14.14 random: pseudo-random number generator

value init : int -> unit

Initialize the generator, using the argument as a seed. The same seed will always yield the same sequence of numbers.

value full_init : int vect -> unit

Same as init but takes more data as seed. It is not useful to give more than 55 integers.

value int : int -> int

random__int bound returns a random number between 0 (inclusive) and bound (exclusive). bound must be positive and smaller than 2^{30}.

value float : float -> float

random__float bound returns a random number between 0 (inclusive) and bound (exclusive).

14.15 set: sets over ordered types

This module implements the set data structure, given a total ordering function over the set elements. All operations over sets are purely applicative (no side-effects). The implementation uses balanced binary trees, and is therefore reasonably efficient: insertion and membership take time logarithmic in the size of the set, for instance.

type 'a t

The type of sets containing elements of type 'a.

value empty: ('a -> 'a -> int) -> 'a t

The empty set. The argument is a total ordering function over the set elements. This is a two-argument function f such that f e1 e2 is zero if the elements e1 and e2 are equal, f e1 e2 is strictly negative if e1 is smaller than e2, and f e1 e2 is strictly positive if e1 is greater than e2. Examples: a suitable ordering function for type int is prefix -. You can also use the generic structural comparison function eq__compare.

value is_empty: 'a t -> bool

Test whether a set is empty or not.

value mem: 'a -> 'a t -> bool

mem x s tests whether x belongs to the set s.

value add: 'a -> 'a t -> 'a t

add x s returns a set containing all elements of s, plus x. If x was already in s, s is returned unchanged.

value remove: 'a -> 'a t -> 'a t

remove x s returns a set containing all elements of s, except x. If x was not in s, s is returned unchanged.

value union: 'a t -> 'a t -> 'a t
value inter: 'a t -> 'a t -> 'a t
value diff: 'a t -> 'a t -> 'a t

Union, intersection and set difference.

value equal: 'a t -> 'a t -> bool

equal s1 s2 tests whether the sets s1 and s2 are equal, that is, contain the same elements.

value compare: 'a t -> 'a t -> int

Total ordering between sets. Can be used as the ordering function for doing sets of sets.

value elements: 'a t -> 'a list

Return the list of all elements of the given set. The elements appear in the list in some non-specified order.

value iter: ('a -> unit) -> 'a t -> unit

iter f s applies f in turn to all elements of s, and discards the results. The elements of s are presented to f in a non-specified order.

value fold: ('a -> 'b -> 'b) -> 'a t -> 'b -> 'b
fold f s a computes (f xN ... (f x2 (f x1 a))...), where x1 ... xN are the

elements of s. The order in which elements of s are presented to f is not specified.

value choose: 'a t -> 'a

Return one element of the given set, or raise Not_found if the set is empty. Which element is chosen is not specified, but equal elements will be chosen for equal sets.

14.16 sort: sorting and merging lists

value sort : ('a -> 'a -> bool) -> 'a list -> 'a list

Sort a list in increasing order according to an ordering predicate. The predicate should return true if its first argument is less than or equal to its second argument.

value merge : ('a -> 'a -> bool) -> 'a list -> 'a list -> 'a list

Merge two lists according to the given predicate. Assuming the two argument lists are sorted according to the predicate, merge returns a sorted list containing the elements from the two lists. The behavior is undefined if the two argument lists were not sorted.

Chapter 14. The standard library 159

14.17 stack: stacks

This module implements stacks (LIFOs), with in-place modification.

type 'a t

The type of stacks containing elements of type 'a.

exception Empty

Raised when pop is applied to an empty stack.

value new: unit -> 'a t

Return a new stack, initially empty.

value push: 'a -> 'a t -> unit

push x s adds the element x at the top of stack s.

value pop: 'a t -> 'a

pop s removes and returns the topmost element in stack s, or raises Empty if the stack is empty.

value clear : 'a t -> unit

Discard all elements from a stack.

value length: 'a t -> int

Return the number of elements in a stack.

value iter: ('a -> unit) -> 'a t -> unit

iter f s applies f in turn to all elements of s, from the element at the top of the stack to the element at the bottom of the stack. The stack itself is unchanged.

14.18 sys: system interface

This module provides a simple interface to the operating system.

exception Sys_error of string

Raised by some functions in the sys and io modules, when the underlying system calls fail. The argument to Sys_error is a string describing the error. The texts of the error messages are implementation-dependent, and should not be relied upon to catch specific system errors.

value command_line : string vect

The command line arguments given to the process. The first element is the command name used to invoke the program.

value interactive: bool

True if we’re running under the toplevel system. False if we’re running as a standalone program.

type file_perm == int
value s_irusr : file_perm
value s_iwusr : file_perm
value s_ixusr : file_perm
value s_irgrp : file_perm
value s_iwgrp : file_perm
value s_ixgrp : file_perm
value s_iroth : file_perm
value s_iwoth : file_perm
value s_ixoth : file_perm
value s_isuid : file_perm
value s_isgid : file_perm
value s_irall : file_perm
value s_iwall : file_perm
value s_ixall : file_perm

Access permissions for files. r is reading permission, w is writing permission, x is execution permission. usr means permissions for the user owning the file, grp for the group owning the file, oth for others. isuid and isgid are for set-user-id and set-group-id files, respectively. The remaining are combinations of the permissions above.

type open_flag =
    O_RDONLY                       (* open read-only *)
  | O_WRONLY                       (* open write-only *)
  | O_RDWR                         (* open for reading and writing *)
  | O_APPEND                       (* open for appending *)
  | O_CREAT                        (* create the file if nonexistent *)
  | O_TRUNC                        (* truncate the file to 0 if it exists *)
  | O_EXCL                         (* fails if the file exists *)
  | O_BINARY                       (* open in binary mode *)
  | O_TEXT                         (* open in text mode *)

The commands for open.

value exit : int -> 'a

Terminate the program and return the given status code to the operating system. In contrast with the function exit from module io, this exit function does not flush the standard output and standard error channels.

Chapter 14. The standard library 161

value open : string -> open_flag list -> file_perm -> int

Open a file. The second argument is the opening mode. The third argument is the permissions to use if the file must be created. The result is a file descriptor opened on the file.

value close : int -> unit

Close a file descriptor.

value remove : string -> unit

Remove the given file name from the file system.

value rename : string -> string -> unit

Rename a file. The first argument is the old name and the second is the new name.

value getenv : string -> string

Return the value associated to a variable in the process environment. Raise Not_found if the variable is unbound.

value chdir : string -> unit

Change the current working directory of the process. Note that there is no easy way of getting the current working directory from the operating system.

value system_command : string -> int

Execute the given shell command and return its exit code.

value time : unit -> float

Return the processor time, in seconds, used by the program since the beginning of execution.

exception Break

Exception Break is raised on user interrupt if catch_break is on.

value catch_break : bool -> unit

catch_break governs whether user interrupt terminates the program or raises the Break exception. Call catch_break true to enable raising Break, and catch_break false to let the system terminate the program on user interrupt.

162

The graphics library

This chapter describes the portable graphics primitives that come standard in the implementation of Caml Light on micro-computers.

Unix: On Unix workstations running the X11 windows system, an implementation of the graphics primitives is available in the directory contrib/libgraph in the distribution. See the file README in this directory for information on building and using camlgraph, a toplevel system that includes the graphics primitives, and linking standalone programs with the library. Drawing takes place in a separate window that is created when open_graph is called.

Mac: The graphics primitive are available from the standalone application that runs the toplevel system. They are not available from programs compiled by camlc and run under the MPW shell. Drawing takes place in a separate window, that can be made visible with the “Show graphics window” menu entry.

PC: The graphics primitive are available from the Windows application that runs the toplevel system. They are not available from programs compiled by camlc and run in a DOS command window. Drawing takes place in a separate window.

The screen coordinates are interpreted as shown in the figure below. Notice that the coordinate system used is the same as in mathematics: y increases from the bottom of the screen to the top of the screen, and angles are measured counterclockwise (in degrees). Drawing is clipped to the screen.

[figure: x; y; Screen; size_x(); size_y(); pixel at (x, y); x; y]

Here are the graphics mode specifications supported by open_graph on the various implementations of this library.

Unix: The argument to open_graph has the format "display-name geometry", where display-name is the name of the X-windows display to connect to, and geometry is a standard X-windows geometry specification. The two components are separated by a space. Either can be omitted, or both. Examples:

open_graph "foo:0"

connects to the display foo:0 and creates a window with the default geometry

open_graph "foo:0 300x100+50-0"

connects to the display foo:0 and creates a window 300 pixels wide by 100 pixels tall, at location (50, 0)

open_graph " 300x100+50-0"

connects to the default display and creates a window 300 pixels wide by 100 pixels tall, at location (50, 0)

open_graph ""

connects to the default display and creates a window with the default geometry.

Mac: The argument to open_graph is ignored.

PC: The argument to open_graph has the format "widthxheight" or "widthxheight+x+y", where width and height are the initial dimensions of the graphics windows, and x and y are the position of the upper-left corner of the graphics window. If omitted, (width,height) default to (600,400) and (x,y) default to (10, 10).

15.1 graphics: machine-independent graphics primitives

exception Graphic_failure of string

Raised by the functions below when they encounter an error.

Initializations

value open_graph: string -> unit

Show the graphics window or switch the screen to graphic mode. The graphics window is cleared. The string argument is used to pass optional information on the desired graphics mode, the graphics window size, and so on. Its interpretation is implementation-dependent. If the empty string is given, a sensible default is selected.

value close_graph: unit -> unit

Delete the graphics window or switch the screen back to text mode.

value clear_graph : unit -> unit

Erase the graphics window.

value size_x : unit -> int
value size_y : unit -> int

Return the size of the graphics window. Coordinates of the screen pixels range over 0 .. size_x()-1 and 0 .. size_y()-1. Drawings outside of this rectangle are clipped, without causing an error. The origin (0,0) is at the lower left corner.

Colors

type color == int

A color is specified by its R, G, B components. Each component is in the range 0..255. The three components are packed in an int: 0xRRGGBB, where RR are the two hexadecimal digits for the red component, GG for the green component, BB for the blue component.

value rgb: int -> int -> int -> color

rgb r g b returns the integer encoding the color with red component r, green component g, and blue component b. r, g and b are in the range 0..255.

value set_color : color -> unit

Set the current drawing color.

value black : color
value white : color
value red : color
value green : color
value blue : color
value yellow : color
value cyan : color
value magenta : color

Some predefined colors.

value background: color
value foreground: color

Default background and foreground colors (usually, either black foreground on a white background or white foreground on a black background). clear_graph fills the screen with the background color. The initial drawing color is foreground.

Point and line drawing

value plot : int -> int -> unit

Plot the given point with the current drawing color.

value point_color : int -> int -> color

Return the color of the given point.

value moveto : int -> int -> unit

Position the current point.

value current_point : unit -> int * int

Return the position of the current point.

value lineto : int -> int -> unit

Draw a line with endpoints the current point and the given point, and move the current point to the given point.

value draw_arc : int -> int -> int -> int -> int -> int -> unit

draw_arc x y rx ry a1 a2 draws an elliptical arc with center x,y, horizontal radius rx, vertical radius ry, from angle a1 to angle a2 (in degrees). The current point is unchanged.

value draw_ellipse : int -> int -> int -> int -> unit

draw_ellipse x y rx ry draws an ellipse with center x,y, horizontal radius rx and vertical radius ry. The current point is unchanged.

value draw_circle : int -> int -> int -> unit

draw_circle x y r draws a circle with center x,y and radius r. The current point is unchanged.

value set_line_width : int -> unit

Set the width of points and lines drawn with the functions above. Under X Windows, set_line_width 0 selects a width of 1 pixel and a faster, but less precise drawing algorithm than the one used when set_line_width 1 is specified.

Text drawing

value draw_char : char -> unit
value draw_string : string -> unit

Draw a character or a character string with lower left corner at current position. After drawing, the current position is set to the lower right corner of the text drawn.

value set_font : string -> unit
value set_text_size : int -> unit

Set the font and character size used for drawing text. The interpretation of the arguments to set_font and set_text_size is implementation-dependent.

value text_size : string -> int * int

Return the dimensions of the given text, if it were drawn with the current font and size.

Chapter 15. The graphics library 167

Filling

value fill_rect : int -> int -> int -> int -> unit

fill_rect x y w h fills the rectangle with lower left corner at x,y, width w and heigth h, with the current color.

value fill_poly : (int * int) vect -> unit

Fill the given polygon with the current color. The array contains the coordinates of the vertices of the polygon.

value fill_arc : int -> int -> int -> int -> int -> int -> unit

Fill an elliptical pie slice with the current color. The parameters are the same as for

draw_arc.
value fill_ellipse : int -> int -> int -> int -> unit

Fill an ellipse with the current color. The parameters are the same as for draw_ellipse.

value fill_circle : int -> int -> int -> unit

Fill a circle with the current color. The parameters are the same as for draw_circle.

Images

type image

The abstract type for images, in internal representation. Externally, images are represented as matrices of colors.

value transp : color

In matrices of colors, this color represent a “transparent” point: when drawing the corresponding image, all pixels on the screen corresponding to a transparent pixel in the image will not be modified, while other points will be set to the color of the corresponding point in the image. This allows superimposing an image over an existing background.

value make_image : color vect vect -> image

Convert the given color matrix to an image. Each sub-array represents one horizontal line. All sub-arrays must have the same length; otherwise, exception Graphic_failure is raised.

value dump_image : image -> color vect vect

Convert an image to a color matrix.

value draw_image : image -> int -> int -> unit

Draw the given image with lower left corner at the given point.

value get_image : int -> int -> int -> int -> image

Capture the contents of a rectangle on the screen as an image. The parameters are the same

as for fill_rect.
value create_image : int -> int -> image

create_image w h returns a new image w pixels wide and h pixels tall, to be used in conjunction with blit_image. The initial image contents are random.

value blit_image : image -> int -> int -> unit

blit_image img x y copies screen pixels into the image img, modifying img in-place. The pixels copied are those inside the rectangle with lower left corner at x,y, and width and height equal to those of the image.

Mouse and keyboard events

type status =
{ mouse_x : int;              (* X coordinate of the mouse *)
mouse_y : int;              (* Y coordinate of the mouse *)
button : bool;              (* true if a mouse button is pressed *)
keypressed : bool;          (* true if a key has been pressed *)
key : char }                (* the character for the key pressed *)

To report events.

type event =
Button_down                 (* A mouse button is pressed *)
| Button_up                   (* A mouse button is released *)
| Key_pressed                 (* A key is pressed *)
| Mouse_motion                (* The mouse is moved *)
| Poll                        (* Don't wait; return immediately *)

To specify events to wait for.

value wait_next_event : event list -> status

Wait until one of the events specified in the given event list occurs, and return the status of the mouse and keyboard at that time. If Poll is given in the event list, return immediately with the current status. If the mouse cursor is outside of the graphics window, the mouse_x and mouse_y fields of the event are outside the range 0..size_x()-1, 0..size_y()-1. Keypresses are queued, and dequeued one by one when the Key_pressed event is specified.

Chapter 15. The graphics library 169

Mouse and keyboard polling

value mouse_pos : unit -> int * int

Return the position of the mouse cursor, relative to the graphics window. If the mouse cursor is outside of the graphics window, mouse_pos() returns a point outside of the range

0..size_x()-1, 0..size_y()-1.
value button_down : unit -> bool

Return true if the mouse button is pressed, false otherwise.

value read_key : unit -> char

Wait for a key to be pressed, and return the corresponding character. Keypresses are queued.

value key_pressed : unit -> bool

Return true if a keypress is available; that is, if read_key would not block.

Sound

value sound : int -> int -> unit

sound freq dur plays a sound at frequency freq (in hertz) for a duration dur (in milliseconds). On the Macintosh, the frequency is rounded to the nearest note in the equal-tempered scale.

170

The unix library: Unix system calls

The unix library (distributed in contrib/libunix) makes many Unix system calls and systemrelated library functions available to Caml Light programs. This chapter describes briefly the functions provided. Refer to sections 2 and 3 of the Unix manual for more details on the behavior of these functions.

Not all functions are provided by all Unix variants. If some functions are not available, they will raise Invalid_arg when called.

Programs that use the unix library must be linked in “custom runtime” mode, as follows:

      camlc -custom other options unix.zo other files -lunix

For interactive use of the unix library, run camllight camlunix.

Mac: This library is not available.

PC: This library is not available.

16.1 unix: interface to the Unix system

Error report

type error =
  ENOERR
| EPERM               (* Not owner *)
| ENOENT              (* No such file or directory *)
| ESRCH               (* No such process *)
| EINTR               (* Interrupted system call *)
| EIO                 (* I/O error *)
| ENXIO               (* No such device or address *)
| E2BIG               (* Arg list too long *)
| ENOEXEC             (* Exec format error *)
| EBADF               (* Bad file number *)
| ECHILD              (* No children *)
| EAGAIN              (* No more processes *)

| ENOMEM              (* Not enough core *)
| EACCES              (* Permission denied *)
| EFAULT              (* Bad address *)
| ENOTBLK             (* Block device required *)
| EBUSY               (* Mount device busy *)
| EEXIST              (* File exists *)
| EXDEV               (* Cross-device link *)
| ENODEV              (* No such device *)
| ENOTDIR             (* Not a directory*)
| EISDIR              (* Is a directory *)
| EINVAL              (* Invalid argument *)
| ENFILE              (* File table overflow *)
| EMFILE              (* Too many open files *)
| ENOTTY              (* Not a typewriter *)
| ETXTBSY             (* Text file busy *)
| EFBIG               (* File too large *)
| ENOSPC              (* No space left on device *)
| ESPIPE              (* Illegal seek *)
| EROFS               (* Read-only file system *)
| EMLINK              (* Too many links *)
| EPIPE               (* Broken pipe *)
| EDOM                (* Argument too large *)
| ERANGE              (* Result too large *)
| EWOULDBLOCK         (* Operation would block *)
| EINPROGRESS         (* Operation now in progress *)
| EALREADY            (* Operation already in progress *)
| ENOTSOCK            (* Socket operation on non-socket *)
| EDESTADDRREQ        (* Destination address required *)
| EMSGSIZE            (* Message too long *)
| EPROTOTYPE          (* Protocol wrong type for socket *)
| ENOPROTOOPT         (* Protocol not available *)
| EPROTONOSUPPORT     (* Protocol not supported *)
| ESOCKTNOSUPPORT     (* Socket type not supported *)
| EOPNOTSUPP          (* Operation not supported on socket *)
| EPFNOSUPPORT        (* Protocol family not supported *)
| EAFNOSUPPORT        (* Address family not supported by protocol family *)
| EADDRINUSE          (* Address already in use *)
| EADDRNOTAVAIL       (* Can't assign requested address *)
| ENETDOWN            (* Network is down *)
| ENETUNREACH         (* Network is unreachable *)
| ENETRESET           (* Network dropped connection on reset *)
| ECONNABORTED        (* Software caused connection abort *)
| ECONNRESET          (* Connection reset by peer *)
| ENOBUFS             (* No buffer space available *)
| EISCONN             (* Socket is already connected *)

| ENOTCONN            (* Socket is not connected *)
| ESHUTDOWN           (* Can't send after socket shutdown *)
| ETOOMANYREFS        (* Too many references: can't splice *)
| ETIMEDOUT           (* Connection timed out *)
| ECONNREFUSED        (* Connection refused *)
| ELOOP               (* Too many levels of symbolic links *)
| ENAMETOOLONG        (* File name too long *)
| EHOSTDOWN           (* Host is down *)
| EHOSTUNREACH        (* No route to host *)
| ENOTEMPTY           (* Directory not empty *)
| EPROCLIM            (* Too many processes *)
| EUSERS              (* Too many users *)
| EDQUOT              (* Disc quota exceeded *)
| ESTALE              (* Stale NFS file handle *)
| EREMOTE             (* Too many levels of remote in path *)
| EIDRM               (* Identifier removed *)
| EDEADLK             (* Deadlock condition. *)
| ENOLCK              (* No record locks available. *)
| ENOSYS              (* Function not implemented *)
| EUNKNOWNERR

The type of error codes.

exception Unix_error of error * string * string

Raised by the system calls below when an error is encountered. The first component is the error code; the second component is the function name; the third component is the string parameter to the function, if it has one, or the empty string otherwise.

value error_message : error -> string

Return a string describing the given error code.

value handle_unix_error : ('a -> 'b) -> 'a -> 'b

handle_unix_error f x applies f to x and returns the result. If the exception Unix_error is raised, it prints a message describing the error and exits with code 2.

Interface with the parent process

value environment : unit -> string vect

Return the process environment, as an array of strings with the format “variable=value”.

   See also sys__getenv.

Process handling

type process_status =
WEXITED of int
| WSIGNALED of int * bool
| WSTOPPED of int

The termination status of a process. WEXITED means that the process terminated normally by exit; the argument is the return code. WSIGNALED means that the process was killed by a signal; the first argument is the signal number, the second argument indicates whether a “core dump” was performed. WSTOPPED means that the process was stopped by a signal; the argument is the signal number.

type wait_flag =
WNOHANG
| WUNTRACED

Flags for waitopt and waitpid. WNOHANG means do not block if no child has died yet, but immediately return with a pid equal to 0. WUNTRACED means report also the children that receive stop signals.

value execv : string -> string vect -> unit

execv prog args execute the program in file prog, with the arguments args, and the current process environment.

value execve : string -> string vect -> string vect -> unit

Same as execv, except that the third argument provides the environment to the program executed.

value execvp : string -> string vect -> unit

Same as execv, except that the program is searched in the path.

value fork : unit -> int

Fork a new process. The returned integer is 0 for the child process, the pid of the child process for the parent process.

value wait : unit -> int * process_status

Wait until one of the children processes die, and return its pid and termination status.

value waitopt : wait_flag list -> int * process_status

Same as wait, but takes a list of options to avoid blocking, or also report stopped children. The pid returned is 0 if no child has changed status.

value waitpid : wait_flag list -> int -> int * process_status

Same as waitopt, but waits for the process whose pid is given. Negative pid arguments represent process groups.

Chapter 16. The unix library: Unix system calls 175

value system : string -> process_status

Execute the given command, wait until it terminates, and return its termination status. The string is interpreted by the shell /bin/sh and therefore can contain redirections, quotes, variables, etc. The result WEXITED 127 indicates that the shell couldn’t be executed.

value getpid : unit -> int

Return the pid of the process.

value getppid : unit -> int

Return the pid of the parent process.

value nice : int -> int

Change the process priority. The integer argument is added to the “nice” value. (Higher values of the “nice” value mean lower priorities.) Return the new nice value.

Basic file input/output

type file_descr

The abstract type of file descriptors.

value stdin : file_descr
value stdout : file_descr
value stderr : file_descr

File descriptors for standard input, standard output and standard error.

type open_flag =
O_RDONLY                            (* Open for reading *)
| O_WRONLY                            (* Open for writing *)
| O_RDWR                              (* Open for reading and writing *)
| O_NDELAY                            (* Open in non-blocking mode *)
| O_APPEND                            (* Open for append *)
| O_CREAT                             (* Create if nonexistent *)
| O_TRUNC                             (* Truncate to 0 length if existing *)
| O_EXCL                              (* Fail if existing *)

The flags to open.

type file_perm == int

The type of file access rights.

value open : string -> open_flag list -> file_perm -> file_descr

Open the named file with the given flags. Third argument is the permissions to give to the file if it is created. Return a file descriptor on the named file.

value close : file_descr -> unit

Close a file descriptor.

value read : file_descr -> string -> int -> int -> int

read fd buff start len reads len characters from descriptor fd, storing them in string buff, starting at position ofs in string buff. Return the number of characters actually read.

value write : file_descr -> string -> int -> int -> int

write fd buff start len writes len characters to descriptor fd, taking them from string buff, starting at position ofs in string buff. Return the number of characters actually written.

Interfacing with the standard input/output library (module io).

value in_channel_of_descr : file_descr -> in_channel

Create an input channel reading from the given descriptor.

value out_channel_of_descr : file_descr -> out_channel

Create an output channel writing on the given descriptor.

value descr_of_in_channel : in_channel -> file_descr

Return the descriptor corresponding to an input channel.

value descr_of_out_channel : out_channel -> file_descr

Return the descriptor corresponding to an output channel.

Seeking and truncating

type seek_command =
SEEK_SET
| SEEK_CUR
| SEEK_END

Positioning modes for lseek. SEEK_SET indicates positions relative to the beginning of the file, SEEK_CUR relative to the current position, SEEK_END relative to the end of the file.

value lseek : file_descr -> int -> seek_command -> int

Set the current position for a file descriptor

value truncate : string -> int -> unit

Truncates the named file to the given size.

value ftruncate : file_descr -> int -> unit

Truncates the file corresponding to the given descriptor to the given size.

File statistics

type file_kind =
S_REG                               (* Regular file *)
| S_DIR                               (* Directory *)
| S_CHR                               (* Character device *)
| S_BLK                               (* Block device *)
| S_LNK                               (* Symbolic link *)
| S_FIFO                              (* Named pipe *)
| S_SOCK                              (* Socket *)
type stats =
{ st_dev : int;                       (* Device number *)
st_ino : int;                       (* Inode number *)
st_kind : file_kind;                (* Kind of the file *)
st_perm : file_perm;                (* Access rights *)
st_nlink : int;                     (* Number of links *)
st_uid : int;                       (* User id of the owner *)
st_gid : int;                       (* Group id of the owner *)
st_rdev : int;                      (* Device minor number *)
st_size : int;                      (* Size in bytes *)
st_atime : int;                     (* Last access time *)
st_mtime : int;                     (* Last modification time *)
st_ctime : int }                    (* Last status change time *)

The informations returned by the stat calls.

value stat : string -> stats

Return the information for the named file.

value lstat : string -> stats

Same as stat, but in case the file is a symbolic link, return the information for the link itself.

value fstat : file_descr -> stats

Return the information for the file associated with the given descriptor.

Operations on file names

value unlink : string -> unit

Removes the named file

value rename : string -> string -> unit

rename old new changes the name of a file from old to new.

value link : string -> string -> unit

link source dest creates a hard link named dest to the file named new.

File permissions and ownership

type access_permission =
R_OK                                (* Read permission *)
| W_OK                                (* Write permission *)
| X_OK                                (* Execution permission *)
| F_OK                                (* File exists *)

Flags for the access call.

value chmod : string -> file_perm -> unit

Change the permissions of the named file.

value fchmod : file_descr -> file_perm -> unit

Change the permissions of an opened file.

value chown : string -> int -> int -> unit

Change the owner uid and owner gid of the named file.

value fchown : file_descr -> int -> int -> unit

Change the owner uid and owner gid of an opened file.

value umask : int -> int

Set the process creation mask, and return the previous mask.

value access : string -> access_permission list -> unit

Check that the process has the given permissions over the named file. Raise Unix_error otherwise.

File descriptor hacking

value fcntl_int : file_descr -> int -> int -> int

Interface to fcntl in the case where the argument is an integer. The first integer argument is the command code; the second is the integer parameter.

value fcntl_ptr : file_descr -> int -> string -> int

Interface to fcntl in the case where the argument is a pointer. The integer argument is the command code. A pointer to the string argument is passed as argument to the command.

Chapter 16. The unix library: Unix system calls 179

Directories

value mkdir : string -> file_perm -> unit

Create a directory with the given permissions.

value rmdir : string -> unit

Remove an empty directory.

value chdir : string -> unit

Change the process working directory.

value getcwd : unit -> string

Return the name of the current working directory.

type dir_handle

The type of descriptors over opened directories.

value opendir : string -> dir_handle

Open a descriptor on a directory

value readdir : dir_handle -> string

Return the next entry in a directory. Raise End_of_file when the end of the directory has been reached.

value rewinddir : dir_handle -> unit

Reposition the descriptor to the beginning of the directory

value closedir : dir_handle -> unit

Close a directory descriptor.

Pipes and redirections

value pipe : unit -> file_descr * file_descr

Create a pipe. The first component of the result is opened for reading, that’s the exit to the pipe. The second component is opened for writing, that’s the entrace to the pipe.

value dup : file_descr -> file_descr

Duplicate a descriptor.

value dup2 : file_descr -> file_descr -> unit

dup2 fd1 fd2 duplicates fd1 to fd2, closing fd2 if already opened.

value open_process_in: string -> in_channel
value open_process_out: string -> out_channel
value open_process: string -> in_channel * out_channel

High-level pipe and process management. These functions run the given command in parallel with the program, and return channels connected to the standard input and/or the standard output of the command. The command is interpreted by the shell /bin/sh (cf. system). Warning: writes on channels are buffered, hence be careful to call flush at the right times to ensure correct synchronization.

value close_process_in: in_channel -> process_status
value close_process_out: out_channel -> process_status
value close_process: in_channel * out_channel -> process_status
Close channels opened by open_process_in, open_process_out and open_process,

respectively, wait for the associated command to terminate, and return its termination status.

Symbolic links

value symlink : string -> string -> unit

symlink source dest creates the file dest as a symbolic link to the file source.

value readlink : string -> string

Read the contents of a link.

Named pipes

value mkfifo : string -> file_perm -> unit

Create a named pipe with the given permissions.

Special files

value ioctl_int : file_descr -> int -> int -> int

Interface to ioctl in the case where the argument is an integer. The first integer argument is the command code; the second is the integer parameter.

value ioctl_ptr : file_descr -> int -> string -> int

Interface to ioctl in the case where the argument is a pointer. The integer argument is the command code. A pointer to the string argument is passed as argument to the command.

Polling

value select :
file_descr list -> file_descr list -> file_descr list -> float ->
      file_descr list * file_descr list * file_descr list

Wait until some input/output operations become possible on some channels. The three list arguments are, respectively, a set of descriptors to check for reading (first argument), for writing (second argument), or for exceptional conditions (third argument). The fourth argument is the maximal timeout, in seconds; a negative fourth argument means no timeout (unbounded wait). The result is composed of three sets of descriptors: those ready for reading (first component), ready for writing (second component), and over which an exceptional condition is pending (third component).

Locking

type lock_command =
  F_ULOCK               (* Unlock a region *)
| F_LOCK                (* Lock a region, and block if already locked *)
| F_TLOCK               (* Lock a region, or fail if already locked *)
| F_TEST                (* Test a region for other process' locks *)

Commands for lockf.

value lockf : file_descr -> lock_command -> int -> unit

lockf fd cmd size puts a lock on a region of the file opened as fd. The region starts at the current read/write position for fd (as set by lseek), and extends size bytes forward if size is positive, size bytes backwards if size is negative, or to the end of the file if size is zero.

Signals

type signal =
  SIGHUP              (* hangup *)
| SIGINT              (* interrupt *)
| SIGQUIT             (* quit *)
| SIGILL              (* illegal instruction (not reset when caught) *)
| SIGTRAP             (* trace trap (not reset when caught) *)
| SIGABRT             (* used by abort *)
| SIGEMT              (* EMT instruction *)
| SIGFPE              (* floating point exception *)
| SIGKILL             (* kill (cannot be caught or ignored) *)
| SIGBUS              (* bus error *)
| SIGSEGV             (* segmentation violation *)
| SIGSYS              (* bad argument to system call *)
| SIGPIPE             (* write on a pipe with no one to read it *)

| SIGALRM             (* alarm clock *)
| SIGTERM             (* software termination signal from kill *)
| SIGURG              (* urgent condition on IO channel *)
| SIGSTOP             (* sendable stop signal not from tty *)
| SIGTSTP             (* stop signal from tty *)
| SIGCONT             (* continue a stopped process *)
| SIGCHLD             (* to parent on child stop or exit *)
| SIGIO               (* input/output possible signal *)
| SIGXCPU             (* exceeded CPU time limit *)
| SIGXFSZ             (* exceeded file size limit *)
| SIGVTALRM           (* virtual time alarm *)
| SIGPROF             (* profiling time alarm *)
| SIGWINCH            (* window changed *)
| SIGLOST             (* resource lost (eg, record-lock lost) *)
| SIGUSR1             (* user defined signal 1 *)
| SIGUSR2             (* user defined signal 2 *)

The type of signals.

type signal_handler =
  Signal_default                      (* Default behavior for the signal *)
| Signal_ignore                       (* Ignore the signal *)
| Signal_handle of (unit -> unit)     (* Call the given function
                                         when the signal occurs. *)

The behavior on receipt of a signal

value kill : int -> signal -> unit

Send a signal to the process with the given process id.

value signal : signal -> signal_handler -> unit

Set the behavior to be taken on receipt of the given signal.

value pause : unit -> unit

Wait until a non-ignored signal is delivered.

Time functions

type process_times =
{ tms_utime : float;          (* User time for the process *)
  tms_stime : float;          (* System time for the process *)
  tms_cutime : float;         (* User time for the children processes *)
  tms_cstime : float }        (* System time for the children processes *)

The execution times (CPU times) of a process.

Chapter 16. The unix library: Unix system calls 183

type tm =
{ tm_sec : int;                       (* Seconds 0..59 *)
tm_min : int;                       (* Minutes 0..59 *)
tm_hour : int;                      (* Hours 0..23 *)
tm_mday : int;                      (* Day of month 1..31 *)
tm_mon : int;                       (* Month of year 0..11 *)
tm_year : int;                      (* Year - 1900 *)
tm_wday : int;                      (* Day of week (Sunday is 0) *)
tm_yday : int;                      (* Day of year 0..365 *)
tm_isdst : bool }                   (* Daylight time savings in effect *)

The type representing wallclock time and calendar date.

value time : unit -> int

Return the current time since 00:00:00 GMT, Jan. 1, 1970, in seconds.

value gettimeofday : unit -> float

Same as time, but with resolution better than 1 second.

value gmtime : int -> tm

Convert a time in seconds, as returned by time, into a date and a time. Assumes Greenwich meridian time zone.

value localtime : int -> tm

Convert a time in seconds, as returned by time, into a date and a time. Assumes the local time zone.

value alarm : int -> int

Schedule a SIGALRM signals after the given number of seconds.

value sleep : int -> unit

Stop execution for the given number of seconds.

value times : unit -> process_times

Return the execution times of the process.

value utimes : string -> int -> int -> unit

Set the last access time (second arg) and last modification time (third arg) for a file. Times are expressed in seconds from 00:00:00 GMT, Jan. 1, 1970.

User id, group id

value getuid : unit -> int

Return the user id of the user executing the process.

value geteuid : unit -> int

Return the effective user id under which the process runs.

value setuid : int -> unit

Set the real user id and effective user id for the process.

value getgid : unit -> int

Return the group id of the user executing the process.

value getegid : unit -> int

Return the effective group id under which the process runs.

value setgid : int -> unit

Set the real group id and effective group id for the process.

value getgroups : unit -> int vect

Return the list of groups to which the user executing the process belongs.

type passwd_entry =
  { pw_name : string;
    pw_passwd : string;
    pw_uid : int;
    pw_gid : int;
    pw_gecos : string;
    pw_dir : string;
    pw_shell : string }

Structure of entries in the passwd database.

type group_entry =
  { gr_name : string;
    gr_passwd : string;
    gr_gid : int;
    gr_mem : string vect }

Structure of entries in the groups database.

value getlogin : unit -> string

Return the login name of the user executing the process.

Chapter 16. The unix library: Unix system calls 185

value getpwnam : string -> passwd_entry

Find an entry in passwd with the given name, or raise Not_found.

value getgrnam : string -> group_entry

Find an entry in group with the given name, or raise Not_found.

value getpwuid : int -> passwd_entry

Find an entry in passwd with the given user id, or raise Not_found.

value getgrgid : int -> group_entry

Find an entry in group with the given group id, or raise Not_found.

Internet addresses

type inet_addr

The abstract type of Internet addresses.

value inet_addr_of_string : string -> inet_addr
value string_of_inet_addr : inet_addr -> string

Conversions between string with the format XXX.YYY.ZZZ.TTT and Internet addresses. inet_addr_of_string raises Failure when given a string that does not match this format.

Sockets

type socket_domain =
PF_UNIX                             (* Unix domain *)
| PF_INET                             (* Internet domain *)

The type of socket domains.

type socket_type =
SOCK_STREAM                         (* Stream socket *)
| SOCK_DGRAM                          (* Datagram socket *)
| SOCK_RAW                            (* Raw socket *)
| SOCK_SEQPACKET                      (* Sequenced packets socket *)

The type of socket kinds, specifying the semantics of communications.

type sockaddr =
ADDR_UNIX of string
| ADDR_INET of inet_addr * int

The type of socket addresses. ADDR_UNIX name is a socket address in the Unix domain; name is a file name in the file system. ADDR_INET(addr,port) is a socket address in the Internet domain; addr is the Internet address of the machine, and port is the port number.

type shutdown_command =
SHUTDOWN_RECEIVE                    (* Close for receiving *)
| SHUTDOWN_SEND                       (* Close for sending *)
| SHUTDOWN_ALL                        (* Close both *)

The type of commands for shutdown.

type msg_flag =
MSG_OOB
| MSG_DONTROUTE
| MSG_PEEK
The flags for recv, recvfrom, send and sendto.
value socket : socket_domain -> socket_type -> int -> file_descr

Create a new socket in the given domain, and with the given kind. The third argument is the protocol type; 0 selects the default protocol for that kind of sockets.

value socketpair :
   socket_domain -> socket_type -> int -> file_descr * file_descr

Create a pair of unnamed sockets, connected together.

value accept : file_descr -> file_descr * sockaddr

Accept connections on the given socket. The returned descriptor is a socket connected to the client; the returned address is the address of the connecting client.

value bind : file_descr -> sockaddr -> unit

Bind a socket to an address.

value connect : file_descr -> sockaddr -> unit

Connect a socket to an address.

value listen : file_descr -> int -> unit

Set up a socket for receiving connection requests. The integer argument is the maximal number of pending requests.

value shutdown : file_descr -> shutdown_command -> unit

Shutdown a socket connection. SHUTDOWN_SEND as second argument causes reads on the other end of the connection to return an end-of-file condition. SHUTDOWN_RECEIVE causes writes on the other end of the connection to return a closed pipe condition (SIGPIPE signal).

value getsockname : file_descr -> sockaddr

Return the address of the given socket.

value getpeername : file_descr -> sockaddr

Return the address of the host connected to the given socket.

value recv : file_descr -> string -> int -> int -> msg_flag list -> int
value recvfrom :
   file_descr -> string -> int -> int -> msg_flag list -> int * sockaddr

Receive data from an unconnected socket.

value send : file_descr -> string -> int -> int -> msg_flag list -> int
value sendto :
   file_descr -> string -> int -> int -> msg_flag list -> sockaddr -> int

Send data over an unconnected socket.

High-level network connection functions

value open_connection : sockaddr -> in_channel * out_channel

Connect to a server at the given address. Return a pair of buffered channels connected to the server. Remember to call flush on the output channel at the right times to ensure correct synchronization.

value shutdown_connection : in_channel -> unit

“Shut down” a connection established with open_connection; that is, transmit an end-of-file condition to the server reading on the other side of the connection.

value establish_server : (in_channel -> out_channel -> unit) -> sockaddr -> unit

Establish a server on the given address. The function given as first argument is called for each connection with two buffered channels connected to the client. A new process is created for each connection. The function establish_server never returns normally.

Host and protocol databases

type host_entry =
{ h_name : string;
h_aliases : string vect;
h_addrtype : socket_domain;
h_addr_list : inet_addr vect }

Structure of entries in the hosts database.

type protocol_entry =
{ p_name : string;
p_aliases : string vect;
p_proto : int }

Structure of entries in the protocols database.

type service_entry =
{ s_name : string;
s_aliases : string vect;
s_port : int;
s_proto : string }

Structure of entries in the services database.

value gethostname : unit -> string

Return the name of the local host.

value gethostbyname : string -> host_entry

Find an entry in hosts with the given name, or raise Not_found.

value gethostbyaddr : inet_addr -> host_entry

Find an entry in hosts with the given address, or raise Not_found.

value getprotobyname : string -> protocol_entry

Find an entry in protocols with the given name, or raise Not_found.

value getprotobynumber : int -> protocol_entry

Find an entry in protocols with the given protocol number, or raise Not_found.

value getservbyname : string -> string -> service_entry

Find an entry in services with the given name, or raise Not_found.

value getservbyport : int -> string -> service_entry

Find an entry in services with the given service number, or raise Not_found.

Terminal interface

The following functions implement the POSIX standard terminal interface. They provide control over asynchronous communication ports and pseudo-terminals. Refer to the termios man page for a complete description.

type terminal_io = {

Input modes:

mutable c_ignbrk: bool;  (* Ignore the break condition. *)
mutable c_brkint: bool;  (* Signal interrupt on break condition. *)
mutable c_ignpar: bool;  (* Ignore characters with parity errors. *)
mutable c_parmrk: bool;  (* Mark parity errors. *)
mutable c_inpck: bool;   (* Enable parity check on input. *)
mutable c_istrip: bool;  (* Strip 8th bit on input characters. *)
mutable c_inlcr: bool;   (* Map NL to CR on input. *)
mutable c_igncr: bool;   (* Ignore CR on input. *)
mutable c_icrnl: bool;   (* Map CR to NL on input. *)
mutable c_ixon: bool;    (* Recognize XON/XOFF characters on input. *)
mutable c_ixoff: bool;   (* Emit XON/XOFF chars to control input flow. *)

Output modes:

mutable c_opost: bool;   (* Enable output processing. *)

Control modes:

mutable c_obaud: int;    (* Output baud rate (0 means close connection).*)
mutable c_ibaud: int;    (* Input baud rate. *)
mutable c_csize: int;    (* Number of bits per character (5-8). *)
mutable c_cstopb: int;   (* Number of stop bits (1-2). *)
mutable c_cread: bool;   (* Reception is enabled. *)
mutable c_parenb: bool;  (* Enable parity generation and detection. *)
mutable c_parodd: bool;  (* Specify odd parity instead of even. *)
mutable c_hupcl: bool;   (* Hang up on last close. *)
mutable c_clocal: bool;  (* Ignore modem status lines. *)

Local modes:

mutable c_isig: bool;    (* Generate signal on INTR, QUIT, SUSP. *)
mutable c_icanon: bool;  (* Enable canonical processing
                            (line buffering and editing) *)
mutable c_noflsh: bool;  (* Disable flush after INTR, QUIT, SUSP. *)
mutable c_echo: bool;    (* Echo input characters. *)
mutable c_echoe: bool;   (* Echo ERASE (to erase previous character). *)
mutable c_echok: bool;   (* Echo KILL (to erase the current line). *)
mutable c_echonl: bool;  (* Echo NL even if c_echo is not set. *)

Control characters:

mutable c_vintr: char;   (* Interrupt character (usually ctrl-C). *)
mutable c_vquit: char;   (* Quit character (usually ctrl-\). *)
mutable c_verase: char;  (* Erase character (usually DEL or ctrl-H). *)
mutable c_vkill: char;   (* Kill line character (usually ctrl-U). *)
mutable c_veof: char;    (* End-of-file character (usually ctrl-D). *)
mutable c_veol: char;    (* Alternate end-of-line char. (usually none). *)
mutable c_vmin: int;     (* Minimum number of characters to read

                           before the read request is satisfied. *)
mutable c_vtime: int;    (* Maximum read wait (in 0.1s units). *)
mutable c_vstart: char;  (* Start character (usually ctrl-Q). *)
mutable c_vstop: char    (* Stop character (usually ctrl-S). *)
}
value tcgetattr: file_descr -> terminal_io

Return the status of the terminal referred to by the given file descriptor.

type setattr_when = TCSANOW | TCSADRAIN | TCSAFLUSH
value tcsetattr: file_descr -> setattr_when -> terminal_io -> unit

Set the status of the terminal referred to by the given file descriptor. The second argument indicates when the status change takes place: immediately (TCSANOW), when all pending output has been transmitted (TCSADRAIN), or after flushing all input that has been received but not read (TCSAFLUSH). TCSADRAIN is recommended when changing the output parameters; TCSAFLUSH, when changing the input parameters.

value tcsendbreak: file_descr -> int -> unit

Send a break condition on the given file descriptor. The second argument is the duration of the break, in 0.1s units; 0 means standard duration (0.25s).

value tcdrain: file_descr -> unit

Waits until all output written on the given file descriptor has been transmitted.

type flush_queue = TCIFLUSH | TCOFLUSH | TCIOFLUSH
value tcflush: file_descr -> flush_queue -> unit

Discard data written on the given file descriptor but not yet transmitted, or data received but not yet read, depending on the second argument: TCIFLUSH flushes data received but not read, TCOFLUSH flushes data written but not transmitted, and TCIOFLUSH flushes both.

type flow_action = TCOOFF | TCOON | TCIOFF | TCION
value tcflow: file_descr -> flow_action -> unit

Suspend or restart reception or transmission of data on the given file descriptor, depending on the second argument: TCOOFF suspends output, TCOON restarts output, TCIOFF transmits a STOP character to suspend input, and TCION transmits a START character to restart input.

The num library: arbitrary-precision rational arithmetic

The num library (distributed in contrib/libnum) implements exact-precision rational arithmetic. It is built upon the state-of-the-art BigNum arbitrary-precision integer arithmetic package, and therefore achieves very high performance.

The functions provided in this library are fully documented in The CAML Numbers Reference Manual by Valérie Ménissier-Morain, technical report 141, INRIA, july 1992 (available by

anonymous FTP from ftp.inria.fr, directory INRIA/publications/RT, file RT-0141.ps.Z). A

summary of the functions is given below.

Programs that use the num library must be linked in “custom runtime” mode, as follows:

        camlc -custom other options nums.zo other files -lnums

For interactive use of the num library, run camllight camlnum.

Mac: This library is not available.

PC: This library is available by default in the standard runtime system and in the toplevel system. Programs that use this library can be linked normally, without the -custom option.

17.1 num: operations on numbers

Numbers (type num) are arbitrary-precision rational numbers, plus the special elements 1/0 (infinity) and 0/0 (undefined).

type num = Int of int | Big_int of big_int | Ratio of ratio

The type of numbers.

value normalize_num : num -> num
value numerator_num : num -> num
value denominator_num : num -> num

Arithmetic operations

value prefix +/ : num -> num -> num
value add_num : num -> num -> num

Addition

value minus_num : num -> num

Unary negation.

value prefix -/ : num -> num -> num
value sub_num : num -> num -> num

Subtraction

value prefix */ : num -> num -> num
value mult_num : num -> num -> num

Multiplication

value square_num : num -> num

Squaring

value prefix // : num -> num -> num
value div_num : num -> num -> num

Division

value quo_num : num -> num -> num
value mod_num : num -> num -> num

Euclidean division: quotient and remainder

value prefix **/ : num -> num -> num
value power_num : num -> num -> num

Exponentiation

value is_integer_num : num -> bool

Test if a number is an integer

value integer_num : num -> num
value floor_num : num -> num
value round_num : num -> num
value ceiling_num : num -> num

Approximate a number by an integer. floor_num n returns the largest integer smaller or equal to n. ceiling_num n returns the smallest integer bigger or equal to n. integer_num n returns the integer closest to n. In case of ties, rounds towards zero. round_num n returns the integer closest to n. In case of ties, rounds off zero.

Chapter 17. The num library: arbitrary-precision rational arithmetic 193

value sign_num : num -> int

Return -1, 0 or 1 according to the sign of the argument.

value prefix =/ : num -> num -> bool
value prefix </ : num -> num -> bool
value prefix >/ : num -> num -> bool
value prefix <=/ : num -> num -> bool
value prefix >=/ : num -> num -> bool
value prefix <>/ : num -> num -> bool
value eq_num : num -> num -> bool
value lt_num : num -> num -> bool
value le_num : num -> num -> bool
value gt_num : num -> num -> bool
value ge_num : num -> num -> bool

Usual comparisons between numbers

value compare_num : num -> num -> int

Return -1, 0 or 1 if the first argument is less than, equal to, or greater than the second argument.

value max_num : num -> num -> num
value min_num : num -> num -> num

Return the greater (resp. the smaller) of the two arguments.

value abs_num : num -> num

Absolute value.

value succ_num: num -> num
     succ n is n+1
value pred_num: num -> num
     pred n is n-1
value incr_num: num ref -> unit

incr r is r:=!r+1, where r is a reference to a number.

value decr_num: num ref -> unit

decr r is r:=!r-1, where r is a reference to a number.

Coercions with strings

value string_of_num : num -> string

Convert a number to a string, using fractional notation.

value approx_num_fix : int -> num -> string
value approx_num_exp : int -> num -> string

Approximate a number by a decimal. The first argument is the required precision. The second argument is the number to approximate. approx_fix uses decimal notation; the first argument is the number of digits after the decimal point. approx_exp uses scientific (exponential) notation; the first argument is the number of digits in the mantissa.

value num_of_string : string -> num

Convert a string to a number.

Coercions between numerical types

value int_of_num : num -> int
value num_of_int : int -> num
value nat_of_num : num -> nat__nat
value num_of_nat : nat__nat -> num
value num_of_big_int : big_int -> num
value big_int_of_num : num -> big_int
value ratio_of_num : num -> ratio
value num_of_ratio : ratio -> num
value float_of_num : num -> float
value num_of_float : float -> num
value sys_print_num : int -> string -> num -> string -> unit
value print_num : num -> unit

17.2 arith_status: flags that control rational arithmetic

value arith_status: unit -> unit

Print the current status of the arithmetic flags.

value get_error_when_null_denominator : unit -> bool
value set_error_when_null_denominator : bool -> unit

Get or set the flag null_denominator. When on, attempting to create a rational with a null denominator raises an exception. When off, rationals with null denominators are accepted. Initially: on.

value get_normalize_ratio : unit -> bool
value set_normalize_ratio : bool -> unit

Get or set the flag normalize_ratio. When on, rational numbers are normalized after each operation. When off, rational numbers are not normalized until printed. Initially: off.

value get_normalize_ratio_when_printing : unit -> bool
value set_normalize_ratio_when_printing : bool -> unit

Get or set the flag normalize_ratio_when_printing. When on, rational numbers are normalized before being printed. When off, rational numbers are printed as is, without normalization. Initially: on.

value get_approx_printing : unit -> bool
value set_approx_printing : bool -> unit

Get or set the flag approx_printing. When on, rational numbers are printed as a decimal approximation. When off, rational numbers are printed as a fraction. Initially: off.

value get_floating_precision : unit -> int
value set_floating_precision : int -> unit

Get or set the parameter floating_precision. This parameter is the number of digits displayed when approx_printing is on. Initially: 12.

196

The str library: regular expressions and string processing

The str library (distributed in contrib/libstr) provides high-level string processing functions, some based on regular expressions. It is intended to support the kind of file processing that is usually performed with scripting languages such as awk, perl or sed.

Programs that use the str library must be linked in “custom runtime” mode, as follows:

   camlc -custom other options str.zo other files -lstr

For interactive use of the str library, run camllight camlstr.

Mac: This library is not available.

PC: This library is not available.

18.1 str: regular expressions and high-level string processing

Regular expressions

type regexp

The type of compiled regular expressions.

value regexp: string -> regexp

Compile a regular expression. The syntax for regular expressions is the same as in Gnu Emacs. The special characters are \$^.*+?[]. The following constructs are recognized: . matches any character except newline

* (postfix) matches the previous expression zero, one or several times + (postfix) matches the previous expression one or several times ? (postfix) matches the previous expression once or not at all [..] character set; ranges are denoted with -, as in a-z; an initial ^, as in ^0-9, complements the set ^matches at beginning of line $ matches at end of line \| (infix) alternative between two expressions \(..\) grouping and naming of the enclosed expression \1 the text matched by the first \(...\) expression (\2 for the second expression, etc) \b matches word boundaries \ quotes special characters.

value regexp_case_fold: string -> regexp

Same as regexp, but the compiled expression will match text in a case-insensitive way: uppercase and lowercase letters will be considered equivalent.

String matching and searching

value string_match: regexp -> string -> int -> bool

string_match r s start tests whether the characters in s starting at position start match the regular expression r. The first character of a string has position 0, as usual.

value search_forward: regexp -> string -> int -> int

search_forward r s start searchs the string s for a substring matching the regular expression r. The search starts at position start and proceeds towards the end of the string. Return the position of the first character of the matched substring, or raise Not_found if no substring matches.

value search_backward: regexp -> string -> int -> int

Same as search_forward, but the search proceeds towards the beginning of the string.

value matched_string: string -> string

matched_string s returns the substring of s that was matched by the latest

string_match, search_forward or search_backward. The user must make sure that the

parameter s is the same string that was passed to the matching or searching function.

value match_beginning: unit -> int
value match_end: unit -> int

match_beginning() returns the position of the first character of the substring that was

matched by string_match, search_forward or search_backward. match_end() returns

the position of the character following the last character of the matched substring.

value matched_group: int -> string -> string

matched_group n s returns the substring of s that was matched by the nth group \(...\) of the regular expression during the latest string_match, search_forward or search_backward. The user must make sure that the parameter s is the same string that was passed to the matching or searching function.

value group_beginning: int -> int
value group_end: int -> int

group_beginning n returns the position of the first character of the substring that was matched by the nth group of the regular expression. group_end n returns the position of the character following the last character of the matched substring.

Replacement

value global_replace: regexp -> string -> string -> string

global_replace regexp repl s returns a string identical to s, except that all substrings of s that match regexp have been replaced by repl. The replacement text repl can contain \1, \2, etc; these sequences will be replaced by the text matched by the corresponding group in the regular expression. \0 stands for the text matched by the whole regular expression.

value replace_first: regexp -> string -> string -> string

Same as global_replace, except that only the first substring matching the regular expression is replaced.

value global_substitute: regexp -> (string -> string) -> string -> string

global_substitute regexp subst s returns a string identical to s, except that all substrings of s that match regexp have been replaced by the result of function subst. The function subst is called once for each matching substring, and receives s (the whole text) as argument.

value substitute_first: regexp -> (string -> string) -> string -> string

Same as global_substitute, except that only the first substring matching the regular expression is replaced.

Splitting

value split: regexp -> string -> string list

split r s splits s into substrings, taking as delimiters the substrings that match r, and returns the list of substrings. For instance, split (regexp "[ \t]+") s splits s into blank-separated words.

value bounded_split: regexp -> string -> int -> string list

Same as split, but splits into at most n substrings, where n is the extra integer parameter.

Joining

value concat: string list -> string

Same as string__concat: catenate a list of string.

value join: string -> string list -> string

Catenate a list of string. The first argument is a separator, which is inserted between the strings.

Extracting substrings

value string_before: string -> int -> string

string_before s n returns the substring of all characters of s that precede position n (excluding the character at position n).

value string_after: string -> int -> string

string_after s n returns the substring of all characters of s that follow position n (including the character at position n).

value first_chars: string -> int -> string

first_chars s n returns the first n characters of s. This is the same function as

string_before.
value last_chars: string -> int -> string

last_chars s n returns the last n characters of s.

Formatting

value format: ('a, unit, string) printf__format -> 'a
Same as printf__sprintf.

Part V Appendix

201

Further reading

For the interested reader, we list below some references to books and reports related (sometimes loosely) to Caml Light.

19.1 Programming in ML

The books below are programming courses taught in ML. Their main goal is to teach programming, not to describe ML in full details — though most contain fairly good introductions to the ML language. Some of those books use the Standard ML dialect instead of the Caml dialect, so you will have to keep in mind the differences in syntax and in semantics.

• Pierre Weis and Xavier Leroy. Le langage Caml. InterÉditions, 1993.

The natural companion to this manual, provided you read French. This book is a step-by-step introduction to programming in Caml, and presents many realistic examples of Caml programs.

• Guy Cousineau and Michel Mauny. Approche fonctionnelle de la programmation. Ediscience, 1995.

Another Caml programming course written in French, with many original examples.

• Lawrence C. Paulson. ML for the working programmer. Cambridge University Press, 1991.

A good introduction to programming in Standard ML. Develops a theorem prover as a complete example. Contains a presentation of the module system of Standard ML.

• Jeffrey D. Ullman. Elements of ML programming. Prentice Hall, 1993.

Another good introduction to programming in Standard ML. No realistic examples, but a very detailed presentation of the language constructs.

• Ryan Stansifer. ML primer. Prentice-Hall, 1992.

A short, but nice introduction to programming in Standard ML.

• Thérèse Accart Hardin and Véronique Donzeau-Gouge Viguié. Concepts et outils de la programmation. Du fonctionnel à l’impératif avec Caml et Ada. InterÉditions, 1992.

A first course in programming, that first introduces the main programming notions in Caml, then shows them underlying Ada. Intended for beginners; slow-paced for the others.

• Rachel Harrison. Abstract Data Types in Standard ML. John Wiley & Sons, 1993.

A presentation of Standard ML from the standpoint of abstract data types. Uses intensively the Standard ML module system.

• Harold Abelson and Gerald Jay Sussman. Structure and Interpretation of Computer Programs. The MIT press, 1985. (French translation: Structure et interprétation des programmes informatiques, InterÉditions, 1989.)

An outstanding course on programming, taught in Scheme, the modern dialect of Lisp. Well worth reading, even if you are more interested in ML than in Lisp.

19.2 Descriptions of ML dialects

The books and reports below are descriptions of various programming languages from the ML family. They assume some familiarity with ML.

• Xavier Leroy and Pierre Weis. Manuel de référence du langage Caml. InterÉditions, 1993.

The French edition of the present reference manual and user’s manual.

• Robert Harper. Introduction to Standard ML. Technical report ECS-LFCS-86-14, University of Edinburgh, 1986.

An overview of Standard ML, including the module system. Terse, but still readable.

• Robin Milner, Mads Tofte and Robert Harper. The definition of Standard ML. The MIT press, 1990.

A complete formal definition of Standard ML, in the framework of structured operational semantics. This book is probably the most mathematically precise definition of a programming language ever written. It is heavy on formalism and extremely terse, so even readers who are thoroughly familiar with ML will have major difficulties with it.

• Robin Milner and Mads Tofte. Commentary on Standard ML. The MIT Press, 1991.

A commentary on the book above, that attempts to explain the most delicate parts and motivate the design choices. Easier to read than the Definition, but still rather involving.

• Guy Cousineau and Gérard Huet. The CAML primer. Technical report 122, INRIA, 1990.

A short description of the original Caml system, from which Caml Light has evolved. Some familiarity with Lisp is assumed.

• Pierre Weis et al. The CAML reference manual, version 2.6.1. Technical report 121, INRIA, 1990.

The manual for the original Caml system, from which Caml Light has evolved.

• Michael J. Gordon, Arthur J. Milner and Christopher P. Wadsworth. Edinburgh LCF. Lecture Notes in Computer Science volume 78, Springer-Verlag, 1979.

This is the first published description of the ML language, at the time when it was nothing more than the control language for the LCF system, a theorem prover. This book is now obsolete, since the ML language has much evolved since then; but it is still of historical interest.

• Paul Hudak, Simon Peyton-Jones and Philip Wadler. Report on the programming language Haskell, version 1.1. Technical report, Yale University, 1991.

Haskell is a purely functional language with lazy semantics that shares many important points with ML (full functionality, polymorphic typing), but has interesting features of its own (dynamic overloading, also called type classes).

19.3 Implementing functional programming languages

The references below are intended for those who are curious to learn how a language like Caml Light is compiled and implemented.

• Xavier Leroy. The ZINC experiment: an economical implementation of the ML language. Technical report 117, INRIA, 1990. (Available by anonymous FTP on ftp.inria.fr.)

A description of the ZINC implementation, the prototype ML implementation that has evolved into Caml Light. Large parts of this report still apply to the current Caml Light system, in particular the description of the execution model and abstract machine. Other parts are now obsolete. Yet this report still gives a complete overview of the implementation techniques used in Caml Light.

• Simon Peyton-Jones. The implementation of functional programming languages. Prentice-Hall, 1987. (French translation: Mise en œuvre des langages fonctionnels de programmation, Masson, 1990.)

An excellent description of the implementation of purely functional languages with lazy semantics, using the technique known as graph reduction. The part of the book that deals with the transformation from ML to enriched lambda-calculus directly applies to Caml Light. You will find a good description of how pattern-matching is compiled and how types are inferred. The remainder of the book does not apply directly to Caml Light, since Caml Light is not purely functional (it has side-effects), has strict semantics, and does not use graph reduction at all.

• Andrew W. Appel. Compiling with continuations. Cambridge University Press, 1992.

A complete description of an optimizing compiler for Standard ML, based on an intermediate representation called continuation-passing style. Shows how many advanced program optimizations can be applied to ML. Not directly relevant to the Caml Light system, since Caml Light does not use continuation-passing style at all, and makes little attempts at optimizing programs.

19.4 Applications of ML

The following reports show ML at work in various, sometimes unexpected, areas.

• Emmanuel Chailloux and Guy Cousineau. The MLgraph primer. Technical report 92-15, École Normale Supérieure, 1992. (Available by anonymous FTP on ftp.ens.fr.)

Describes a Caml Light library that produces Postscript pictures through high-level drawing functions.

• Xavier Leroy. Programmation du système Unix en Caml Light. Technical report 147, INRIA, 1992. (Available by anonymous FTP on ftp.inria.fr.)

A Unix systems programming course, demonstrating the use of the Caml Light library that gives access to Unix system calls.

• John H. Reppy. Concurrent programming with events — The concurrent ML manual. Cornell University, 1990. (Available by anonymous FTP on research.att.com.)

Concurrent ML extends Standard ML of New Jersey with concurrent processes that communicate through channels and events.

• Jeannette M. Wing, Manuel Faehndrich, J. Gregory Morrisett and Scottt Nettles. Extensions to Standard ML to support transactions. Technical report CMU-CS-92-132, Carnegie-Mellon University, 1992. (Available by anonymous FTP on reports.adm.cs.cmu.edu.)

How to integrate the basic database operations to Standard ML.

• Emden R. Gansner and John H. Reppy. eXene. Bell Labs, 1991. (Available by anonymous

FTP on research.att.com.)

An interface between Standard ML of New Jersey and the X Windows windowing system.

Index to the library

! (infix), 130 != (infix), 116 & (infix), 114 && (infix), 114

* (infix), 117, 120 ** (infix), 118 **. (infix), 118 **/ (infix), 192 *. (infix), 117 */ (infix), 192 + (infix), 117, 120 +. (infix), 117 +/ (infix), 192

- (infix), 117, 120 -. (infix), 117 -/ (infix), 192 / (infix), 117, 120 /. (infix), 117 // (infix), 192 < (infix), 115 <. (infix), 118 </ (infix), 193 <= (infix), 115 <=. (infix), 118 <=/ (infix), 193 <> (infix), 115 <>. (infix), 118 <>/ (infix), 193 = (infix), 115 =. (infix), 118 =/ (infix), 193 == (infix), 116 > (infix), 115 >. (infix), 118 >/ (infix), 193 >= (infix), 115 >=. (infix), 118 >=/ (infix), 193 @ (infix), 127 ^(infix), 132 || (infix), 114

abs, 121

                                            abs_float, 119
                                            abs_num, 193
                                            accept, 186
                                            access, 178
                                            acos, 118

add, 139, 150, 153, 156, 158

                                            add_float, 117
                                            add_int, 120
                                            add_num, 192
                                            alarm, 183
                                            approx_num_exp, 194
                                            approx_num_fix, 194

arg (module), 137

                                            arith_status, 194
                                            arith_status (module), 194
                                            asin, 118

asr (infix), 121

                                            assoc, 129
                                            assq, 129
                                            atan, 118
                                            atan2, 118
                                            background, 165

Bad (exception), 138 baltree (module), 138

                                            basename, 140
                                            big_int_of_num, 194
                                            bind, 186
                                            black, 165
                                            blit_image, 168
                                            blit_string, 133

blit_vect, 135
blue, 165

bool (module), 113

bounded_split, 199

Break (exception), 161 builtin (module), 114

button_down, 169
catch_break, 161

cd, 57

ceil, 118
ceiling_num, 192

char (module), 115

char_for_read, 115
char_of_int, 115

chdir, 161, 179

check_suffix, 140
chmod, 178
choose, 158
chop_suffix, 140
chown, 178

clear, 150, 156, 159

clear_graph, 165
clear_parser, 154

close, 161, 176

close_box, 141
close_graph, 164
close_in, 127
close_out, 125
close_process, 180
close_process_in, 180
close_process_out, 180
close_tbox, 144
closedir, 179
combine, 130
command_line, 160

compare, 115, 139, 158

compare_num, 193
compare_strings, 133
compile, 56

concat, 132, 140, 200

concat_vect, 134
connect, 186
contains, 139
copy_vect, 134

cos, 118

                                            cosh, 118
                                            create_image, 168
                                            create_lexer, 152
                                            create_lexer_channel, 152
                                            create_lexer_string, 152
                                            create_string, 132
                                            current_dir_name, 140
                                            current_point, 166
                                            cyan, 165
                                            debug_mode, 57
                                            decr, 130
                                            decr_num, 193
                                            denominator_num, 191
                                            descr_of_in_channel, 176
                                            descr_of_out_channel, 176
                                            diff, 158
                                            directory, 57
                                            dirname, 140
                                            div_float, 117
                                            div_int, 120
                                            div_num, 192
                                            Division_by_zero (exception), 120
                                            do_list, 128
                                            do_list2, 128
                                            do_list_combine, 130
                                            do_stream, 131
                                            do_table, 151
                                            do_table_rev, 151
                                            do_vect, 135
                                            draw_arc, 166
                                            draw_char, 166
                                            draw_circle, 166
                                            draw_ellipse, 166
                                            draw_image, 167
                                            draw_string, 166
                                            dump_image, 167

dup, 179

                                            dup2, 179
                                            elements, 158

empty, 153, 157

Empty (exception), 156, 159 End_of_file (exception), 122

                                            end_of_stream, 131

environment, 173
eprint, 155

eprintf, 147, 155 eq (module), 115

eq_float, 118
eq_int, 120
eq_num, 193
eq_string, 133
equal, 158
err_formatter, 145
error_message, 173
establish_server, 187

exc (module), 116

except, 129
exceptq, 129
execv, 174
execve, 174
execvp, 174
exists, 129

exit, 122, 160 Exit (exception), 116 exp, 118

Failure (exception), 116

failwith, 116

fchar (module), 117

fchmod, 178
fchown, 178
fcntl_int, 178
fcntl_ptr, 178

filename (module), 140

fill_arc, 167
fill_circle, 167
fill_ellipse, 167
fill_poly, 167
fill_rect, 167
fill_string, 132
fill_vect, 135

find, 139, 150, 153

find_all, 150
first_chars, 200
flat_map, 129
float, 157

float (module), 117

float_of_int, 117

                                            float_of_num, 194
                                            float_of_string, 119
                                            floor, 118
                                            floor_num, 192
                                            flush, 124
                                            fold, 158
                                            for_all, 129
                                            force_newline, 142
                                            foreground, 165
                                            fork, 174
                                            format, 200

format (module), 140

                                            fprint, 155

fprintf, 146, 154

                                            frexp, 119

fst, 130

                                            fstat, 177

fstring (module), 119

                                            ftruncate, 176
                                            full_init, 157
                                            full_major, 149

fvect (module), 119

gc (module), 147

                                            ge_float, 118
                                            ge_int, 121
                                            ge_num, 193
                                            ge_string, 133

genlex (module), 149

get, 148

                                            get_approx_printing, 195
                                            get_ellipsis_text, 144
                                            get_error_when_null_denominator, 194
                                            get_floating_precision, 195
                                            get_formatter_output_functions, 144
                                            get_image, 168
                                            get_lexeme, 152
                                            get_lexeme_char, 152
                                            get_lexeme_end, 152
                                            get_lexeme_start, 152
                                            get_margin, 142
                                            get_max_boxes, 143
                                            get_max_indent, 143
                                            get_normalize_ratio, 195
                                            get_normalize_ratio_when_printing, 195

getcwd, 179
getegid, 184
getenv, 161
geteuid, 184
getgid, 184
getgrgid, 185
getgrnam, 185
getgroups, 184
gethostbyaddr, 188
gethostbyname, 188
gethostname, 188
getlogin, 184
getpeername, 187
getpid, 175
getppid, 175
getprotobyname, 188
getprotobynumber, 188
getpwnam, 185
getpwuid, 185
getservbyname, 188
getservbyport, 188
getsockname, 186
gettimeofday, 183
getuid, 184
global_replace, 199
global_substitute, 199
gmtime, 183
Graphic_failure (exception), 164

graphics (module), 164

green, 165
group_beginning, 199
group_end, 199
gt_float, 118
gt_int, 121
gt_num, 193
gt_string, 133
handle_unix_error, 173
hash, 151
hash_param, 151

hashtbl (module), 150 hd, 128

in_channel_length, 127
in_channel_of_descr, 176
include, 56

                                            incr, 130
                                            incr_num, 193
                                            index, 129
                                            index_char, 133
                                            index_char_from, 133
                                            inet_addr_of_string, 185
                                            init, 157
                                            init_vect, 134
                                            input, 126
                                            input_binary_int, 127
                                            input_byte, 127
                                            input_char, 126
                                            input_line, 126
                                            input_value, 127
                                            install_printer, 57

int, 157

int (module), 120

                                            int_of_char, 115
                                            int_of_float, 117
                                            int_of_num, 194
                                            int_of_string, 122
                                            integer_num, 192
                                            inter, 158
                                            interactive, 160
                                            intersect, 129
                                            invalid_arg, 116
                                            Invalid_argument (exception), 116

io (module), 122

                                            ioctl_int, 180
                                            ioctl_ptr, 180
                                            is_absolute, 140
                                            is_empty, 157
                                            is_integer_num, 192
                                            it_list, 128
                                            it_list2, 128

iter, 153, 156, 158, 159

                                            join, 200
                                            key_pressed, 169
                                            kill, 182

land (infix), 121

                                            last_chars, 200
                                            ldexp, 119
                                            le_float, 118

le_int, 121
le_num, 193
le_string, 133

length, 156, 159 lexing (module), 151

lineto, 166
link, 177

list (module), 127

list_it, 128
list_it2, 128
list_length, 127
list_of_vect, 135
listen, 186
lnot, 121
load, 56
load_object, 56
localtime, 183
lockf, 181

log, 118

log10, 118

lor (infix), 121

lseek, 176
lshift_left, 121
lshift_right, 121

lsl (infix), 121 lsr (infix), 121

lstat, 177
lt_float, 118
lt_int, 121
lt_num, 193
lt_string, 133

lxor (infix), 121

magenta, 165
major, 149
make_formatter, 145
make_image, 167
make_lexer, 149
make_matrix, 134
make_string, 132
make_vect, 134

map, 128 map (module), 153

map2, 128
map_combine, 130

                                            map_vect, 135
                                            map_vect_list, 135
                                            match_beginning, 198
                                            match_end, 198

Match_failure (exception), 24–26, 114

                                            matched_group, 198
                                            matched_string, 198

max, 116

                                            max_int, 121
                                            max_num, 193

mem, 129, 157

                                            mem_assoc, 129
                                            memq, 129
                                            merge, 158

min, 116

                                            min_int, 121
                                            min_num, 193
                                            minor, 148

minus, 117, 120

                                            minus_float, 117
                                            minus_int, 120
                                            minus_num, 192
                                            mkdir, 179
                                            mkfifo, 180

mod (infix), 120

                                            mod_float, 119
                                            mod_num, 192
                                            modf, 119
                                            modify, 139
                                            mouse_pos, 169
                                            moveto, 166
                                            mult_float, 117
                                            mult_int, 120
                                            mult_num, 192
                                            nat_of_num, 194
                                            neq_float, 118
                                            neq_int, 120
                                            neq_string, 133

new, 150, 156, 159

                                            nice, 175
                                            normalize_num, 191

not (infix), 114

Not_found (exception), 116

                                            nth_char, 132

num (module), 191

num_of_big_int, 194
num_of_float, 194
num_of_int, 194
num_of_nat, 194
num_of_ratio, 194
num_of_string, 194
numerator_num, 191

open, 161, 175

open_box, 141
open_connection, 187
open_descriptor_in, 126
open_descriptor_out, 124
open_graph, 164
open_hbox, 143
open_hovbox, 143
open_hvbox, 143
open_in, 126
open_in_bin, 126
open_in_gen, 126
open_out, 124
open_out_bin, 124
open_out_gen, 124
open_process, 180
open_process_in, 180
open_process_out, 180
open_tbox, 144
open_vbox, 143
opendir, 179

or (infix), 114

out_channel_length, 125
out_channel_of_descr, 176

Out_of_memory (exception), 116

output, 125
output_binary_int, 125
output_byte, 125
output_char, 125
output_compact_value, 125
output_string, 125
output_value, 125
over_max_boxes, 143

pair (module), 130

parse, 138

Parse_error (exception), 131, 154 Parse_failure (exception), 131 parsing (module), 153

                                            pause, 182
                                            peek, 156
                                            pipe, 179
                                            plot, 166
                                            point_color, 166

pop, 159

                                            pos_in, 127
                                            pos_out, 125
                                            power, 118
                                            power_num, 192
                                            pp_close_box, 146
                                            pp_close_tbox, 146
                                            pp_force_newline, 146
                                            pp_get_ellipsis_text, 146
                                            pp_get_formatter_output_functions, 146
                                            pp_get_margin, 146
                                            pp_get_max_boxes, 146
                                            pp_get_max_indent, 146
                                            pp_open_box, 146
                                            pp_open_hbox, 146
                                            pp_open_hovbox, 146
                                            pp_open_hvbox, 146
                                            pp_open_tbox, 146
                                            pp_open_vbox, 146
                                            pp_over_max_boxes, 146
                                            pp_print_as, 146
                                            pp_print_bool, 146
                                            pp_print_break, 146
                                            pp_print_char, 146
                                            pp_print_cut, 146
                                            pp_print_float, 146
                                            pp_print_flush, 146
                                            pp_print_if_newline, 146
                                            pp_print_int, 146
                                            pp_print_newline, 146
                                            pp_print_space, 146
                                            pp_print_string, 146
                                            pp_print_tab, 146
                                            pp_print_tbreak, 146
                                            pp_set_ellipsis_text, 146
                                            pp_set_formatter_out_channel, 146
                                            pp_set_formatter_output_functions, 146
                                            pp_set_margin, 146

pp_set_max_boxes, 146
pp_set_max_indent, 146
pp_set_tab, 146
pred, 120
pred_num, 193
prerr_char, 123
prerr_endline, 123
prerr_float, 123
prerr_int, 123
prerr_string, 123
print, 155
print_as, 141
print_bool, 141
print_break, 142
print_char, 122, 141
print_cut, 142
print_endline, 123
print_float, 123, 141
print_flush, 142
print_if_newline, 142
print_int, 123, 141
print_newline, 123, 142
print_num, 194
print_space, 142
print_stat, 148
print_string, 123, 141
print_tab, 144
print_tbreak, 144

printexc (module), 154 printf, 147, 155 printf (module), 154

push, 159

queue (module), 156

quit, 56

quo (infix), 120

quo_num, 192
raise, 116

random (module), 157

ratio_of_num, 194
read, 176
read_float, 124
read_int, 124
read_key, 169
read_line, 124

                                            readdir, 179
                                            readlink, 180
                                            really_input, 126
                                            recv, 187
                                            recvfrom, 187

red, 165

ref (module), 130

                                            regexp, 197
                                            regexp_case_fold, 198

remove, 139, 150, 153, 158, 161

                                            remove_printer, 57

rename, 161, 177

                                            replace_first, 199
                                            replace_string, 133

rev, 128

                                            rewinddir, 179

rgb, 165

                                            rhs_end, 154
                                            rhs_start, 154
                                            rindex_char, 133
                                            rindex_char_from, 133
                                            rmdir, 179
                                            round_num, 192
                                            s_irall, 160
                                            s_irgrp, 160
                                            s_iroth, 160
                                            s_irusr, 160
                                            s_isgid, 160
                                            s_isuid, 160
                                            s_iwall, 160
                                            s_iwgrp, 160
                                            s_iwoth, 160
                                            s_iwusr, 160
                                            s_ixall, 160
                                            s_ixgrp, 160
                                            s_ixoth, 160
                                            s_ixusr, 160
                                            search_backward, 198
                                            search_forward, 198
                                            seek_in, 127
                                            seek_out, 125
                                            select, 181
                                            send, 187
                                            sendto, 187

set, 148 set (module), 157

set_approx_printing, 195
set_color, 165
set_ellipsis_text, 144
set_error_when_null_denominator, 194
set_floating_precision, 195
set_font, 166
set_formatter_out_channel, 144
set_formatter_output_functions, 144
set_line_width, 166
set_margin, 142
set_max_boxes, 143
set_max_indent, 143
set_normalize_ratio, 195
set_normalize_ratio_when_printing, 195
set_nth_char, 132
set_print_depth, 57
set_print_length, 57
set_tab, 144
set_text_size, 166
setgid, 184
setuid, 184
shutdown, 186
shutdown_connection, 187
sign_num, 193
signal, 182

sin, 118

sinh, 118
size_x, 165
size_y, 165
sleep, 183

snd, 130

socket, 186
socketpair, 186
sort, 158

sort (module), 158

sound, 169

split, 130, 139, 199

sprintf, 155
sqrt, 118
square_num, 192

stack (module), 159 stat, 148, 177

std_err, 122

                                            std_formatter, 145
                                            std_in, 122
                                            std_out, 122

stderr, 122, 175 stdin, 122, 175 stdout, 122, 175 str (module), 197 stream (module), 131

                                            stream_check, 131
                                            stream_from, 131
                                            stream_get, 131
                                            stream_next, 131
                                            stream_of_channel, 131
                                            stream_of_string, 131

string (module), 132

                                            string_after, 200
                                            string_before, 200
                                            string_for_read, 133
                                            string_length, 132
                                            string_match, 198
                                            string_of_bool, 114
                                            string_of_char, 115
                                            string_of_float, 119
                                            string_of_inet_addr, 185
                                            string_of_int, 122
                                            string_of_num, 194
                                            sub_float, 117
                                            sub_int, 120
                                            sub_num, 192
                                            sub_string, 132
                                            sub_vect, 134
                                            substitute_first, 199
                                            subtract, 129
                                            succ, 120
                                            succ_num, 193
                                            symbol_end, 153
                                            symbol_start, 153
                                            symlink, 180

sys (module), 159

Sys_error (exception), 159

                                            sys_print_num, 194
                                            system, 175
                                            system_command, 161
                                            take, 156

tan, 118

tanh, 118
tcdrain, 190
tcflow, 190
tcflush, 190
tcgetattr, 190
tcsendbreak, 190
tcsetattr, 190
text_size, 166

time, 161, 183

times, 183

tl, 128 toplevel (module), 55

trace, 56
transp, 167
truncate, 176
umask, 178

union, 129, 158 unix (module), 171 Unix_error (exception), 173

unlink, 177
untrace, 57
utimes, 183

vect (module), 134

vect_assign, 134
vect_item, 134
vect_length, 134
vect_of_list, 135
verbose_mode, 57
wait, 174
wait_next_event, 168
waitopt, 174
waitpid, 174
white, 165
write, 176
yellow, 165

Index of keywords

and, see let, type, exception, value, where

as, 21

begin, 22, 23

do, see while, for
done, see while, for
downto, see for
else, see if

end, 22, 23

exception, 31, 32

for, 22, 26

fun, 22
function, 22, 35

if, 22, 25

in, see let

let, 22, 24

match, 22, 26, 35

mutable, 30, 37
not, 22
of, see type, exception

or, 22, 26

prefix, 22, 28, 39

rec, see let, where
then, see if
to, see for

try, 22, 27 type, 30, 32

value, 32

                                            when, 36
                                            where, 37
                                            while, 26
                                            with, see match, try